### v1

# DoubleAgents: Exploring Mechanisms of Building Trust with Proactive AI

Abstract.
Agentic workflows promise efficiency, but adoption hinges on whether people actually trust systems that act on their behalf. We present DoubleAgents, an agentic planning tool that embeds transparency and control through user intervention, value-reflecting policies, rich state visualizations, and uncertainty flagging for human coordination tasks. A built-in respondent simulation generates realistic scenarios, allowing users to rehearse, refine policies, and calibrate their reliance before live use. We evaluate DoubleAgents in a two-day lab study (n=10), two deployments (n=2), and a technical evaluation. Results show that participants initially hesitated to delegate but grew more reliant as they experienced transparency, control, and adaptive learning during simulated cases. Deployment results demonstrate DoubleAgents’ real-world relevance and usefulness, showing that the effort required scaled appropriately with task complexity and contextual data. We contribute trust-by-design patterns and mechanisms for proactive AI—consistency, controllability, and explainability—along with simulation as a safe path to build and calibrate trust over time.
generative AI, AI agents, trust, reliance, transparency, state tracking, seminar organizing, control, planning, human coordination, reasoning, constitutional AI
copyright: acmcopyrightjournalyear: 2025conference: ; ; booktitle: Proceedings of the 2025price: 15.00isbn: 978-1-4503-XXXX-X/18/06ccs: Human-centered computing Interactive systems and tools
1. Introduction
Coordinating human effort is essential for tasks at every scale. Event planning involves finding suitable times, locations, vendors, and transportation. Warehouses must assign workers with the right skills and experience to shifts they are available for. Even organizing a speaker series requires emailing potential speakers, collecting their availability, and assigning them to appropriate slots. These problems involve both discovering people’s availability and solving the constraints of assigning individuals to limited roles. Human coordinators face time-consuming and intricate communication challenges: people may hesitate to disclose their full availability, delay responses, or provide only tentative answers such as “I might be able to come Saturday, but I’m not sure yet.” Coordinators must also balance group dynamics, make exceptions for important individuals, and continuously replan when last-minute changes arise. Coordination is thus both tedious and delicate, requiring continuous monitoring and re-planning while respecting social norms and relationships.
Generative AI now possesses several key capabilities that could help with coordination, including deep reasoning to evaluate constraints and produce plans (Guo et al., 2025; Yao et al., 2023b, a; Stechly et al., 2025), and API tool-calling to take concrete actions (Trivedi et al., 2024; Yao et al., 2024; Masterman et al., 2024). However, people do not yet trust AI to act proactively on their behalf for complex and delicate tasks like human coordination problems. AI systems are prone to hallucinations, often lack sufficient context to understand the problem (Zhang et al., 2025; Athaluri et al., 2023), and may mishandle delicate social interactions (Zhou et al., 2023; Wang et al., 2024). Most importantly, they lack an awareness of when to stop and escalate to human oversight. As a result, current AI interfaces largely wait for users to issue prompts and micromanage responses. This limits productivity and prevents AI from relieving people of burdensome procedural work. To unlock proactive AI, we need frameworks that let humans observe, guide, and calibrate trust in AI over time.
Trust is not established by statements alone—it develops through repeated reliable behavior over time (Razin and Feigh, 2024; Fouragnan et al., 2013). Game theory suggests that trust between humans is built by repeated observation of behavior in varied circumstances (Kollock, 1998). Additionally, in the AI trust and safety literature, Constitutional AI (Bai et al., 2022) is a paradigm that uses human-articulated policies (called a constitution) that can align an AI’s behavior with human expectations and values. Policies in making expectations explicit - so that humans and AI have a shared understanding of the expected behavior. Repeated interactions and clear policies are two pillars of establishing trust that we build upon to help build trust in proactive AI.
Figure 1. System diagram of DoubleAgents illustrating a day-by-day ReAct workflow that couples policy-guided planning with human oversight and LLM simulation. Inputs (top left) include the organizer’s goal, seminar slots, speaker personas, and a toolbox of callable functions. The coordination agent (center) iteratively: summarizes state; selects applicable policies; proposes a plan and action that the user can Regenerate or Approve; executes by drafting and sending emails, or waiting for responses. If a reply falls outside policy coverage, issueFlag escalates an edge case to the user for clarification before proceeding. The context management (right) maintains time information, past plans/actions, email logs, user preferences, and persona data, continuously updating to ground subsequent steps. Simulated respondent agents (bottom right) generate realistic, persona-consistent behaviors and replies that drive the loop across days. For a detailed walkthrough, please refer to Section 4.
[…]
Diagram showing the architecture of the DoubleAgents system, with components for goal setting, agent coordination, context management (including edge case detection), and various submodules such as toolbox, orchestrator, and simulated respondents.
We propose to build trust in proactive AI through simulation and policy refinement. Human behavior can be unpredictable, but AI-driven simulations can generate a wide range of realistic scenarios (Park et al., 2023), including late, incomplete, or uncertain replies. These allow coordinators to test and refine policies in safe environments. Policies specify how the AI should act (e.g., when to follow up, when to stop, how to phrase outreach) and when to escalate to human oversight. Hence, we propose to use simulation to expose AI behavior across repeated interactions, letting users correct and evolve policies iteratively. Visualizations of system state—goal progress, a timeline view, and person-level interaction histories—provide transparency and help humans intervene efficiently when needed.
We present DoubleAgents, shown in Figure 1, an interactive agentic system designed to automate aspects of human coordination while supporting user oversight and trust-building. DoubleAgents integrates generative AI agents in two complementary roles: (1) a ReAct Loop (Yao et al., 2023b) with AI planning and execution agents that can call tools to get appropriate context, query the current state of the system, and execute actions such as emailing potential speakers. This loop is extended to provide interaction for human interactivity and oversight at every step. (2) A LLM-based simulation of respondents that generate the varied and realistic human responses for various respondent personas (e.g. a professor who is very slow to respond, or a speaker who won’t commit). The LLM simulations allow users to experience repeated execution of the system to understand how it works, evolve the policies over time to suit their preferences, and potentially build trust that the system will act according to their wishes. By combining policy generation and refinement, action execution, and multi-agent LLM simulation within a user-centered feedback process, DoubleAgents helps users delegate repetitive aspects of coordination while retaining control of high-stakes interactions.
In a lab study (n=10) of people using the system with simulated responses over two study sessions, participants began hesitant to delegate but, through visible state, explicit reasoning based from policies, strong intervention points, and rehearsal on realistic cases, grew more comfortable relying on the agent for both planning and execution. In two real-world deployments of the system, we found that organizers adopted DoubleAgents as transparency, policy controls, and edge‑case flagging built trust; however, required more upfront context setup and tighter email integration.
Our work demonstrates how agentic systems can support reliable, human-centered coordination by embedding simulation, transparency, and interactive oversight into the agentic workflow. Specifically, we contribute:
  • A policy-driven, intervention‑friendly approach to govern proactive AI behavior in an agentic and integactive ReAct framework to build trust with users.
  • An AI-driven simulation framework of human responses to support repeated observations of reliable behavior over time to test and tune policies for AI to follow.
  • A user study (n=10) and a deployment with two users to evaluate how users perceive and gradually trust proactive AI for a human coordination task of scheduling a speaker seminar.
2. Related Works
2.1. Agentic AI in Task Planning and Execution
“AI Agents” typically refers to AI that can perform tasks for people. This can be done through tool use (namely, calling an API) (Schick et al., 2023), or operating a computer (Anthropic, 2025; OpenAI, 2025; Manus AI, 2025; Labs, 2025). Although AI Agents show impressive potential to act, trusting agents to act on behalf of a human is a critical missing piece (Afroogh et al., 2024). The delegation of control to AI raises questions not only about competence, but also about alignment, transparency, and user oversight.
Planning and execution remain central yet unresolved challenges in the development of AI agents. A growing body of work highlights both the difficulty and the importance of task planning, as seen in benchmarks and simulation environments such as Tau-Bench (Yao et al., 2024), NaturalPlan (Zheng et al., 2024), TravelPlanner (Xie et al., 2024), and AppWorld (Trivedi et al., 2024). These studies underscore that robust, goal-directed planning, particularly in open-ended or multi-step domains, is not only technically demanding but also essential for delivering meaningful user value.
To address this challenge, recent work has explored how LLMs can scaffold user planning by decomposing high-level goals into actionable steps (Zhang et al., 2024). Techniques such as ReAct (Yao et al., 2023b), which interleave reasoning and acting, have gained traction in multi-step environments for improving agent adaptability and responsiveness. These approaches are often embedded within multi-agent systems to support domains like travel, scheduling, and personal task management (Lal et al., 2023). However, autonomous systems such as AutoGPT (Significant Gravitas, [n. d.]) often fall short in transparency, adaptability, and robustness, limiting their usability in real-world settings (Zheng et al., 2024; Yao et al., 2024). Furthermore, prior cognitive science research shows that humans rarely engage in exhaustive upfront planning due to cognitive load (Krieger et al., 2009; Teevan et al., 2016), highlighting the need for AI systems that can support incremental, iterative, and contextualized planning over time. Additionally, recent studies suggest that human–AI co-planning, particularly when agents expose their intermediate reasoning and allow users to guide or intervene, can improve both task performance and user trust (He et al., 2025). When users can monitor and influence agent decisions at each step, they are better able to understand, calibrate, and ultimately trust the system, especially in domains requiring high-stakes delegation or multi-party coordination.
2.2. Human Simulations with LLMs
Recent work has explored the use of LLMs in multi-agent simulations  (Park et al., 2023), revealing their surprising capacity to model (Horton, 2023), predict, and even exhibit human-like behavior - including taking actions they were not explicitly instructed to perform in order to achieve high-level goals such as forming routines, organizing events, and collaborating spontaneously (Park et al., 2022, 2023). These behaviors often mirror human reasoning and coordination strategies. Other studies have shown that these behaviors closely align with empirical observations of humans across various domains, including economic games (Sreedhar et al., 2025; Sreedhar and Chilton, 2025; Aher et al., 2023; Horton, 2023), U.S. Supreme Court judgments (Hamilton, 2023), and over 80% of published and unpublished psychology experiments (Hewitt et al., 2024). These results suggest that LLM agents can serve as realistic proxies for humans in testing coordination, negotiation, and reasoning behaviors at scale.
AI systems that silate human behavior have already shown a striking capability to engage in complex, social behaviors competitively well with humans, for real world purposes. . Beyond simulation, AI systems like CICERO (Team et al., 2022) have shown how language agents can engage directly with humans in dynamic, multi-party settings. In online Diplomacy games, CICERO achieved a top 10% ranking among human players - illustrating the potential for LLM-based systems to engage in trust-sensitive planning, social reasoning, and strategic communication. The MATRIX framework (Pang et al., 2024) improved alignment in constitutional AI evaluations by useing simulated AI dialogues to anticipate a range of perspectives and surface important value trade-offs. Many systems have successfully used LLM red teaming  (Purpura et al., 2025) to identify potential threats in systems. Although AI simulations might not perfectly replicate individual user behaviors (yet) - they can be used to explore edge cases, probe systems, and engage in sensitive dialogs with humans.
2.3. Constitutional AI
Constitutional AI is a framework for aligning AI systems with human values by training models to self-critique and revise their outputs based on a set of guiding principles, or a ”constitution”(Bai et al., 2022). Unlike reinforcement learning with human feedback (RLHF)(Ouyang et al., 2022), which requires human evaluation for each model response, Constitutional AI allows the model to evaluate its own behavior using pre-defined rules. These constitutions are typically expressed as human-readable policies such as “Don’t promote illegal activity” or “Be helpful and respectful,” and can be authored by anyone. Building on this idea, Collective Constitutional AI (Huang et al., 2024) introduced a participatory approach in which over 1,000 people contributed rules to a shared constitution. Models trained on this crowd-sourced constitution demonstrated qualitatively different behaviors compared to those trained on developer-authored rules—highlighting the diversity of human values and community norms, and the implications this has for model alignment. To create constitutions, interactive systems like ConstitutionMaker (Petridis et al., 2024) have shown that LLMs can help transform users’ natural language feedback into generalized alignment policies. By providing real-time AI assistance in policy creation, such systems help users articulate their values more effectively, reduce cognitive burden, and make the process of shaping AI behavior more accessible and inclusive. These advances open up opportunities for systems that incorporate user-driven policy creation and alignment into workflows that have real-world impacts.
3. Seminar Planning: A Human Coordination Problem
Human coordination—such as assigning work shifts or scheduling job interviews—is both cognitively demanding and socially delicate. Unlike tasks with stable inputs and predictable outcomes, coordination requires managing uncertainty, reconciling multiple perspectives, and sustaining relationships over time. Even small lapses—missed follow-ups, unclear tone, or misaligned expectations—can erode trust and derail collective progress. These challenges make coordination both essential and persistently difficult in many areas of knowledge work.
We study these challenges through one concrete and consequential setting: organizing an academic seminar series. A typical series spans an entire semester and involves inviting a new external speaker nearly every week, amounting to 8–12 speakers per term. Early on, organizers must identify potential speakers, gather their availability constraints, and balance their own, such as ensuring topical diversity across weeks and securing highly sought-after senior researchers who are harder to schedule.
However, the most unpredictable and challenging part is dealing with responses. Some people do not reply at all and require multiple follow-ups to elicit a response. Some give only vague and tentative replies that need repeated confirmation. Some can be flakey and unreliable, offering slots but later canceling. And many unexpected problem cases arise: people might be on parental leave, ask for visa letters or funding support, recommend a different speaker, request a different date or talk modality, or raise logistical demands. In such human coordination tasks, you can rarely predict what is coming. Therefore, conventional scheduling tools such as Calendly [^1] https://calendly.com/ or When2Meet [^2] https://www.when2meet.com/ are ill-suited to this setting. They assume stable constraints and one-shot coordination, yet seminar scheduling requires ongoing negotiation, persuasion, and relationship maintenance. Unlike ordinary meeting scheduling, seminar speakers are doing you a favor, which means you cannot simply send them a link to pick a time—you must persuade and negotiate. Social dynamics and etiquette also play a large role: organizers must time their emails carefully to avoid overwhelming, annoying, or losing momentum with potential speakers, who are often busy scholars. Knowing when to nudge, follow up, or wait is subtle. Email remains the standard medium for professional and academic outreach, allowing organizers to adapt tone—formal, casual, or enthusiastic—to different speakers in ways scheduling links cannot. It also supports sharing rich context, such as talk topics, background, past speakers, or travel plans—far beyond simply choosing a calendar slot.
Thus, experienced organizers adopt personal strategies to cope with such a demanding task. For example, some follow a rule of “wait two days, then follow up” to nudge busy people who may not read emails right away, while others wait a full week before reaching out again. Some prefer contacting close collaborators or junior researchers first, while others prioritize securing prominent professors or distant speakers early to reduce travel constraints. When conflicts arise between speakers, organizers might “give faculty priority and ask students to be more flexible,” carefully deciding whom to approach for alternative availability to minimize disruption. For non-responders or last-minute cancellations, they must judge when and how often to follow up before moving on to new options—and sometimes even decide when it is appropriate to “beg” busy invitees for additional availability. Because plans are constantly evolving, organizers become locked in a cognitively and time-intensive loop of tracking shifting states, replanning, and executing revised plans as new information arrives—yet they often cannot proceed with the next steps or finalize allocations until replies are received.
Since seminar scheduling requires nuanced, context-dependent decisions that reflect each organizer’s values and preferences, it is difficult to fully automate. An effective seminar scheduling tool should not attempt to handle the process independently; instead, it must learn from and adapt to the organizer, respecting social norms, relational dynamics, and the inherent unpredictability of responses.
4. DoubleAgents System
Figure 2. A screenshot of DoubleAgents coordinating the assignment of four speakers to four seminar slots. (A) Policy Panel – Displays coordination policies that reflect user values and preferences. Users can add, edit, or delete policies, which guide the reasoning and planning of the coordination agent. (B) Interactive Chat Interface – Shows the full history of plans and actions, allowing users to monitor planning and execution, confirm or modify agent suggestions, and intervene when necessary. (C) Plan and Action Pop-up – A focused view within the chat history where users can review, edit, regenerate, or approve individual plans and actions. (D) Assignment Tracker – Visualizes tentative speaker-to-slot assignments and indicates each speaker’s availability. Dark purple denotes confirmed availability; light purple signals ambiguous availability requiring user confirmation. (E) Communication History Viewer – Displays the messaging history between the coordination agent and each speaker, organized by date and interaction. (F) Calendar View – Provides temporal context by showing the current simulation and upcoming seminar dates, color-coded for clarity.
We present DoubleAgents, an interactive agentic system designed to investigate how users build and calibrate trust in proactive AI systems. Built on the ReAct framework (Yao et al., 2023b), DoubleAgents enables iterative planning and execution through a team of AI agents that assist users in complex coordination tasks. In this work, we focus on the use case of assigning seminar speakers to available time slots. The system takes as input a user-defined goal (e.g., “Organize a seminar with four speakers and four slots.”) and produces as output a finalized assignment of speakers to slots. DoubleAgents consists of six core components, grouped into the following three functional categories. First, it contains the Core ReAct Agents that carry out the reasoning, planning, and simulation logic that drive the coordination process:
  • Coordination Agent - Responsible for reasoning over the current problem state using a set of predefined policies. It proposes plans and actions, and invokes appropriate tools to execute them.
  • Simulated Respondents - Represent background agents that emulate human respondents’ behavior by responding to user-generated emails based on predefined personas and descriptions.
Second, DoubleAgents incorporates the System Orchestration Modules, which manage the internal context and ensure robustness during decision-making:
  • Context Management Module - Maintains all relevant contextual information related to the task, including simulated respondents’ details (e.g., personas and descriptions), email logs from prior time steps, and the history of user-approved plans, actions, and emails.
  • Edge Case Detector - Identifies situations that fall outside the coverage of existing policies and escalates them to the user for clarification and additional instructions.
Lastly, the User Interaction Interfaces components enable transparency, user control, and real-time feedback during the coordination process:
  • Interactive Chat Interface - Facilitates communication between the user and the system. Users can confirm or revise generated plans, actions, and emails, and respond to flagged edge cases by providing natural language feedback.
  • Visualization Dashboard - Provides real-time feedback on the state of the task, including temporal progress, assignment status, and an activity tracker to help users monitor the overall coordination process.
In this section, we provide a walkthrough example to illustrate how DoubleAgents operates in practice. We also describe the implementation details of DoubleAgents. Consider the example of Lee, a PhD student organizing the HCI seminar series at his university. Lee plans to invite four speakers–Prof. A (a famous professor), Prof. B (a rising star professor), Student X (a senior PhD student), and Student Y (a PhD student who is a friend of Lee) – to deliver talks during four available time slots: Sep 5, Oct 4, Nov 7, and Dec 5.
Lee begins at the landing page, where he specifies his goals to the system. He types: “I am trying to organize a seminar with four speakers and four slots, help me reach out”. He then provides additional details, including the seminar location, and the dates of the four time slots. Next, Lee enters persona information for each speaker he wants to invite, which the system uses to generate simulated speaker profiles. The landing page is shown in Figure 3. To further customize communication, Lee uploads five of his past emails, allowing the system to adapt its generated messages to his style; for instance, casual for friends and formal for faculty. Finally, Lee initiates the simulation by clicking the “Start Planning” button, prompting the system to set up the scenario with the specified speakers and time slots. At this point, he is also shown a list of preset policies (see Figure 2(A)), which serve as guidelines for the coordination agent to begin planning. Lee reviews the policies, all of which appear reasonable. He may modify existing policies or introduce new ones to better guide the agent’s planning process.
Figure 3. The landing page to specify the user goals, seminar details, and the persona info for the speakers.
4.1. Context Management Module
At the beginning of the scheduling process, all information provided by Lee on the landing page is compiled into a structured problem context that the system consults at each step to solve the assignment task. DoubleAgents represents this problem context using JSON structures, where each key denotes a specific type of contextual information (e.g., speaker profiles, time constraints, email history, etc.), and each value contains the corresponding content. As shown in Figure 1, an example snapshot of the problem context includes keys such as ”time info”, which reflects the seminar slots specified by Lee. The context is dynamically updated throughout the interaction, incorporating user feedback, agent actions, and simulated responses to maintain an up-to-date representation of the task state.
4.2. Interactive Chat Interface
Upon clicking “Start Planning,” Lee is directed to the main Interactive Chat Interface (Figure 2(B)), where he can view the full history of system interactions and monitor the planning and execution progress carried out by the core ReAct agents. All plans, actions, and emails generated by the system are accessible through this interface. Each system update appears as an interactive pop-up window (Figure 2(C) shows a pop-up plan item), enabling Lee to inspect and confirm the system output in a controllable and traceable way. The interface also supports natural language queries, giving Lee full control over what to view. For example, he can type “What did Prof. B reply in the last email?” to retrieve the relevant message history directly within the chat. This chat functionality is powered by GPT-4o, with prompt details provided in the Appendix A.4.
4.3. Coordination Agent
4.3.1. Progress Understanding
When Lee initiates the planning process, the coordination agent assesses the overall state of the problem by generating a concise summary of the progress so far. To do that, it first reviews the current state of the scheduling process by analyzing the problem context, including speakers’ info, email logs from previous days, and the history of plans and actions previously approved by Lee. Based on this information, the coordination agent generates a summary of the current status, indicating which days have been scheduled, what actions have been taken, and the most recent contact date for each speaker to help monitor responses and follow-ups. The prompt used in generating the summary is illustrated in the Appendix A.1.1.
4.3.2. Policy Selection
To design a set of general policies applicable to a broad range of planning problems, five experts were recruited to brainstorm and define a framework of rules tailored to address common scheduling challenges. These challenges include: (1) individuals being reluctant to share their full availability, (2) ambiguous or unclear responses, (3) the need for multiple follow-ups to elicit responses, (4) resolving scheduling conflicts between two speakers, and (5) initiating the process by broadly reaching out to potential participants. Based on these identified challenges, we developed six initial ”seed policies” as a core set of guidelines for the system to reference and select from when generating plans. Each policy is designed to address a specific challenge, with two policies dedicated to handling non-responsive participants: one focused on strategies to cope with the lack of response (follow-up frequency) and another providing a framework for when to give up on follow-ups.
In Lee’s example, since it is the first day of scheduling, the progress summary informs the system that no speaker availabilities have been collected yet. As a result, the coordination agent invokes the policy selection tool, which identifies the appropriate policies based on the current problem state. Drawing from the list of ”seed policies”, the system selects: “If we are just beginning and have not yet contacted any speakers for availability, then ask each speaker for their availability across all slots.” The selected policy is then highlighted in the Policy Panel (as shown in Figure 2(A)), providing Lee with the rationale to generate the plan. The policy acts as the planning guideline, informing the agent’s next steps. Details of the prompts used for the policy selection module are provided in the Appendix A.1.2.
(a) The pop-up window for a generated plan.
(b) The pop-up window for generated actions following the plan.
(c) The pop-up window for a generated email.
Figure 4. Examples for plan and action generation.
4.3.3. Plan Generation
Guided by this policy, the coordination agent generates the following plan (shown in Figure 4(a)) for Lee: “Send email to Prof. A, Prof. B, Student X, and Student Y to request availability for slots on Sep 5, Oct 4, Nov 7, and Dec 5.”. Lee considers the plan and the rationale logical, since scheduling cannot proceed without first collecting availability information, and clicks ’Approve & Continue’ to move forward. The plan and its rationale are automatically recorded in the interactive chat interface (shown in Figure 1(B)), allowing Lee to review past activities and understand the rationale behind the plan. If Lee is dissatisfied with the proposed plan, he can either request the system to regenerate an alternative or directly edit the current plan himself, as shown in Figure 4(a).
4.3.4. Action Generation
To operationalize the plan, the coordination agent then generates a set of concrete actions to be executed (see Figure 4(b)). Specifically, it proposes sending individual emails to each speaker, with the stated goal: “Request for availability for Seminar Slots (Sep 5, Oct 3, Nov 7, Dec 5).” These proposed actions are presented to Lee for review. He examines each action and its corresponding goal, deems them appropriate, and clicks Approve to proceed. Importantly, Lee retains full control over the suggested actions: he may edit the goals, remove actions that are misaligned, or instruct the system to regenerate the set based on revised input. In general, there are three types of actions: (1) drafting emails to selected respondents, (2) sending emails to specified recipients using their provided addresses, and (3) awaiting responses from certain respondents. The detailed prompts used for plan and action generation are provided in the Appendix A.1.3.
4.3.5. Action Execution
Following approval, the coordination agent proceeds to instantiate each action by generating email drafts for the designated speakers, including both subject lines and message bodies (Figure 4(c)). This is done by calling the email writing tool, as specified in the tool box shown in Figure 1. Specifically, it calls the writeEmail() API, which drafts emails given the current problem state, seminar details, user-provided templates, and historical email edits. These elements allow the system to learn and adapt to the Lee’s preferred communication style, capturing tone, level of formality, and domain-specific phrasing, when generating new emails. Note that email drafting is performed using GPT‑4o, which is invoked by the coordination agent during the execution of the writeEmail() tool.
Lee reviews each draft and identifies one instance where the tone is not appropriate—the email to Prof. A, a distinguished professor, feels overly informal. To address this, Lee types: “Please make the email sound more formal, as Prof. A is a distinguished professor I’d like to invite.” The system then regenerates a more suitable version. Lee further refines the message by editing the first paragraph to better reflect his personal voice: “My name is Lee, a PhD student in Computer Science. Your work in [XXX] has greatly inspired us, and I would be honored to invite you to speak at our HCI seminar series.” After reviewing and finalizing all drafts, Lee is satisfied and clicks ‘Approve & Send’, prompting the system to dispatch the emails to all invited speakers. This is accomplished by invoking the sendEmail() API by the coordination agent, targeting the appropriate speaker email addresses specified in the problem context. In addition, all email-related activities are logged in the dashboard (Figure 1(E)), enabling Lee to easily monitor the status and history of interactions with each speaker. Besides writing and sending emails, the coordination agent may invoke actions involving delayed responses, such as waiting for a speaker’s reply. This is achieved by calling the waitResponse() tool. In such cases, no message is generated or sent; instead, the agent pauses execution and either proceeds to the next pending action or advances to the next simulated day. Overall, we use GPT-o3 as the backbone of the coordination agent, given its strong reasoning capabilities.
4.3.6. Implicit User Preference Learning
DoubleAgents supports implicit learning of user preferences by observing user-initiated edits to system-generated plans, actions, or emails. For instance, if the coordination agent proposes to reach out to Student X to ask for more availability, Lee may choose to remove the action “Send initial email to Student X”—perhaps because Student X’s prior responses have been vague or unhelpful. The system treats such edits as preference signals. To act on these signals, the coordination agent summarizes the edit into a concise statement, logs it in a user edit history, and incorporates it into the problem context. This updated context is then referenced in subsequent planning steps, enabling the agent to evolve its behavior in alignment with the user’s preferences over time. To streamline the summarization process, we leverage GPT-4o. Details of the prompt used for generating summaries of user edits across various content types can be found in Appendix A.1.5. The user-initiated edits, the corresponding system-generated drafts, and the generated summaries of the edits are all preserved in the user edit history. This comprehensive history serves as an evolving knowledge base, directly influencing the system’s generation of future plans, actions, and emails, and ensuring that it remains responsive to the nuances of user behavior.
4.4. Simulated Respondent Agent
To construct a realistic simulation environment for Lee, we draw inspiration from (Park et al., 2023) and employ LLMs to simulate each respondent agent, generating actions that align with their assigned personas. Concretely, upon receiving Lee’s emails, each simulated respondent agent first reasons over its persona and then generates an action that adheres strictly to those characteristics. For instance, the agent representing Prof. A opts not to respond on the same day (shown in Figure 1), as his persona specifies: “You are very busy and reply very late. Normally, it takes you three follow-up emails to respond.” In contrast, Prof. B’s persona indicates she is prompt in communication, leading her agent to reply immediately. Given that Prof.B’s persona also specifies preferences, such as “prefers presenting over Zoom”, the generated action will explicitly request a Zoom presentation option. The prompt for speaker agent, both for generating actions and drafting email responses, is detailed in the Appendix A.3. To execute the selected actions, the simulated respondent agents compose and send email responses, which are then received by Lee on his end. Figure 5(a) shows an example response received by Lee as shown in the interface. These responses advance the overall scheduling process and enable Lee to continue coordinating subsequent steps. Similar to the coordination agent, we adopt GPT-o3 as the backbone of each respondent agent to perform reasoning and generate plans and actions. GPT-4o is used when invoking the sendEmail() tool to generate email responses.
(a) Email sent by the simulated Prof. A.
(b) Edge case detection given response from Prof. B asking about Zoom options.
Figure 5. The example email sent by the simulated respondents and the system message when detecting edge case.
4.5. Edge Case Detector
Within a day, Lee receives an email from Prof. B stating: “Thanks for reaching out, Lee! I was wondering if there might be an option to present via Zoom instead of in person? That would provide more flexibility for me.” As this type of request had not been previously addressed by any existing policy, representing an edge case, the edge case detector issues a flag to prompt clarification from Lee. Through the Interactive Chat Interface, Lee sees the following system message: “Prof. B requests a Zoom option due to circumstances preventing in-person attendance. Should I confirm Zoom availability for their seminar or insist on in-person participation?” Preferring in-person interactions for seminars, Lee responds with the instruction: “Tell Prof. B we prefer in-person due to the potentially large audience.” The coordination agent then processes this clarification and generates a new action to inform Prof. B accordingly. Lee reviews the suggested action, finds it appropriate, and clicks Approve. An email draft is subsequently generated; after confirming the message aligns with his intent, Lee clicks ’Approve & Send’ to dispatch the response to Prof. B. In general, the edge case detector acts as a tool for the coordination agent to check if the message sent from the respondent agent can be addressed with the existing set of policies in order to make the next plans. When the detector identifies an edge case, the coordination agent escalates the issue to the user by presenting a clarification question in the interactive chat interface, prompting the user to provide guidance on how to proceed. The edge case detector is implemented with GPT-4o (see prompts in the Appendix A.2).
4.6. Dashboard Visualization
To support user understanding and foster trust, DoubleAgents provides clear visual feedback on the task state and system progress throughout the scheduling process. We implement a Flask-based dashboard interface designed to make the planning workflow interactive, legible, and actionable. The dashboard is composed of several layered visual components, shown in Figure 2(D), (E), and (F), which we describe in detail below:
  • Assignment Grid and Progress Tracker (Figure 2D): This panel visualizes tentative and confirmed speaker-to-slot assignments. It highlights which seminar slots remain unfilled and which speakers have been successfully scheduled. This layer helps Lee assess how close the current plan is to completion at a glance.
  • Calendar and Timelines (Figure 2F): Anchored in simulated days, this temporal layer helps Lee reason about timing, deadlines, and follow-up actions. The calendar highlights both the current simulated day and the seminar dates, allowing Lee to prioritize actions based on urgency and scheduling constraints.
  • Communication Visualization (Figure 2E): This panel displays the message history between the coordination agent and each speaker, offering a transparent view of the outreach and response status. It enables Lee to track ongoing communications, detect gaps, and assess follow-up needs.
These visual layers operate in parallel, allowing the system to efficiently track and update the state of each element in real time. Throughout the process, Lee interacts with the coordination agent and the respondent agents to collaboratively establish speaker-to-slot assignments for the seminar. Together, these visualization features provide Lee with a clear overview of assignment progress and a sense of what needs to be done on the current and subsequent days.
At the end of each simulated day, the system generates a daily summary that highlights key events and decisions. Specifically, we use GPT-4o to generate the daily summaries, with prompt details provided in the Appendix A.5. As shown in Figure 2(B), one example summary is presented within the Interactive Chat Interface and reads: “Invitations for the HCI seminar were sent to Profs. A, B, and Students X, Y. Prof. B and Student Y expressed interest, with scheduling pending. Both Prof. B and Student X inquired about Zoom options instead of in-person attendance.” After reviewing the summary, Lee confirms the day’s progress and clicks Continue to Next Day to proceed. This iterative, day-by-day process continues until all seminar slots are successfully assigned.
5. Technical Evaluation
DoubleAgents enables users to engage in iterative task planning and action execution through two core functionalities: (1) policy selection, which guides the generation of plans aligned with predefined coordination strategies; and (2) robust edge case detection, which identifies problematic or ambiguous situations and escalates them to the user for clarification and resolution. In this section, we describe the technical evaluation of the two components.
5.1. Policy Selection Experiments
5.1.1. Experiment settings
Policy selection refers to the process of identifying relevant coordination policies that guide the generation of high-quality, context-appropriate plans. Specifically, given a particular problem state (e.g., unfilled time slots and unresponsive speakers), the policy selection module is expected to retrieve all applicable policies that can serve as guidelines for the next planning step. For example, if no speakers have been contacted yet, a suitable policy would be: “If we are just beginning and have not yet contacted any speakers for availability, then ask each speaker for their availability across all slots.” In our implementation, we adopt a seed set of general-purpose coordination policies curated by domain experts. To inform policy selection, we use the output of the getProgressSummary() API, which provides a high-level snapshot of the current task state. An example summary reads: “One slot has been secured: speaker 4 confirmed for Nov 7 (2025-11-07) and received a confirmation email on 2025-08-07. All other speakers are still pending. Last outreach dates: speaker 1 and speaker 2 were last emailed on 2025-08-08; speaker 3 was last emailed on 2025-08-07.
We evaluated the effectiveness of our approach by comparing it against three baseline variants, each using a reduced context: (1) speaker-only context, which includes only information about speaker responsiveness; (2) slot-only context, which includes only slot-level availability (e.g., filled vs. unfilled); and (3) combined speaker and slot context, which merges the two but omits higher-level summaries. Prompt templates for all conditions are provided in Appendix B. To evaluate the performance of each approach, we constructed a test suite of 30 examples, all sourced from real user study sessions and curated to reflect a diverse set of scheduling scenarios. For each example, we treat the set of policies approved by users during the study as the ground truth for evaluation. All evaluations were conducted using the o4-mini model with a default temperature of 1. We report the average F1 score across three independent runs for each experimental setting.
5.1.2. Results and Findings
Prompting
Techniques
F1 Score
Mean SD
Speaker-only .27 .054
Slot-only .11 .038
Slot + Speaker .27 .017
Summary-based .70*** .035
Table 1. The technical evaluation results for Policy Section comparing four prompting techniques. The best scores are shown in bold. Statistically significant improvements (independent two-sample t-test, $p<0.001$) over all baselines are marked with ***.
As summarized in Table 1, our summary-based context approach achieves the highest F1 score for predicting relevant policies given a specific problem state. The improvement is statistically significant when compared to all three baseline conditions, with an average F1 score of 0.70, demonstrating the effectiveness of using structured progress summaries to guide policy selection. Notably, simply aggregating speaker- and slot-related context without summarization does not improve performance; the combined-context baseline yields only 0.27 average F1, indicating that raw context alone is insufficient for accurate policy retrieval. Based on these results, we adopt the summary-based approach as the default mechanism for selecting policies, which are then passed to the orchestrator agent to guide plan generation in a policy-aligned manner.
5.2. Edge Case Detection Experiments
5.2.1. Experiment settings
Edge case detection involves identifying whether a respondent’s message presents a scenario that falls outside the coverage of existing coordination policies. For example, a speaker requesting to present via Zoom, when no policy has addressed virtual presentations, should be flagged to the user for clarification and further instruction. To enable reliable detection of such edge cases, we adopt a few-shot prompting strategy with access to existing policies (denoted as ”few-shot w/ policy”). We compare this approach against two baselines: (1) zero-shot prompting without any demonstration examples, and (2) few-shot prompting without access to policy references (denoted as ”few-shot w/o policy”). Concrete prompt templates for each setup are provided in Appendix B.2. To evaluate performance, we constructed a test suite of 100 examples, manually validated by the lead authors to ensure coverage and quality. The construction methodology for this test suite is detailed in Appendix B.2. All experiments were conducted using GPT-4o, with temperature set to 0.7 and a maximum token limit of 100. We report the average accuracy across three independent runs for each condition.
5.2.2. Results and Findings
Prompting Techniques
Accuracy
Mean SD
Zero-shot .79 .021
Few-shot w/o Policy .88† .000
Few-shot w Policy .93** .015
Table 2. The technical evaluation results for Edge Case Detection comparing different prompting techniques. The best scores are shown in bold. Statistically significant improvements (independent two-sample t-test, p ¡ 0.01) over both baselines are marked with **, improvements over one baseline are marked with †.
As shown in Table 2, our few-shot prompting with policy method significantly outperforms both baselines, with statistically significant improvements across evaluation runs. These results demonstrate the effectiveness of incorporating policy context into the detection process, enabling the model to first analyze existing policies before determining whether a given case falls outside their scope. Based on these findings, DoubleAgents adopts the policy-aware few-shot approach as its default mechanism for edge case detection. This ensures that novel or ambiguous cases are reliably flagged to users, enabling timely clarification and maintaining a high-quality user experience.
6. User Study
6.1. Goal and Setup
To investigate the user-perceived effectiveness of DoubleAgents’ features and the dynamics of trust in human-AI collaboration, we recruited 10 participants to use DoubleAgents across two separate sessions. Our goal was to assess users’ perceived system usefulness, their reliance on the systems, and their trust and comfort in using AI to automate tasks had shifted. During each session, participants worked in a simulated environment to schedule a seminar with four dates and four speakers, while the system gradually adapted to their preferences. Participants completed a questionnaire assessing their trust in 15 different system features at three points: before the first session, after the first session, and after the second session. This allowed us to examine whether and when their trust and perceptions of the system changed over time.
6.2. Participants
We recruited 10 participants (7 female, 3 male; aged 21–30, M = 25.7) from our university community. All participants (10 of 10) reported using generative AI automation tools daily. Participants primarily used tools such as AI writing assistants (10 of 10), AI coding assistants (7 of 10), and AI design/creative tools (5 of 10). All participants reported solid understanding of generative AI: 7 described having in-depth knowledge of models and tools, while 3 reported a good general understanding of key concepts. Participants’ experience with human coordination tasks varied widely. 2 participants had moderate to extensive experience, having led team schedules or group planning multiple times. 3 participants had some experience, having organized small tasks or events independently. 4 participants had minimal experience, primarily helping or shadowing others in coordination tasks. Finally, 1 participant had no prior experience, having never hosted or coordinated any tasks.
6.3. Procedures and Data Analysis
Test Scenario Preparation.
We select eight distinct personas identified by three organizing experts, representing realistic and diverse roles, research areas, and backgrounds (e.g., a fifth-year PhD student studying VR currently on the job market; a busy junior professor in NLP who spends much of her time on her startup). For each persona, we specify reply behaviors (responsiveness, email writing style; e.g., slow responder, writes only short emails) and optional issues (potential concerns they might raise with speakers; e.g., requesting additional funding even though their in-person visit is already funded). We split them into two groups and created seminar organizing instances. We set the timeline as a monthly seminar schedule, with the first seminar one month after the simulation start date.
Day 1, Preparation and Setup (30min)
Participants first provided informed consent and completed a short AI perception survey assessing their perceptions and experiences with AI automation tools, human coordination tasks, and demographics. They also edited a few seeded templated seminars organizing emails according to their tone to feed the system their writing style. They were then introduced to the DoubleAgents interface and task. One researcher gave a brief walkthrough using a visual diagram and demonstration, covering the system’s core components, including the planning-execution orchestrator agent loo, respondent agent simulation, policy and context modules, visualization, and the flag mechanism.
Day 1, First System Usage and Evaluation (60min)
Participants then completed a seminar scheduling task (40 minutes), interacting with four simulated speakers with diverse personas, response behaviors, constraints, and coordination tensions. During the task, participants articulated and shared their mental models and usage strategies for DoubleAgents. Afterward, they completed another time of AI perception survey and a system experience survey assessing experience, workload, and attitudes toward specific components such as control, usefulness, and trust. Survey questions were adapted from NASA TLX (Hart, 2006) and prior AI perception and long-term evaluation studies (Ross et al., 2023; Long et al., 2024). Finally, a short interview asked participants to reflect on their overall experience, including any insights or policies they would like the system to incorporate before Day 2.
Day 2, Second System Usage and Evaluation (60min)
Before users’ next DoubleAgent usages, the researcher first ensured that the system incorporated participants’ first-day planning context and any preferences they had established via policies or system edits. Participants then completed another seminar scheduling task (40 minutes), followed by the same AI perception survey, the system experience survey, and a short interview. Day 2 interviews included follow-up questions comparing participants’ mental models, perceptions, experiences, and expectations of the system between Day 1 and Day 2.
Data Collection and Analysis
We logged participants’ interactions with AI suggestions, including regenerations, rejections, and edits, which were incorporated into DoubleAgents’ implicit learning module and our analysis. Interview data were analyzed using thematic analysis to understand how users perceive the usefulness and trust drive about each feature, and quantitative data were analyzed using paired-samples t-tests to examine whether participants’ experiences and perceptions of the DoubleAgents and AI tools across usage and days.
6.4. Result: Users’ Perceptions in AI Systems Shifted After Usages
Figure 6. Average user perception scores toward proactive AI and comfort with automating the planning stage in seminar organizing, across three stages (average scores across 10 users: before Session 1, after Session 1, and after Session 2). denotes statistical significance at $p<0.1$; * at $p<0.05$; ** at $p<0.01$; and *** at $p<0.005$.
Increasing Trust and Desire for Proactive Control.
Figure 6(a) shows statistically significant changes in participants’ trust in AI systems to act on their behalf, from 3.20/7 (before using the system, based on their previous AI experiences), to 5.00/7 after two times of using DoubleAgents. P1 noted that even though DoubleAgents represented just one specific experience with a particular AI system, “I felt more optimistic… [and] in seeing similar systems like DoubleAgents in the future, I would generally trust such AI systems more.
Additionally, there was a significant increase in wanting detailed explanations before the system acts, from 5.00/7 (before) to 6.00/7 (after). This suggests that seeing explanations from DoubleAgents made similar explanations from AI systems more desirable. P6 mentioned, “[With the system explanation] I felt like I could tell exactly what it’s doing at every stage so there’s a lot of transparency, [which] I think is pretty good… if I didn’t get that, I would feel lost [and] not know what’s going on…
Figure 7. Average user perception scores toward comfort with automating the execution stage in seminar organizing, across three stages (average scores across 10 users: before Session 1, after Session 1, and after Session 2). denotes statistical significance at $p<0.1$ and * at $p<0.05$.
Increasing Comfort with Automating the Planning Stage.
Figure 6(b) shows participants’ attitudes towards proactive AI in planning became significantly more positive. All of these metrics showed statistically significant differences: they felt more comfortable letting AI visualize the current status (4.80/7 to 6.50/7), figure out potential time assignments in the planning (5.10/7 to 6.38/7), decide daily next steps (4.70/7 to 6.50/7), and act as an assistant to monitor or flag issues (5.50/7 to 6.50/7). For example, P6 found the Communication Visualization section of the interface useful in showing the full email exchange histories between the organizer and each speaker. She mentioned, “Everything is consistent inside the system—when some speakers were not followed up, I could map it back to the visualization and confirm accuracy… [It] helped me build trust… For tomorrow, I’ll spend less time checking the email replies [inside the chat interface] one by one, and instead [just check] the Communication Visualization on the side.
For letting AI help monitor and point out potential issues, P1 described flagging as “one of the most important features of the system” and shared, “[With the issue flagging,] I skim [through the other features] now… It made me trust the system because I know it won’t just execute something random… It flags when it doesn’t know… That helps elicit my values and how I want to handle things—and [then DoubleAgents] respects that…
Increasing Comfort with Automating the Execution Stage.
Figure 7 shows that participants’ attitudes toward proactive AI in execution also became significantly more positive. For the tasks involving only drafting emails, as noted in the left column, all attitudes significantly improved, with comfort levels starting high and rising across the four types of emails (from around 6.0 before to nearly 7.0/7 after): initial outreach, follow-ups, clarification, and final confirmation. For the right column, draft-and-send tasks on users’ behalf, comfort also increased over time—rising for initial outreach from around 3.0/7 before to about 4.5/7 after, as well as for follow-ups and final confirmation. Participants demonstrated this after seeing DoubleAgents’ consistent performance over sessions: “I think [the system] acted quite reliably across two days, similar in terms of how it drafts emails [to the speakers], how it organizes language…” (P5). Participants also appreciated that the emails became more aligned to their tone and style over time—their email edits carried over to subsequent messages. For example, P4 mentioned that the system was learning her email writing style and that it remembered the phrasing she favored in email starters: “Compared to last time, I edited fewer times this time.,” which contributed to their trust in offloading more execution tasks to the AI, as well as a significantly increased desire for a system that aligns with and adapts to their styles over time (6.10/7 to 7.00/7, in Figure 6a).
One notable exception is drafting-and-sending clarification emails without approval, which showed little change (3.00/7 to 3.90/7 to 3.90/7). P10 explained that clarification emails often require nuanced, user-specific context or subjective judgment: “For clarification, I might want to double-check… because clarification sort of means that AI is also encountering a situation that was usually not in the plan [or policies]… I’m not comfortable [offloading that] to AI… AI is already very good at pretending like it knows what it is talking about…” As a result, they preferred to remain in the loop, reviewing or guiding these clarification messages before sending.
Additional Trust Driver: Simulation
All 10 participants emphasized that the run-throughs of simulated test cases helped establish confidence in the AI system’s robustness, contributing to trust. They found the simulations useful for surfacing a variety of problems that were realistic and representative of real-life cases. For example, P3 mentioned “The simulations [over the 2 days] felt realistic and true to the person [they] were simulating, like their personality and emails… The simulation elicited possible problems that would happen in real life… Like [having] people asking for funding or [needing] to reschedule — it forced me to think about how I would handle that [in real life], and that was good. I can trust the system to handle that in the future for me if scenarios like that occur, since it already happens in the simulation.
7. Deployment Study
As the in-lab study demonstrated, the simulation was helpful, and participants felt comfortable trusting AI in planning and executing human coordination tasks. However, would real-world seminar organizers use DoubleAgents? If so, how would they use it—would they plan, draft, and send invitation, follow-up, or clarification emails in different ways? Would trust be established in a similar manner? What additional challenges might arise in real-world use?
To explore these questions and evaluate the real-world deployability of DoubleAgents, we invited two organizers who were managing actual events during the preparation of this work. The goal was to observe how the system functions in realistic settings—including scheduling, tracking, email communication, and adaptive learning—and to understand how comfortable organizers are when the system sends real emails on their behalf.
7.1. Deployment Users and Procedures
Participants
We recruited two experienced seminar organizers (1 male, 1 female; average age 31), both with a good general understanding of generative AI and experience using AI for work—one on a daily basis. They are PhD students who have experience organizing research seminars for their departments. In this deployment, each is responsible for managing speaker events with audiences of over 30 participants: D1 is coordinating the Fall 2025 HCI seminar series at a university, selecting 10 speakers from a potential pool of 23 to fill 10 slots. D2 is organizing a panel discussion for a conference workshop, inviting 5–6 panelists from a potential 18 to fill a single slot. Each organizer was introduced to the system and used it to plan and manage their events.
Introduction, Setup, and Simulation Rounds
Organizers were first informed about the study and provided consent. They were then introduced to the study procedures and the DoubleAgents system through a demonstration, which included planning–execution loops and trust mechanisms, illustrated with a simulation example from the user study. Organizers also completed the initial survey covering AI perception and demographics.
To set up the system, organizers prepared a realistic start sheet by replacing simulated data with real organizing data (real potential speakers’ names, first with fake email addresses and later with the real ones) and provided their Gmail SMTP credentials. Clear instructions and consent procedures ensured again that participants understood data usage, privacy rights, and the option to withdraw. Before connecting their real email accounts, participants completed a 2-hour simulation round using pre-generated speaker personas based on the real potential speakers, allowing them to safely familiarize themselves with the system’s workflows. Finally, they tested the real email-sending module twice to confirm that emails were delivered correctly.
Deployment
After the simulation familiarization, organizers used DoubleAgents to initiate the coordination process and track it for over a few days. They continued receiving system updates over the following week. We communicated with participants consistently, logging interactions, usage patterns, and feedback on the system’s support for scheduling, tracking, and learning. Organizers’ system interactions—including email sends, scheduling updates, and system adaptations—were tracked. Qualitative feedback was collected via check-in meetings, focusing on usability, trust, concerns, and code support, while quantitative system logs were collected for system learning over the week.
7.2. Deployment Results
DoubleAgents’ Perceived Usefulness and Relevance in Live Deployment.
After the demonstration and simulation run-through, both organizers (D1 and D2) reported that DoubleAgents was highly relevant to their seminar-organizing workflow. They highlighted that the system fit seamlessly into their email-heavy tasks and that control settings and edge-case flagging enabled handling complex coordination scenarios that conventional scheduling tools like Calendly cannot manage. D1 emphasized that crafting emails is a delicate and important task, as it involves personalizing messages in order to “make the speakers feel they are special,” even when requesting a favor. Both D1 and D2 found DoubleAgents capable of generating good starter emails for edits at the beginning, as well as producing an acceptable final email to send. Organizers specifically appreciated that the simulation included the names of the real speakers and incorporated aspects of their personalities in constructing the respondent simulations and edge cases, which helped them anticipate challenges in live deployment.
System Onboarding, Testing, and Understanding in Live Deployment.
D1’s deployment setup took four days and five meetings. The first two meetings were devoted to debugging and ensuring the system ran smoothly for a scenario involving more than 20 speakers, 10 slots, and substantial context information for each speaker. D1 modified their invitation and follow-up email templates more than four times. Additionally, D1 updated the seminar speakers’ descriptions to enable the email generation module to generate precise, high-quality emails that incorporated extensive context information. After hardening the system and refining our onboarding protocols, D2’s setup took two days and two meetings. D2’s application in panelists invitation were also simpler (e.g., shorter panelist descriptions, invitations for one-day panels), which contributed to the faster setup. Both D1 and D2 reported that the practice simulation rounds were valuable, and that sending testing emails through SMTP allowed them to gradually grasp the system’s capabilities and limitations. This illustrates that system setup and onboarding are consistent during deployment, while the required effort scales with task complexity and the amount of context data.
Specifically, organizers emphasized the importance of supplying sufficient context to generate highly personalized emails. D1 mentioned, “[context management] is really tricky.” For example, one speaker D1 tried to invite had previously corresponded with a different organizer in a prior semester about her seminar availability on Fridays in a separate email thread. This semester, the seminar had shifted to Thursdays. To manage the context from past email threads, schedule updates, and templates, D1 opted to start a new thread with the speaker, as DoubleAgents at the time worked with new threads rather than replying within existing Gmail threads. In another case, D1 had already discussed seminar opportunities with a speaker on Discord but initially omitted this context during the early simulations. When composing the email, they noticed that the system had selected a generic template, so they updated the backend context with the appropriate tone and information to create a suitable follow-up. After practicing with real-speaker data in simulation, D1 finalized a preferred template for live sending. During the actual sending, they continued to refine messages—adding details and small fillers—and observed that these edits were retained and applied to subsequent emails.
Email Initiation, Response Handling, and Edge Cases in Live Deployment.
During live deployment, D1 sent emails to nine speakers. Over the earlier simulation rounds, she carefully perfected her initial templates and preferred writing styles via the implicit style transfer modules. Despite this, she still spent significant time refining the first few real emails. For the remaining speakers, she adopted a common real-world practice: copying previous templates into the text editor and adapting them for each recipient, which sped up the process. She mentioned that “the system’s email suggestions are helpful, though for similar invitation messages, I don’t always need a new draft each time.” D1 also used the system’s built-in AI chat feature seven times during initial email sending for phrasing recommendations. She noted that this support reduced mental demand and felt similar to receiving quick email feedback from a co-organizer, a practice she commonly uses in real-life organizing. D2, on the other hand, relied on the system’s automated drafting and style adaptation, demonstrating that trust and efficiency increased with experience. Interestingly, for some longer emails requiring scrolling to read, D2 skipped reviewing the full content and proceeded directly, trusting the system’s style transfer. At the end of the invitation email sending day, D2 sent emails to five potential panelists, relying directly on system-generated drafts based on his perfected templates and implicit style transfers from the previous simulation rounds.
After the initial invitation email were sent, D1 received responses as quickly as 4 minutes, 1 hour, and 2 hours. As replies came in, organizers encountered more complex edge cases than those seen during simulation: all three responding speakers had overlapping availability on the same dates, creating interlocked scheduling conflicts. DoubleAgents logged the replies in the backend and suggested that D1 wait for the remaining six speakers while coordinating with the student speaker and the faculty member closest to her school to adjust availability. D1 instead opted to wait for three days before responding (one day more than the policy instructed), as she was occupied with other commitments and felt comfortable replying later. In contrast, D2 received three negative responses after a full day and did not encounter similar complications, and the system suggested waiting for the remaining two people to reply.
Trust Establishment in Live Deployment
Over time, D1 emphasized that trust depended on the system’s closeness to ground truth—specifically, whether emails were actually sent and whether responses were logged correctly and efficiently. Accurate alignment between the system’s state visualization and the real state of email threads was critical to their trust. Also, D1 noted that, over time, understanding the system’s capabilities and limitations helped her feel more comfortable using the tool. D2 reported that they appreciated having policies and rationale always visible and editable, in contrast to AI tools that hide their reasoning. After sending real emails, D2 remarked, “I trust AI for specific use cases more,” explaining that they found the system to be dedicated to and capable of generating high-quality email.
At the same time, D2 compared DoubleAgents with other tools they were familiar with, abstracting their trust establishment process and the underlying barriers. He noted that the email interface differed from Gmail, introducing a small trust barrier. They compared Gmail’s richer text editor, easier insertion of tables and visuals, and familiar thread visualization—where they know what to expect when clicking—to our interface. They also expressed concern about formatting consistency for recipients (e.g., indentation, font sizes) in emails sent via DoubleAgents’ Gmail backend, noting that such issues can occur even within Gmail phone apps, which made them unsure how messages would appear to recipients. This concern was addressed through several rounds of email formatting and link embedding tests, during which they sent mails to their personal Gmails and checked them in parallel. Regarding integration, D2 felt that strong state visualization and tracking supported their trust through natural mapping. They believed their current practice of using spreadsheets to track speaker invitation progress could be integrated into the system’s visualization. For example, he usually rank the lists to monitor progress, and instead of serving solely as a display, the spreadsheet also functions as a goal-setting and action-taking tool. Thus, he expressed interest in making DoubleAgents’ visualization more interactive and editable, allowing them to directly initiate system actions. For actual adoption, he emphasized the need for scalability and further automation, requesting dynamic fields features similar to other email client tools like MailChimp. They also compared real email SMTP sending to Zoom microphone testing, emphasizing the value of having a trial, closed, and safe environment with strong validation—such as sending emails to themselves first.
Trust Establishment Mechanism in DoubleAgents How Humans Build Trust With Humans Nielsen Heuristics (Nielsen, 1992) Human-AI Design Challenges (Yang et al., 2020)
(1) Consistent and robust
Policy-based conditional reasoning grounded in human social cognition; reliable handling of varied test suites and edge cases
Trust grows when others are reliable, predictable, and follow shared norms or values. Consistency over time signals dependability. (Lansing et al., 2023; Rempel et al., 1985; Mayer et al., 1995; Beratan, 2007) - Consistency and standards
- Error prevention
- Help users recognize, diagnose, and recover from errors
- Unpredictability of AI behavior: users need repeated, consistent outputs to form expectations
- Data dependency: robustness across edge cases builds confidence
(2) Controllable and aligned with user values
User can confirm and intervene at every step; system adapts to users’ editing history and evolving policies
Trust strengthens when people respect autonomy, allow mutual control and negotiation, and align with shared goals or values (Lansing et al., 2023; Kramer, 1999). - User control and freedom
- Flexibility and efficiency of use
- Conceptual misalignment: user oversight helps align system behavior with goals
- Evolving mental models: adaptation supports shifting user practices
(3) Understandable and explainable
Real-time state tracking and visualization; clear reasoning and a chat window to ask questions
Humans trust others who are transparent, communicate intentions, and explain their actions clearly (Lansing et al., 2023; Kramer, 1999). - Visibility of system status
- Help and documentation
- Match between system and the real world
- Explainability gap: reasoning and state tracking reduce opacity and support accurate mental models
Table 3. How DoubleAgents’ trust-building dimensions map to human–human trust building literature (Rempel et al., 1985; Mayer et al., 1995; Kramer, 1999; Lansing et al., 2023; Beratan, 2007), Nielsen heuristics (Nielsen, 1992), and human-AI interaction design challenges (Yang et al., 2020).
8. Discussion: Trust Establishment with AI Agents Over Time, Consistency, Control, and Understanding
DoubleAgents illustrates how users establish trust in a novel AI system over time. Users initially expressed low trust because they were unfamiliar with the system, but through exploration, testing, learning, and habituation, they gradually began to rely on it. This aligns with prior longitudinal studies on AI tools (Long et al., 2024). Understanding how this trust emerged sheds light on the mechanisms by which human–AI relationships can evolve.
As shown in Table 3, DoubleAgents’ trust-establishment mechanisms center on three interdependent qualities: consistency, control, and understanding. These dimensions align with patterns of human–human trust development, are supported by classic usability heuristics (Nielsen, 1992), and address well-known challenges in human–AI interaction design (Yang et al., 2020).
First, one main design goal of DoubleAgents is to be consistent and robust. The system employs policy-driven, conditional reasoning to stabilize its outputs and ground its behavior in shared human values and social norms. This design enables reliable performance even across diverse testing suites and edge cases, allowing users to form stable expectations—echoing how consistency signals dependability in human relationships and mitigating issues of unpredictability and data dependency in AI systems. Second, DoubleAgents incorporates mechanisms to remain controllable and aligned with user values. The system enforces stepwise confirmations and allows user intervention throughout planning and execution, while adaptively learning from users’ past editing histories and evolving policy additions. These mechanisms promote a sense of autonomy and mutual negotiation, reflecting how humans build trust through control-sharing, and addressing conceptual misalignments and evolving mental models in human–AI collaboration. Third, DoubleAgents is designed to be understandable and explainable. It offers real-time state tracking, clear visualizations, explicit reasoning for its actions, and a chat interface for user queries. These mechanisms increase system transparency and communicability, similar to how open communication fosters trust between humans, and reduce the explainability gap in AI systems.
Collectively, these design mechanisms enabled users to gradually transition from skepticism to reliance. By operationalizing consistency, control, and understanding, DoubleAgents supported the iterative habituation process through which users internalized the system’s behavior and incorporated it into their workflows. This suggests that trust in proactive AI systems is not a fixed property but an emergent outcome—shaped by interaction designs that allow users to explore, test, and progressively align the system with their own goals and values.
Several participants compared building trust in DoubleAgents to their prior experiences with coding automation tools such as Cursor and CLine, reflecting the CS background of most participants. For example, P10 described first trying the tool on a safe project in a controlled environment (not pushing to a live website) to see how well it worked, illustrating the role of consistent system performance in building trust. They also highlighted that Cursor offered strong control and state-reverting capabilities, which allowed them to confirm each step and explore outcomes safely. Over time, they learned how to craft better prompts, demonstrating alignment with their goals and values. Finally, participants emphasized the importance of understandable reasoning: both Cursor and DoubleAgents provide stepwise explanations for each action, which helped them make sense of the system’s behavior and develop confidence in its outputs. As P4 reflected: “Most of the trust I had with Cursor was because it would ask at each step and then explain its reasoning for that step—and it was accurate. The same thing here [in DoubleAgents]: it’s accurate, and each step provides reasoning. It’s from the frequent confirmation and the reasoning always being correct that I built trust.”
These accounts illustrate that consistent performance, controllable interactions, and clear explanations collectively scaffolded the sophisticated trust formation process between AI and users, aligning users’ experiences with the mechanisms designed into DoubleAgents and with broader theories of habituation and technology acceptance.
9. Limitations and Future Works
First, DoubleAgents was initially evaluated through simulated seminar-organizing scenarios. While these simulations were reviewed by domain experts and deemed realistic, they may still lack the richness of real-world settings. Future work could make the ReAct-based coordination agent more user-aligned by incorporating users’ past email or organizing data and modeling their personal communication styles and values. Similarly, the respondent agent simulation could be made more realistic by drawing on a more diverse set of real human behavioral data. Second, our system used different models for different tasks—o3 for reasoning tasks and 4o-mini for other tasks. Although our technical evaluation showed they performed well, future models with stronger social and human cognition capabilities, as well as incorporating users’ data, may achieve better planning and execution. Third, although we deployed the system with two real organizers in real life, we also consulted others whose scenarios introduced additional complexities where the system may not perform perfectly. For example, one seminar organizer mentioned theirs involved bilingual invitations, and others described co-organizing dynamics, both of which introduce challenges not addressed in our current design. Finally, all our participants were experienced users of AI tools. Their familiarity may have positively biased their trust and comfort levels. Future work should include participants with a broader range of AI experience to better understand adoption barriers among less experienced users.
10. Conclusion
We present DoubleAgents, an agentic planning system that incorporates clear human intervention points, rich state visualizations, explicit policies, and uncertainty/edge‑case flagging to automate human coordination. In a two‑session lab study (n=10) and a week‑long deployment with two organizers, participants reported significantly greater comfort delegating planning and execution to DoubleAgents. Simulated respondents provided a safe proving ground to observe policy‑aligned planning and consistent system performance. From these studies, we distill three trust‑by‑design mechanisms—consistency, control, and understanding—operationalized via policy‑driven reasoning, stepwise approvals, implicit system adaptive learning, and a real-time system state. These patterns generalize beyond seminar scheduling to proactive agents design that must act while remaining auditable and aligned.
Appendix A DoubleAgents Prompts
A.1. Coordination Agent
A.1.1. Progress Summary
You are an expert seminar organizer assistant. Your task is to analyze the current progress and return a JSON object. Input information may include: - Speaker details (persona and description) - Possible dates - Scheduling rules - Email log - History of approved plans/actions - Time context - Any user suggestions You should mainly look through Email log and history of approved plans/actions to analyze the current progress. Do not assume any additional actions have been taken. Your JSON output should contain: - summary: A concise summary of the current state, including which days are scheduled, what actions have been taken, and the last date we reached out to each speaker (for tracking and follow-up). - un_or_filled_slots: Two lists of dates — one for filled slots and one for unfilled slots. - pending_speakers: A list of speaker names who have not yet responded or are pending (how many days since we last reached out to them). When mentioning speakers, use the format “English Name (ID)”. Avoid vague terms like *all speakers* or *pending speakers*. ---------- [Task Rules] [Possible Dates] [Time Context] [Speaker Information] Here are the history of Approved Plans and Actions: [Approved Plans and Actions] -------------------------------- Here is the history of past user edits and implicit learning (recent, action edits only): [Implicit Learning] -------------------------------- Past Email Log: [Email Log]
A.1.2. Policy Selection
You are a helpful assistant. Given the following main reasoning result and context, select which of the following coordination policies should be applied. The chosen policies should be 100% applied to the scenario! If you found that we didn’t do anything yet from the reasoning, you should choose the ’we are just beginning’ policy Return a JSON list of the indices of the applicable policies (e.g., [1,2]) or [1]. If none apply, return an empty list []. [Main Reasoning Result] [Coordination Policies] Return ONLY the JSON list of indices. [Implicit Learning/User Edit History (recent, action edits only)]
A.1.3. Action Generation
You are an expert seminar organizer assistant. Your task: Given the organizing progress summary, the selected coordination policies, and the current time context, generate a concise plan for what must be accomplished today only. Output format: Return only one single string that contains: 1. The plan (concise, no bullet points, no unnecessary multiple sentences). 2. A short rationale starting exactly with "Why this plan? ...". The rationale should be about 40 words explaining why this plan was chosen according to the task progress summary context and how the policies/time context informed the decision. Example: "Send email to X to request availability for slots XXX. Why this plan? As we are just beginning the scheduling process and have not reached out to X yet, we followed policy #6 (highlighted on the left) to obtain X’s availability for the slots. This will help jumpstart the process."" Guidelines: - Use each speaker’s real English name. - Focus strictly on today (ignore tomorrow or later). - Valid actions (choose strategically): 1. "Send email to X to request availability for slots XXX" (initial outreach). 2. "Follow up with X" (if email sent but no reply). 3. "Wait for response from X" (if email sent and still within expected waiting window). 4. "Confirm and assign slot for X" (only if X has shared availability; avoid finalizing too early unless time is urgent). 5. "Send clarification email to X" (if clarification on dates/details is needed). - If multiple actions are needed, combine them into one clear sentence (e.g., "Send email to X, Y, and Z to request availability for slots A, B, and C."). - Do not include any other tasks (e.g., slides, addresses, future prep). - Remember what day today is and when the last emailing was done (relevant for some policies). - Consider time urgency: - If few days remain, prioritize immediate action. - If weekend, default to waiting for responses unless a policy specifies otherwise. -------------------------- Time Context: [Time Context] Task Progress Summary: [Progress Summary] -------------------------- Applied Coordination Policies: [Applied Policy] -------------------------- Implicit Learning/User Edit History (recent, action edits only): [Implicit Learning - Action Edits] --------------------------
A.1.4. Email Drafting
I am [Organizer Name], a seminar/workshop organizer. Please help me write an email to [Speaker Name] about [Email Goal]. Please make sure it conveys that. Try to use a casual tone. If I haven’t given you any details, keep it concise and avoid unnecessary formalities. Sometimes it is a full email to invite some, while sometimes it might be just a followup email or a clarification / confirmation email. So you should write different emails according to the situation and past email history. Please return a JSON object with the enhanced email containing: - "from": organzier (just organizer, not names), - "to": recipient (ensure it’s just speaker’s Full name, no other information, dont change) - "subject": subject - "body": enhanced email body that is professional and clear Make sure the email is natural language, professional, and incorporates all context provided. Consider the time urgency when writing the email. If only a few days remain until the seminar dates, convey urgency appropriately. For follow-up emails (like following up / reminder emails -- no need to mention the seminar information.) You need to explicitly mention the dates, don’t say "next seminar/next Thursday" or "Nov 1, 8, 10". You need to list the dates individually. Start the greeting with "Dear [Speaker name]". --------------------------- Here are some information of all the speakers we want to invite Context: [Progress Summary] Past Email History with that Speaker: [Past Email History with Speaker Name] --------------------------- Time Context: [Time Context] --------------------------- Seminar Information (can be opted out if this is not the first email) [Seminar Information] --------------------------- Users’ template examples for invitation, if user has provided some, please try to follow them: User’s history invitation email : [Invitation Email Template] User’s history followup email: [Invitation Followup Template] User’s history confirmation emails : [Invitation Confirmation Template] --------------------------- When necessary, you should adapt some style edit that I have already taken on previous emails. However, be smart on whether they are the correct people to do so. [Implicit Learning - Email Edits]
A.1.5. Implicit User Preference Learning
Plan and Action Edits
You need to describe the user’s edits by comparing the AI/system {item_type} with the user-edited {item_type}, making use of the history context. Output only a brief and concise summary of the edits. Do not infer motives. Include relevant background context from the history — for example, when the change happened (if indicated) and how the text was modified. Format: ’The user deleted XXX or added XXX to the {item_type} when [context about the stage or progress of the allocation].’ Example: The user removed Tracy and Jack from the AI-generated outreach plan but kept others, such as Tanya and Adam, who are both faculty members. Inputs: history_context: {history_context}" "ai_text: {ai_text}" "user_text: {user_text}"
Email Edits
You need to describe the user’s email edits related to format and style by comparing the AI/system item_type with the user-edited item_type, making use of the history context. You need to first analyze if the user edits relate to format and style. If not, you should return "No learning needed as the user edits are not related to format and style." Note that you do not need to analyze the user edits if it relates to content (e.g. specific dates) Output only a brief and concise summary of the edits. Do not infer motives. Include relevant background context from the history — for example, when the change happened (if indicated) and how the text was modified. Format: ’The user deleted XXX or added XXX to the item_type when [context about the stage or progress of the allocation].’ Example: The user changed "Best regards," from the AI-generated outreach plan to "Thanks," in the email when initiating outreach to Jack regarding available dates. Inputs: history_context: {history_context}" "ai_text: {ai_text}" "user_text: {user_text}"
A.2. Edge Case Detection
You are an expert seminar organizer assistant. Analyze the email log, approved plans, time context, and reasoning summary. You need to flag edge-case issues that may require user confirmation or special and urgent attention or the conditions are not covered by the coordination policies (e.g., last-minute availability change, medical or accessibility information, special requests like unusual conditions not covered by the policies). If an edge-case issue is found, output a concise 40-word message starting with "FLAG:" that clearly explains the issue and ends with a direct question that requires user direction so the system can take action. Please note: If the user has already confirmed, is following the plan, or appears on the right track to solve the problem (e.g., they have already contacted speakers, have a plan in place, or are simply waiting for responses), then do not flag—output "Clear". You need to output "Clear" for routine scheduling matters (e.g., unclaimed slots, unresponsive speakers, conflicting availability, or other common slot-assignment issues). These routine tasks can usually be resolved by following up or reaching out again. In comparison, edge-case issues require explicit user guidance and explanation to ensure the situation is addressed appropriately. ### FEWSHOT EXAMPLES ### Clear examples (no special action needed): Input: Many past emails… New Email: “Thanks for the invite, I’$l$ check my schedule and get back to you in a few days.” Output: Clear Input: Many past emails… One email from another speaker: “I want to come to Sep 1! Other slots may also work!” New Email: “I can only come Sep 1.” Output: Clear Flag examples (requires organizer decision): Input: Many past emails… New Email: “I’m on maternity leave this semester, not sure I can attend, maybe someone else from my lab could take the slot.” Output: FLAG: SpeakerName seems unavailable due to maternity leave and suggesting a lab substitute. Should I follow up with the proposed colleague or reassign the slot to another invited speaker? Input: Many past emails… New Email: “I’$l$ be abroad that entire week, is there any chance I can swap to a later slot in May?” Output: FLAG: SpeakerName’s traveling during scheduled week requests swap to May. Should I reschedule them to May or retain current slot and reassign to another speaker? Input: Many past emails… New Email: “Hey! I am sorry!! I know my seminar is at the coming Friday! However, due to some family reason, I’$l$ only know the day before.” Output: FLAG: SpeakerName’s uncertain due to family reason, with final notice only one day prior. Should I secure a backup speaker now or wait for their last-minute confirmation? Here are the policies: [Policies]
A.3. Simulated Respondents
You are simulating a seminar speaker. Your persona: [Speaker Persona]. Name: [Speaker Name]. The organizer’s name is: [Organizer Name]. The current simulation day is: [Day ID]. You have received the following email thread (latest at the bottom): [Email Log] Possible dates to choose from: [Speaker Availability]] As the speaker, you can: - Reply to the organizer’s email, providing your availability, asking questions, or giving other relevant information. - If you have questions for the organizer, you can draft them in your reply and wait for a response. - You may choose to wait for more information before replying, if appropriate. - You can reply to multiple emails in the thread if you wish. - You may also choose to wait for a specific number of days before replying (e.g., "wait 2 days and reply on day 5"). If you choose to wait, specify the simulation day you intend to reply. Decide what you want to do next: - If you want to reply, write a detailed, persona-driven reply (possibly including questions for the organizer). - If you want to ask a question and wait for a response, draft your question(s) and indicate you are waiting. Do not ask questions that are too specific (’like equipment setup’ etc, this is a scheduling task, they only care if you are available on the date.) - If you want to wait and not reply yet, explain why, and specify the simulation day you intend to reply (e.g., "wait_until_day": 5). Return your action as one of: ’reply’, ’ask_question’, ’wait’. If you choose to wait, also return a "wait_until_day" field (integer simulation day to reply). Then, provide your reply email (if any) as a JSON object with keys: from, to, subject, body, timestamp. Remember, your subject should be just a short first-point-of-view summary of your reply (less than 10 words, like "I can make it on Sep 5", "I can’t make it on Sep 5", "Sorry I need a few days to think." etc). Also, provide a short "RATIONALE" section (2-4 sentences) explaining your reasoning and thought process for this turn. Format your response as: ACTION: <reply|ask_question|wait> WAIT_UNTIL_DAY: <integer or ’none’> EMAIL: <JSON object or ’none’> RATIONALE: <your reasoning> EMAIL should be like: "from": "XX", "to": "XX", "subject": "XXX", "body": "XXXXX", "timestamp": "XX" dont output json ‘‘‘ etc
A.4. Free-form QA Chat
You are an expert assistant. The following is an email log from a seminar simulation system. Answer the user’s question based on the email log. If the answer is not directly available, say so, but try to be helpful and concise. [EMAIL LOG] [USER QUESTION]
A.5. Daily Summary
Today is simulation day [Day ID]. Here is the email log for today (may include previous days): [Email Log] Please provide a concise 40-word summary of the progress made today, including any key actions taken, responses received, and remaining issues. The summary should be suitable as a notification for users in a sidebar. Dont include today’s date in the summary -- just summarize the progress.
Appendix B Technical Evaluation Details
B.1. Policy Selection
Here are the prompts used in the technical evaluations for policy selection:
B.1.1. Experiment prompts
speaker-only
You are a helpful assistant. Given the following context, select which of the following coordination policies should be applied. The chosen policies should be 100% applied to the scenario! If you found that we didn’t do anything yet from the reasoning, you should choose the ’we are just beginning’ policy. Return a JSON list of the indices of the applicable policies (e.g., [1,2] or [1]). If none apply, return an empty list []. Current context: Pending Speakers: {pending_speakers_text} Return ONLY the JSON list of indices:
slot-only
You are a helpful assistant. Given the following context, select which of the following coordination policies should be applied. The chosen policies should be 100% applied to the scenario! If you found that we didn’t do anything yet from the reasoning, you should choose the ’we are just beginning’ policy. Return a JSON list of the indices of the applicable policies (e.g., [1,2] or [1]). If none apply, return an empty list []. Current context: Slot Status: - Filled slots: {filled_slots} - Unfilled slots: {unfilled_slots} Return ONLY the JSON list of indices:
slot-only
You are a helpful assistant. Given the following context, select which of the following coordination policies should be applied. The chosen policies should be 100% applied to the scenario! If you found that we didn’t do anything yet from the reasoning, you should choose the ’we are just beginning’ policy. Return a JSON list of the indices of the applicable policies (e.g., [1,2] or [1]). If none apply, return an empty list []. Current context: Slot Status: - Filled slots: {filled_slots} - Unfilled slots: {unfilled_slots} Pending Speakers: {pending_speakers_text} Return ONLY the JSON list of indices:
summary-based
You are a helpful assistant. Given the following main reasoning result and context, select which of the following coordination policies should be applied. The chosen policies should be 100% applied to the scenario! If you found that we didn’t do anything yet from the reasoning, you should choose the ’we are just beginning’ policy. Return a JSON list of the indices of the applicable policies (e.g., [1,2] or [1]). If none apply, return an empty list []. Main reasoning result: {progress_summary} Coordination Policies: {policy_list} Return ONLY the JSON list of indices:
B.2. Edge Case Detection
B.2.1. Test suite construction
We employ GPT‑5 to generate a set of 10 edge-case (flaggable) test cases and 10 non-flagged (clear) test cases, with all examples validated for quality by the lead authors. To enhance linguistic and contextual diversity, each test case is further expanded into five stylistic variations, reflecting different tones and levels of formality. These variations are conditioned on five distinct speaker personas: a senior renowned faculty member, a junior faculty member, a senior PhD student, a junior PhD student, and a close peer researcher. In total, 100 test cases are constructed and used as our test suite for edge case detection.
B.2.2. Prompts to generate linguistic variations and test data
You are a helpful assistant. You need to generate test cases for me to test my flag detection assistant. I am organizing a seminar series, inviting speakers. You will be acting as the simulated speakers, writing response emails to my invitation email. You will be given cases where the system should mark as clear and that are not edge cases. For each case, you should generate exactly five different variations of responses in the role of the following five speaker personas: 1) a famous distinguished professor who is always busy 2) a junior faculty who is trying to establish herself in the field 3) a senior phd student who is able to graduate and is looking for a job 4) a junior phd student who is just starting his phd journey 5) a close friend of mine who is also a phd student researcher. The following are the cases where the system should mark as clear (i.e. labeled as "0"): Please generate the test cases in the following JSON format: """ { "case_description": $\{"persona":"persona\_description","response":"response","flag":"0$"\} } The case_description is the title of each scenario shown above. Please generate exactly five variations of the response for each case
B.2.3. Experiment prompts
zero-shot
You are an expert seminar organizer assistant. Analyze this response from a speaker invitation. Response: response You need to flag edge-case issues that may require user confirmation or special and urgent attention If an edge-case issue is found, simply output "FLAG", otherwise output "Clear". Your Output:
few_shot without policy
You are an expert seminar organizer assistant. Analyze the email log, approved plans, time context, and reasoning summary. You need to flag edge-case issues that may require user confirmation or special and urgent attention or the conditions are not covered by the coordination policies (e.g., last-minute availability change, medical or accessibility information, special requests like unusual conditions not covered by the policies). If an edge-case issue is found, output a concise 40-word message starting with "FLAG:" that clearly explains the issue and ends with a direct question that requires user direction so the system can take action. Please note: If the user has already confirmed, is following the plan, or appears on the right track to solve the problem (e.g., they have already contacted speakers, have a plan in place, or are simply waiting for responses), then do not flag—output "Clear". You need to output "Clear" for routine scheduling matters (e.g., unclaimed slots, unresponsive speakers, conflicting availability, or other common slot-assignment issues). These routine tasks can usually be resolved by following up or reaching out again. In comparison, edge-case issues require explicit user guidance and explanation to ensure the situation is addressed appropriately. ### FEWSHOT EXAMPLES ### Clear examples (no special action needed): Input: Many past emails… New Email: “Thanks for the invite, I’$l$ check my schedule and get back to you in a few days.” Output: Clear Input: Many past emails… One email from another speaker: “I want to come to Sep 1! Other slots may also work!” New Email: “I can only come Sep 1.” Output: Clear Flag examples (requires organizer decision): Input: Many past emails… New Email: “I’m on maternity leave this semester, not sure I can attend, maybe someone else from my lab could take the slot.” Output: FLAG Input: Many past emails… New Email: “I’$l$ be abroad that entire week, is there any chance I can swap to a later slot in May?” Output: FLAG Input: Many past emails… New Email: “Hey! I am sorry!! I know my seminar is at the coming Friday! However, due to some family reason, I’$l$ only know the day before.” Output: FLAG Response: {response} Your Output:
few_shot with policy
You are an expert seminar organizer assistant. Analyze the email log, approved plans, time context, and reasoning summary. You need to flag edge-case issues that may require user confirmation or special and urgent attention or the conditions are not covered by the coordination policies (e.g., last-minute availability change, medical or accessibility information, special requests like unusual conditions not covered by the policies). If an edge-case issue is found, output a concise 40-word message starting with "FLAG:" that clearly explains the issue and ends with a direct question that requires user direction so the system can take action. Please note: If the user has already confirmed, is following the plan, or appears on the right track to solve the problem (e.g., they have already contacted speakers, have a plan in place, or are simply waiting for responses), then do not flag—output "Clear". You need to output "Clear" for routine scheduling matters (e.g., unclaimed slots, unresponsive speakers, conflicting availability, or other common slot-assignment issues). These routine tasks can usually be resolved by following up or reaching out again. In comparison, edge-case issues require explicit user guidance and explanation to ensure the situation is addressed appropriately. ### FEWSHOT EXAMPLES ### Clear examples (no special action needed): Input: Many past emails… New Email: “Thanks for the invite, I’$l$ check my schedule and get back to you in a few days.” Output: Clear Input: Many past emails… One email from another speaker: “I want to come to Sep 1! Other slots may also work!” New Email: “I can only come Sep 1.” Output: Clear Flag examples (requires organizer decision): Input: Many past emails… New Email: “I’m on maternity leave this semester, not sure I can attend, maybe someone else from my lab could take the slot.” Output: FLAG Input: Many past emails… New Email: “I’$l$ be abroad that entire week, is there any chance I can swap to a later slot in May?” Output: FLAG Input: Many past emails… New Email: “Hey! I am sorry!! I know my seminar is at the coming Friday! However, due to some family reason, I’$l$ only know the day before.” Output: FLAG Here are the policies: - "Don’t be shy to ask for more availability if they don’t give you many. They will probably reply more if you beg." - "Some people don’t read or respond to emails right away because they’re busy. You should wait for their response, but it’s okay to follow up every two workday days if you haven’t heard back.", - "If someone hasn’t responded to the email after three reach-outs, then don’t contact them again.", - "Ambiguity is okay. People can typically do the days they are vague about — you can prioritize asking for more availability and just confirm the ambiguity right before assigning them at the end.", - "Some speakers—especially faculty—are more desirable to schedule. We should prioritize their availability by assigning slots to faculty over students and asking students to be more flexible or to provide additional dates to accommodate them.", - "If we are just beginning and have not yet contacted any speakers for availability, then ask each speaker for their availability across all slots." Response: response Your Output:

### v2

@@
-DoubleAgents: Exploring Mechanisms of Building Trust with Proactive AI
+DoubleAgents: Interactive Simulations for Alignment in Agentic AI
 Abstract.
-Agentic workflows promise efficiency, but adoption hinges on whether people actually trust systems that act on their behalf.
+Agentic workflows promise efficiency, but adoption hinges on whether people can align systems that act on their behalf with their goals, values, and situational expectations.
 We present DoubleAgents, an agentic planning tool that embeds transparency and control through user intervention, value-reflecting policies, rich state visualizations, and uncertainty flagging for human coordination tasks.
-A built-in respondent simulation generates realistic scenarios, allowing users to rehearse, refine policies, and calibrate their reliance before live use.
-We evaluate DoubleAgents in a two-day lab study (n=10), two deployments (n=2), and a technical evaluation.
-Results show that participants initially hesitated to delegate but grew more reliant as they experienced transparency, control, and adaptive learning during simulated cases.
-Deployment results demonstrate DoubleAgents’ real-world relevance and usefulness, showing that the effort required scaled appropriately with task complexity and contextual data.
-We contribute trust-by-design patterns and mechanisms for proactive AI—consistency, controllability, and explainability—along with simulation as a safe path to build and calibrate trust over time.
+A built-in respondent simulation generates realistic scenarios, allowing users to rehearse, refine policies, and calibrate their use of agentic behavior before live deployment.
+We evaluate DoubleAgents in a two-day lab study (n=10), three deployment studies, and a technical evaluation.
+Results show that participants initially hesitated to delegate but used simulation to probe system behavior and adjust policies, gradually increasing delegation as agent actions became better aligned with their intentions and context.
+Deployment results demonstrate DoubleAgents’ real-world relevance and usefulness, showing that simulation helps users effectively manage real-world tasks with higher complexity and uncertainty.
+We contribute interactive simulation as a practical pathway for users to iteratively align and calibrate agentic systems.
 generative AI, AI agents, trust, reliance, transparency, state tracking, seminar organizing, control, planning, human coordination, reasoning, constitutional AI
-copyright: acmcopyrightjournalyear: 2025conference: ; ; booktitle: Proceedings of the 2025price: 15.00isbn: 978-1-4503-XXXX-X/18/06ccs: Human-centered computing Interactive systems and tools
+copyright: acmcopyrightjournalyear: 2025conference: 2026; 2026; arxivbooktitle: Proceedings of the 2026price: 15.00isbn: 978-1-4503-XXXX-X/18/06ccs: Human-centered computing Interactive systems and tools
+Figure 1.
+DoubleAgents is an agentic AI coordination tool that uses interactive simulation to let users experience and align the full ReAct cycle before live deployment.
+Starting from an initial setup (left)—including a simulation goal, simulated test-case personas, and an initial set of policies, and email templates—users can undergo multiple simulated ReAct cycles (middle) to rehearse planning, review and approve actions, execute plans, detect new edge cases, and iteratively refine policies via stop hooks, and email templates over repeated test cases and time.
+This process results in (right) the final speaker-slot allocation, as well as a user-aligned set of policies, emails, and stop hooks before the system sends real emails.
 1. Introduction
-Coordinating human effort is essential for tasks at every scale.
-Event planning involves finding suitable times, locations, vendors, and transportation.
-Warehouses must assign workers with the right skills and experience to shifts they are available for.
-Even organizing a speaker series requires emailing potential speakers, collecting their availability, and assigning them to appropriate slots.
-These problems involve both discovering people’s availability and solving the constraints of assigning individuals to limited roles.
-Human coordinators face time-consuming and intricate communication challenges: people may hesitate to disclose their full availability, delay responses, or provide only tentative answers such as “I might be able to come Saturday, but I’m not sure yet.” Coordinators must also balance group dynamics, make exceptions for important individuals, and continuously replan when last-minute changes arise.
-Coordination is thus both tedious and delicate, requiring continuous monitoring and re-planning while respecting social norms and relationships.
-Generative AI now possesses several key capabilities that could help with coordination, including deep reasoning to evaluate constraints and produce plans (Guo et al., 2025; Yao et al., 2023b, a; Stechly et al., 2025), and API tool-calling to take concrete actions (Trivedi et al., 2024; Yao et al., 2024; Masterman et al., 2024).
-However, people do not yet trust AI to act proactively on their behalf for complex and delicate tasks like human coordination problems.
-AI systems are prone to hallucinations, often lack sufficient context to understand the problem (Zhang et al., 2025; Athaluri et al., 2023), and may mishandle delicate social interactions (Zhou et al., 2023; Wang et al., 2024).
-Most importantly, they lack an awareness of when to stop and escalate to human oversight.
-As a result, current AI interfaces largely wait for users to issue prompts and micromanage responses.
-This limits productivity and prevents AI from relieving people of burdensome procedural work.
-To unlock proactive AI, we need frameworks that let humans observe, guide, and calibrate trust in AI over time.
-Trust is not established by statements alone—it develops through repeated reliable behavior over time (Razin and Feigh, 2024; Fouragnan et al., 2013).
-Game theory suggests that trust between humans is built by repeated observation of behavior in varied circumstances (Kollock, 1998).
-Additionally, in the AI trust and safety literature, Constitutional AI (Bai et al., 2022) is a paradigm that uses human-articulated policies (called a constitution) that can align an AI’s behavior with human expectations and values.
-Policies in making expectations explicit - so that humans and AI have a shared understanding of the expected behavior.
-Repeated interactions and clear policies are two pillars of establishing trust that we build upon to help build trust in proactive AI.
 Figure 2.
@@
 For a detailed walkthrough, please refer to Section 4.
-[…]
 Diagram showing the architecture of the DoubleAgents system, with components for goal setting, agent coordination, context management (including edge case detection), and various submodules such as toolbox, orchestrator, and simulated respondents.
-We propose to build trust in proactive AI through simulation and policy refinement.
-Human behavior can be unpredictable, but AI-driven simulations can generate a wide range of realistic scenarios (Park et al., 2023), including late, incomplete, or uncertain replies.
-These allow coordinators to test and refine policies in safe environments.
-Policies specify how the AI should act (e.g., when to follow up, when to stop, how to phrase outreach) and when to escalate to human oversight.
-Hence, we propose to use simulation to expose AI behavior across repeated interactions, letting users correct and evolve policies iteratively.
-Visualizations of system state—goal progress, a timeline view, and person-level interaction histories—provide transparency and help humans intervene efficiently when needed.
-We present DoubleAgents, shown in Figure 1, an interactive agentic system designed to automate aspects of human coordination while supporting user oversight and trust-building.
-DoubleAgents integrates generative AI agents in two complementary roles: (1) a ReAct Loop (Yao et al., 2023b) with AI planning and execution agents that can call tools to get appropriate context, query the current state of the system, and execute actions such as emailing potential speakers.
-This loop is extended to provide interaction for human interactivity and oversight at every step.
-(2) A LLM-based simulation of respondents that generate the varied and realistic human responses for various respondent personas (e.g.
-a professor who is very slow to respond, or a speaker who won’t commit).
-The LLM simulations allow users to experience repeated execution of the system to understand how it works, evolve the policies over time to suit their preferences, and potentially build trust that the system will act according to their wishes.
-By combining policy generation and refinement, action execution, and multi-agent LLM simulation within a user-centered feedback process, DoubleAgents helps users delegate repetitive aspects of coordination while retaining control of high-stakes interactions.
-In a lab study (n=10) of people using the system with simulated responses over two study sessions, participants began hesitant to delegate but, through visible state, explicit reasoning based from policies, strong intervention points, and rehearsal on realistic cases, grew more comfortable relying on the agent for both planning and execution.
-In two real-world deployments of the system, we found that organizers adopted DoubleAgents as transparency, policy controls, and edge‑case flagging built trust; however, required more upfront context setup and tighter email integration.
-Our work demonstrates how agentic systems can support reliable, human-centered coordination by embedding simulation, transparency, and interactive oversight into the agentic workflow.
+Human–AI alignment is one of the defining challenges of the 21st century, especially for agentic AI: systems that go beyond answering questions to taking actions on users’ behalf (Leike et al., 2018; Soares and Fallenstein, 2014; Hendrycks et al., 2021).
+Following the Bidirectional Human–AI Alignment framework defined by Shen et al.
+(2025), we view alignment in agentic settings as an ongoing co-adaptation process between users and systems.
+On the system side, agents must be continuously shaped through interaction history and curated context so their actions reflect an individual user’s goals, preferences, and social norms (Zhang et al., 2025; Gao et al., 2024).
+On the human side, users must continually develop appropriate mental models, usage patterns, and AI literacy so they can effectively direct, evaluate, and calibrate delegation (Vanderlyn et al., 2025; Andrews et al., 2023).
+As agentic AI becomes more common, we need practical ways to assess and improve whether these systems remain aligned as they act in the world.
+Today, many popular agentic use cases, such as agentic coding or “vibe coding”, are relatively low-stakes and operate in closed sandbox environments (Horikawa et al., 2025).
+In these domains, alignment is often easier because systems can leverage large-scale, well-structured public data (e.g., GitHub), and success criteria are comparatively legible (e.g., whether code compiles, passes tests, or produces the desired output) (Chen et al., 2021; Jimenez et al., 2024).
+However, in more complex tasks that must adhere to human and domain knowledge as well as social cognition and norms, such as writing emails to real people to request favors, where etiquette is highly context-dependent, agents cannot reliably align on their own (Afroogh et al., 2024; Gabriel, 2020).
+These tasks often require nuanced, case-by-case judgment that is difficult to infer from generic training data alone and may demand substantial human oversight if deployed naively.
+One representative scenario is human coordination, such as scheduling workers with the right skills and experience for shifts they are available for or inviting researchers to a speaker series.
+In this work, we focus on the speaker-series scheduling setting.
+Organizers must secure a speaker for each slot (e.g., 4-5pm on Fridays), coordinate availability across multiple candidates, and handle uneven responsiveness (e.g., some invitees reply late or not at all).
+Responses can also introduce socially sensitive and context-dependent requests: “Can you reimburse travel?”, “Can I do Zoom?”, “I can’t make it on Fridays, what about Thursdays 4-5pm?” “Sorry for the last-minute cancellation—I now have a conflict,” or “My schedule is uncertain; can you follow up next month?” Determining the right next step is rarely optimal in a mathematical sense; it depends on the organizer’s values and preferences, which we refer to as “policies” (e.g., “Offer Zoom only if in-person is not feasible and no one else can do that slot,” “Prioritize senior faculty over PhD students when conflicts happen,” or “Follow up after two business days unless urgent”).
+In practice, because these “actions” are carried out almost entirely through email communication, the organizer must decide which email to send next (and to whom), how to phrase each email with the right content and right tone, and when to send it.
+While an LLM-based agent may generate plausible actions and write acceptable emails, it often lacks the organizer’s specific norms and cannot reliably anticipate how preferences should change across cases.
+Alignment therefore must happen on both sides: the system needs ways to acquire and apply three user-specific automation refinements: (1) coordination policies that guide planning and action choices, (2) email templates and writing style that match the user’s voice and social intent for each action, and (3) stop hooks that define when the agents need to pause and request for user confirmation; and the user needs support to form a clear mental model and specify these three preferences, especially early on, when they may not yet know what rules, tone, or escalation boundaries they want.
+The central challenge lies in how can we accelerate this bidirectional alignment so the agent can act more autonomously over time, without requiring constant babysitting.
+Our core approach is interactive simulation to support iterative automation refinement, enabling alignment between users and agents.
+As modern AI can simulate diverse human behaviors (Park et al., 2023, 2022), we use simulation as a testbed to expose users to realistic coordination scenarios, including problematic or edge-case behaviors, and to help them define and refine their communication strategies.
+During simulations, the agent proposes plans and actions (e.g., inviting multiple candidates, deciding when to wait versus follow up, or holding a slot for a high-priority speaker), and the user evaluates whether these choices are socially and contextually appropriate.
+This mirrors ideas from software engineering such as unit and integration testing: a system earns confidence not by succeeding on a single happy path, but by handling a range of representative cases.
+In our context, however, completing a task is not enough: agentic AI systems must be user-specific and value-aligned, reflecting personal priorities, norms, and preferences.
+We present DoubleAgents, a system for aligning agentic AI through interactive simulation.
+DoubleAgents integrates generative AI to simulate two complementary roles, as Figure 2 shows.
+First, it uses a ReAct loop (Yao et al., 2023) that pairs coordination agents for planning and execution.
+The agents can call tools to retrieve relevant context, query the current task state, and execute actions such as drafting and sending emails to potential speakers.
+We extend the loop with explicit interaction checkpoints so users can review, intervene, and approve decisions throughout planning and execution.
+Second, DoubleAgents includes an LLM-based simulator of respondents that generates diverse, realistic replies for different personas (e.g., a professor who responds slowly or a respondent who introduces funding or logistics constraints).
+This simulation enables repeated run-throughs of difficult cases, helping users understand how the agent behaves, stress-test policies against edge cases, and iteratively refine preferences and autonomy boundaries.
+Across interactive simulation runs, the primary outputs are not only a candidate schedule (speaker–slot allocation), but also a reusable set of user-specific alignment artifacts: refined policies, updated email templates, and calibrated stop hooks that can be applied when the system is deployed to send real emails.
+We evaluate DoubleAgents through a technical evaluation, a two-day lab study (n=10) in which participants used DoubleAgents with simulated respondents, and three real-world deployments.
+In the lab study, participants initially hesitated to delegate.
+However, over time, through multiple simulation rounds and the two study sessions, participants increasingly delegated, as visible state, policy-grounded reasoning, and clear intervention points (including stop-hook flagging) helped them calibrate autonomy and build confidence in the agent’s planning and execution.
+By the end, participants also produced reusable policies, stop hooks, and email templates aligned with their voice and style.
+Three week-long deployment studies demonstrate DoubleAgents’ usefulness and relevance for real-world, complex human coordination tasks.
+DoubleAgents was connected to organizers’ real email workflows and sending emails to real invitees, as well as incorporated incoming real speakers’ replies into its ongoing coordination state.
+Organizers reported that interactive simulation beforehand helped them prepare for deployment by crafting strong initial policies and email templates for real-world use.
+Our work demonstrates how agentic systems can support reliable, human-centered coordination by embedding simulation, transparency, and interactive oversight into the agentic AI workflow.
 Specifically, we contribute:
-  • A policy-driven, intervention‑friendly approach to govern proactive AI behavior in an agentic and integactive ReAct framework to build trust with users.
-  • An AI-driven simulation framework of human responses to support repeated observations of reliable behavior over time to test and tune policies for AI to follow.
-  • A user study (n=10) and a deployment with two users to evaluate how users perceive and gradually trust proactive AI for a human coordination task of scheduling a speaker seminar.
+  • A design perspective on alignment in agentic AI, realized through a policy-driven, intervention-friendly ReAct framework that supports transparency, human oversight, and alignment with human goals and values.
+  • An interactive simulation mechanism that models realistic human responses, enabling users to safely explore agent behavior, understand system dynamics, rehearse interventions, and iteratively refine policies, email references, and stop hooks before real-world application.
+  • Empirical results from a user study (n=10) and three real-world deployments, examining how interactive simulation and policy refinement support alignment in agentic AI for human coordination.
 2. Related Works
+2.1. Alignment in Human-AI Systems
+As AI agents take on more complex tasks, ensuring they align with human values has become a central research challenge.
+A key strategy is reinforcement learning from human feedback (RLHF), which fine-tunes models to produce outputs consistent with user expectations (Ouyang et al., 2022).
+However, alignment doesn’t end at training.
+Studies show that users often push back on AI responses that conflict with their values, steering behavior through real-time feedback, a process termed user-driven value alignment (Fan et al., 2025).
+Such interactive alignment is critical in high-stakes contexts like medical decision support, where LLM-powered co-pilot systems have helped clinicians catch more errors than either humans or AI alone (Ong et al., 2025).
+These findings show the importance of transparency, oversight, and human agency in AI deployment.
+Another approach, Constitutional AI, replaces constant human feedback with rule-based self-supervision.
+Models are trained to revise outputs using explicit principles like “Be respectful” (Bai et al., 2022).
+Recent work has expanded this into participatory frameworks such as Collective Constitutional AI, where large groups co-author alignment rules to reflect diverse values (Huang et al., 2024).
+Tools like ConstitutionMaker further help users turn natural language feedback into usable policies (Petridis et al., 2024).
+DoubleAgents contributes to this space by embedding alignment into interactive simulation.
+Users iteratively refine coordination policies, email tone, and stop hooks through rehearsed ReAct loops, treating alignment not as a one-time step, but as a continuous, participatory process.
 2.2. Agentic AI in Task Planning and Execution
-“AI Agents” typically refers to AI that can perform tasks for people.
-This can be done through tool use (namely, calling an API) (Schick et al., 2023), or operating a computer (Anthropic, 2025; OpenAI, 2025; Manus AI, 2025; Labs, 2025).
-Although AI Agents show impressive potential to act, trusting agents to act on behalf of a human is a critical missing piece (Afroogh et al., 2024).
-The delegation of control to AI raises questions not only about competence, but also about alignment, transparency, and user oversight.
-Planning and execution remain central yet unresolved challenges in the development of AI agents.
-A growing body of work highlights both the difficulty and the importance of task planning, as seen in benchmarks and simulation environments such as Tau-Bench (Yao et al., 2024), NaturalPlan (Zheng et al., 2024), TravelPlanner (Xie et al., 2024), and AppWorld (Trivedi et al., 2024).
-These studies underscore that robust, goal-directed planning, particularly in open-ended or multi-step domains, is not only technically demanding but also essential for delivering meaningful user value.
-To address this challenge, recent work has explored how LLMs can scaffold user planning by decomposing high-level goals into actionable steps (Zhang et al., 2024).
-Techniques such as ReAct (Yao et al., 2023b), which interleave reasoning and acting, have gained traction in multi-step environments for improving agent adaptability and responsiveness.
-These approaches are often embedded within multi-agent systems to support domains like travel, scheduling, and personal task management (Lal et al., 2023).
-However, autonomous systems such as AutoGPT (Significant Gravitas, [n.
-d.]) often fall short in transparency, adaptability, and robustness, limiting their usability in real-world settings (Zheng et al., 2024; Yao et al., 2024).
-Furthermore, prior cognitive science research shows that humans rarely engage in exhaustive upfront planning due to cognitive load (Krieger et al., 2009; Teevan et al., 2016), highlighting the need for AI systems that can support incremental, iterative, and contextualized planning over time.
-Additionally, recent studies suggest that human–AI co-planning, particularly when agents expose their intermediate reasoning and allow users to guide or intervene, can improve both task performance and user trust (He et al., 2025).
-When users can monitor and influence agent decisions at each step, they are better able to understand, calibrate, and ultimately trust the system, especially in domains requiring high-stakes delegation or multi-party coordination.
+AI agents, systems that perform tasks on users’ behalf, operate through tool use (e.g., APIs) (Schick et al., 2023) or direct interface manipulation (Anthropic, 2025; OpenAI, 2025; Manus AI, 2025; Labs, 2025).
+While these agents show growing capabilities, delegating control to them still raises concerns around trust, transparency, and alignment (Afroogh et al., 2024).
+Planning and execution remain core challenges.
+Benchmarks like Tau-Bench, NaturalPlan, TravelPlanner, and AppWorld (Yao et al., 2024; Zheng et al., 2024; Xie et al., 2024; Trivedi et al., 2024) highlight the difficulty of robust multi-step planning, especially in open-ended tasks.
+To address this, methods like task decomposition (Zhang et al., 2024) and ReAct, which interleaves reasoning and acting, have been shown to improve agent adaptability (Yao et al., 2023), especially when used in collaborative or multi-agent domains (Lal et al., 2023).
+However, fully autonomous agents like AutoGPT often struggle with robustness and transparency in real-world use (Zheng et al., 2024; Yao et al., 2024).
+Cognitive research suggests humans rarely plan exhaustively upfront (Krieger et al., 2009; Teevan et al., 2016), pointing to a need for agents that support incremental and context-sensitive planning.
+Recent studies show that exposing agent reasoning and allowing user intervention improves both performance and trust (He et al., 2025).
+In sum, planning in agentic AI is not just a technical hurdle; it is a design problem.
+Transparent reasoning, iterative refinement, and human-in-the-loop oversight are essential for building trustworthy agents that can meaningfully support real-world tasks.
 2.3. Human Simulations with LLMs
-Recent work has explored the use of LLMs in multi-agent simulations  (Park et al., 2023), revealing their surprising capacity to model (Horton, 2023), predict, and even exhibit human-like behavior - including taking actions they were not explicitly instructed to perform in order to achieve high-level goals such as forming routines, organizing events, and collaborating spontaneously (Park et al., 2022, 2023).
-These behaviors often mirror human reasoning and coordination strategies.
-Other studies have shown that these behaviors closely align with empirical observations of humans across various domains, including economic games (Sreedhar et al., 2025; Sreedhar and Chilton, 2025; Aher et al., 2023; Horton, 2023), U.S.
-Supreme Court judgments (Hamilton, 2023), and over 80% of published and unpublished psychology experiments (Hewitt et al., 2024).
-These results suggest that LLM agents can serve as realistic proxies for humans in testing coordination, negotiation, and reasoning behaviors at scale.
-AI systems that silate human behavior have already shown a striking capability to engage in complex, social behaviors competitively well with humans, for real world purposes.
-.
-Beyond simulation, AI systems like CICERO (Team et al., 2022) have shown how language agents can engage directly with humans in dynamic, multi-party settings.
-In online Diplomacy games, CICERO achieved a top 10% ranking among human players - illustrating the potential for LLM-based systems to engage in trust-sensitive planning, social reasoning, and strategic communication.
-The MATRIX framework (Pang et al., 2024) improved alignment in constitutional AI evaluations by useing simulated AI dialogues to anticipate a range of perspectives and surface important value trade-offs.
-Many systems have successfully used LLM red teaming  (Purpura et al., 2025) to identify potential threats in systems.
-Although AI simulations might not perfectly replicate individual user behaviors (yet) - they can be used to explore edge cases, probe systems, and engage in sensitive dialogs with humans.
-2.3. Constitutional AI
-Constitutional AI is a framework for aligning AI systems with human values by training models to self-critique and revise their outputs based on a set of guiding principles, or a ”constitution”(Bai et al., 2022).
-Unlike reinforcement learning with human feedback (RLHF)(Ouyang et al., 2022), which requires human evaluation for each model response, Constitutional AI allows the model to evaluate its own behavior using pre-defined rules.
-These constitutions are typically expressed as human-readable policies such as “Don’t promote illegal activity” or “Be helpful and respectful,” and can be authored by anyone.
-Building on this idea, Collective Constitutional AI (Huang et al., 2024) introduced a participatory approach in which over 1,000 people contributed rules to a shared constitution.
-Models trained on this crowd-sourced constitution demonstrated qualitatively different behaviors compared to those trained on developer-authored rules—highlighting the diversity of human values and community norms, and the implications this has for model alignment.
-To create constitutions, interactive systems like ConstitutionMaker (Petridis et al., 2024) have shown that LLMs can help transform users’ natural language feedback into generalized alignment policies.
-By providing real-time AI assistance in policy creation, such systems help users articulate their values more effectively, reduce cognitive burden, and make the process of shaping AI behavior more accessible and inclusive.
-These advances open up opportunities for systems that incorporate user-driven policy creation and alignment into workflows that have real-world impacts.
-3. Seminar Planning: A Human Coordination Problem
-Human coordination—such as assigning work shifts or scheduling job interviews—is both cognitively demanding and socially delicate.
-Unlike tasks with stable inputs and predictable outcomes, coordination requires managing uncertainty, reconciling multiple perspectives, and sustaining relationships over time.
-Even small lapses—missed follow-ups, unclear tone, or misaligned expectations—can erode trust and derail collective progress.
-These challenges make coordination both essential and persistently difficult in many areas of knowledge work.
-We study these challenges through one concrete and consequential setting: organizing an academic seminar series.
-A typical series spans an entire semester and involves inviting a new external speaker nearly every week, amounting to 8–12 speakers per term.
-Early on, organizers must identify potential speakers, gather their availability constraints, and balance their own, such as ensuring topical diversity across weeks and securing highly sought-after senior researchers who are harder to schedule.
-However, the most unpredictable and challenging part is dealing with responses.
-Some people do not reply at all and require multiple follow-ups to elicit a response.
-Some give only vague and tentative replies that need repeated confirmation.
-Some can be flakey and unreliable, offering slots but later canceling.
-And many unexpected problem cases arise: people might be on parental leave, ask for visa letters or funding support, recommend a different speaker, request a different date or talk modality, or raise logistical demands.
-In such human coordination tasks, you can rarely predict what is coming.
-Therefore, conventional scheduling tools such as Calendly [^1] https://calendly.com/ or When2Meet [^2] https://www.when2meet.com/ are ill-suited to this setting.
-They assume stable constraints and one-shot coordination, yet seminar scheduling requires ongoing negotiation, persuasion, and relationship maintenance.
-Unlike ordinary meeting scheduling, seminar speakers are doing you a favor, which means you cannot simply send them a link to pick a time—you must persuade and negotiate.
-Social dynamics and etiquette also play a large role: organizers must time their emails carefully to avoid overwhelming, annoying, or losing momentum with potential speakers, who are often busy scholars.
-Knowing when to nudge, follow up, or wait is subtle.
-Email remains the standard medium for professional and academic outreach, allowing organizers to adapt tone—formal, casual, or enthusiastic—to different speakers in ways scheduling links cannot.
-It also supports sharing rich context, such as talk topics, background, past speakers, or travel plans—far beyond simply choosing a calendar slot.
-Thus, experienced organizers adopt personal strategies to cope with such a demanding task.
-For example, some follow a rule of “wait two days, then follow up” to nudge busy people who may not read emails right away, while others wait a full week before reaching out again.
-Some prefer contacting close collaborators or junior researchers first, while others prioritize securing prominent professors or distant speakers early to reduce travel constraints.
-When conflicts arise between speakers, organizers might “give faculty priority and ask students to be more flexible,” carefully deciding whom to approach for alternative availability to minimize disruption.
-For non-responders or last-minute cancellations, they must judge when and how often to follow up before moving on to new options—and sometimes even decide when it is appropriate to “beg” busy invitees for additional availability.
-Because plans are constantly evolving, organizers become locked in a cognitively and time-intensive loop of tracking shifting states, replanning, and executing revised plans as new information arrives—yet they often cannot proceed with the next steps or finalize allocations until replies are received.
-Since seminar scheduling requires nuanced, context-dependent decisions that reflect each organizer’s values and preferences, it is difficult to fully automate.
-An effective seminar scheduling tool should not attempt to handle the process independently; instead, it must learn from and adapt to the organizer, respecting social norms, relational dynamics, and the inherent unpredictability of responses.
+Recent work has shown that LLMs in multi-agent simulations can exhibit surprisingly human-like behaviors, forming routines, organizing events, and coordinating spontaneously without explicit instructions (Park et al., 2023; Horton, 2023; Park et al., 2022).
+These behaviors often reflect human reasoning and social strategies, aligning with patterns observed in economic games (Sreedhar et al., 2025; Sreedhar and Chilton, 2025; Aher et al., 2023), judicial decisions (Hamilton, 2023), and over 80% of psychology experiments (Hewitt et al., 2024).
+This suggests LLM agents can serve as useful proxies for modeling human behavior at scale.
+Beyond simulation, LLMs are demonstrating real-time social competence.
+CICERO, a dialogue-enabled Diplomacy agent, ranked in the top 10% of human players by integrating language with strategic planning (Team et al., 2022).
+Similarly, the MATRIX framework used simulated AI dialogues to surface moral tradeoffs in Constitutional AI evaluations (Pang et al., 2024), showing how LLMs can assist in value alignment.
+LLM simulations also support red teaming and adversarial testing, enabling safe, scalable evaluation of edge cases and vulnerabilities (Purpura et al., 2025).
+While imperfect, these agents provide a reproducible and low-risk method for probing system behavior in socially sensitive contexts, offering a powerful complement to human-centered alignment and evaluation.
 3. DoubleAgents System
@@
 (F) Calendar View – Provides temporal context by showing the current simulation and upcoming seminar dates, color-coded for clarity.
-We present DoubleAgents, an interactive agentic system designed to investigate how users build and calibrate trust in proactive AI systems.
-Built on the ReAct framework (Yao et al., 2023b), DoubleAgents enables iterative planning and execution through a team of AI agents that assist users in complex coordination tasks.
-In this work, we focus on the use case of assigning seminar speakers to available time slots.
-The system takes as input a user-defined goal (e.g., “Organize a seminar with four speakers and four slots.”) and produces as output a finalized assignment of speakers to slots.
-DoubleAgents consists of six core components, grouped into the following three functional categories.
-First, it contains the Core ReAct Agents that carry out the reasoning, planning, and simulation logic that drive the coordination process:
+We present DoubleAgents, a system for aligning agentic AI through interactive simulation.
+Built on the ReAct framework (Yao et al., 2023), DoubleAgents supports iterative planning and execution via a team of AI agents that collaborate with users on complex coordination tasks.
+In this paper, we focus on a concrete use case: assigning seminar speakers to available time slots (see Appendix A for additional background).
+DoubleAgents takes as input a user-defined goal (e.g., “Organize a seminar with four speakers and four slots”), persona descriptions for simulated respondent agents, the user’s existing email templates, and a set of generic scheduling policies.
+Through iterative simulations, the system produces a finalized assignment of speakers to slots, along with refined policies, updated email templates, and executable stop hooks that users can later deploy in real-world agentic workflows.
+The system comprises six core components organized into three functional categories.
+First, DoubleAgents includes the Core ReAct Agents, which implement the reasoning, planning, and simulation logic that drives the coordination process:
   • Coordination Agent - Responsible for reasoning over the current problem state using a set of predefined policies. It proposes plans and actions, and invokes appropriate tools to execute them.
@@
   • Edge Case Detector - Identifies situations that fall outside the coverage of existing policies and escalates them to the user for clarification and additional instructions.
+These escalations are recorded as stop hooks, enabling iterative refinement and extension of the policy set.
 Lastly, the User Interaction Interfaces components enable transparency, user control, and real-time feedback during the coordination process:
   • Interactive Chat Interface - Facilitates communication between the user and the system.
-Users can confirm or revise generated plans, actions, and emails, and respond to flagged edge cases by providing natural language feedback.
+Users can approve or revise generated plans, actions, and emails, and respond to flagged edge cases by providing natural language feedback.
   • Visualization Dashboard - Provides real-time feedback on the state of the task, including temporal progress, assignment status, and an activity tracker to help users monitor the overall coordination process.
@@
 B reply in the last email?” to retrieve the relevant message history directly within the chat.
-This chat functionality is powered by GPT-4o, with prompt details provided in the Appendix A.4.
+This chat functionality is powered by GPT-4o, with prompt details provided in the Appendix C.4.
+(a) The pop-up window for a generated plan.
+(b) The pop-up window for generated actions following the plan.
+(c) The pop-up window for a generated email.
+Figure 5. Examples for plan and action generation.
 3.3. Coordination Agent
@@
 Based on this information, the coordination agent generates a summary of the current status, indicating which days have been scheduled, what actions have been taken, and the most recent contact date for each speaker to help monitor responses and follow-ups.
-The prompt used in generating the summary is illustrated in the Appendix A.1.1.
+The prompt used in generating the summary is illustrated in the Appendix C.1.1.
 3.3.2. Policy Selection
-To design a set of general policies applicable to a broad range of planning problems, five experts were recruited to brainstorm and define a framework of rules tailored to address common scheduling challenges.
+To design a set of general policies applicable to a broad range of planning problems, the lead authors brainstorm and define rules tailored to address common scheduling challenges.
 These challenges include: (1) individuals being reluctant to share their full availability, (2) ambiguous or unclear responses, (3) the need for multiple follow-ups to elicit responses, (4) resolving scheduling conflicts between two speakers, and (5) initiating the process by broadly reaching out to potential participants.
@@
 The policy acts as the planning guideline, informing the agent’s next steps.
-Details of the prompts used for the policy selection module are provided in the Appendix A.1.2.
-(a) The pop-up window for a generated plan.
-(b) The pop-up window for generated actions following the plan.
-(c) The pop-up window for a generated email.
-Figure 4. Examples for plan and action generation.
+Details of the prompts used for the policy selection module are provided in the Appendix C.1.2.
 3.3.3. Plan Generation
@@
 In general, there are three types of actions: (1) drafting emails to selected respondents, (2) sending emails to specified recipients using their provided addresses, and (3) awaiting responses from certain respondents.
-The detailed prompts used for plan and action generation are provided in the Appendix A.1.3.
+The detailed prompts used for plan and action generation are provided in the Appendix C.1.3.
 3.3.5. Action Execution
@@
 To streamline the summarization process, we leverage GPT-4o.
-Details of the prompt used for generating summaries of user edits across various content types can be found in Appendix A.1.5.
+Details of the prompt used for generating summaries of user edits across various content types can be found in Appendix C.1.5.
 The user-initiated edits, the corresponding system-generated drafts, and the generated summaries of the edits are all preserved in the user edit history.
@@
 Given that Prof.B’s persona also specifies preferences, such as “prefers presenting over Zoom”, the generated action will explicitly request a Zoom presentation option.
-The prompt for speaker agent, both for generating actions and drafting email responses, is detailed in the Appendix A.3.
+The prompt for speaker agent, both for generating actions and drafting email responses, is detailed in the Appendix C.3.
 To execute the selected actions, the simulated respondent agents compose and send email responses, which are then received by Lee on his end.
@@
 B.
+Lee’s clarification is recorded as a stop hook and used to refine the policy set, explicitly specifying that Zoom is not preferred when speakers request remote participation.
+In this way, user interventions during edge cases are converted into reusable system policies.
 In general, the edge case detector acts as a tool for the coordination agent to check if the message sent from the respondent agent can be addressed with the existing set of policies in order to make the next plans.
 When the detector identifies an edge case, the coordination agent escalates the issue to the user by presenting a clarification question in the interactive chat interface, prompting the user to provide guidance on how to proceed.
-The edge case detector is implemented with GPT-4o (see prompts in the Appendix A.2).
+The edge case detector is implemented with GPT-4o (see prompts in the Appendix C.2).
 3.6. Dashboard Visualization
-To support user understanding and foster trust, DoubleAgents provides clear visual feedback on the task state and system progress throughout the scheduling process.
+To support user understanding and foster alignment, DoubleAgents provides clear visual feedback on the task state and system progress throughout the scheduling process.
 We implement a Flask-based dashboard interface designed to make the planning workflow interactive, legible, and actionable.
@@
 At the end of each simulated day, the system generates a daily summary that highlights key events and decisions.
-Specifically, we use GPT-4o to generate the daily summaries, with prompt details provided in the Appendix A.5.
+Specifically, we use GPT-4o to generate the daily summaries, with prompt details provided in the Appendix C.5.
 As shown in Figure 3(B), one example summary is presented within the Interactive Chat Interface and reads: “Invitations for the HCI seminar were sent to Profs.
@@
 unfilled); and (3) combined speaker and slot context, which merges the two but omits higher-level summaries.
-Prompt templates for all conditions are provided in Appendix B.
+Prompt templates for all conditions are provided in Appendix D.
 To evaluate the performance of each approach, we constructed a test suite of 30 examples, all sourced from real user study sessions and curated to reflect a diverse set of scheduling scenarios.
@@
 4.1.2. Results and Findings
-Prompting
-Techniques
-F1 Score
+As summarized in Table 1, our summary-based context approach achieves the highest F1 score for predicting relevant policies given a specific problem state.
+The improvement is statistically significant when compared to all three baseline conditions, with an average F1 score of 0.70, demonstrating the effectiveness of using structured progress summaries to guide policy selection.
+Notably, simply aggregating speaker- and slot-related context without summarization does not improve performance; the combined-context baseline yields only 0.27 average F1, indicating that raw context alone is insufficient for accurate policy retrieval.
+Based on these results, we adopt the summary-based approach as the default mechanism for selecting policies, which are then passed to the orchestrator agent to guide plan generation in a policy-aligned manner.
+Prompting Techniques F1 Score
 Mean SD
@@
 Statistically significant improvements (independent two-sample t-test, $p<0.001$) over all baselines are marked with ***.
-As summarized in Table 1, our summary-based context approach achieves the highest F1 score for predicting relevant policies given a specific problem state.
-The improvement is statistically significant when compared to all three baseline conditions, with an average F1 score of 0.70, demonstrating the effectiveness of using structured progress summaries to guide policy selection.
-Notably, simply aggregating speaker- and slot-related context without summarization does not improve performance; the combined-context baseline yields only 0.27 average F1, indicating that raw context alone is insufficient for accurate policy retrieval.
-Based on these results, we adopt the summary-based approach as the default mechanism for selecting policies, which are then passed to the orchestrator agent to guide plan generation in a policy-aligned manner.
-5.2. Edge Case Detection Experiments
-5.2.1. Experiment settings
-Edge case detection involves identifying whether a respondent’s message presents a scenario that falls outside the coverage of existing coordination policies.
-For example, a speaker requesting to present via Zoom, when no policy has addressed virtual presentations, should be flagged to the user for clarification and further instruction.
-To enable reliable detection of such edge cases, we adopt a few-shot prompting strategy with access to existing policies (denoted as ”few-shot w/ policy”).
-We compare this approach against two baselines: (1) zero-shot prompting without any demonstration examples, and (2) few-shot prompting without access to policy references (denoted as ”few-shot w/o policy”).
-Concrete prompt templates for each setup are provided in Appendix B.2.
-To evaluate performance, we constructed a test suite of 100 examples, manually validated by the lead authors to ensure coverage and quality.
-The construction methodology for this test suite is detailed in Appendix B.2.
-All experiments were conducted using GPT-4o, with temperature set to 0.7 and a maximum token limit of 100.
-We report the average accuracy across three independent runs for each condition.
-5.2.2. Results and Findings
-Prompting Techniques
-Accuracy
+Prompting Techniques Accuracy
 Mean SD
@@
 Statistically significant improvements (independent two-sample t-test, p ¡ 0.01) over both baselines are marked with **, improvements over one baseline are marked with †.
+4.2. Edge Case Detection Experiments
+4.2.1. Experiment Settings
+Edge case detection involves identifying whether a respondent’s message presents a scenario that falls outside the coverage of existing coordination policies.
+For example, a speaker requesting to present via Zoom, when no policy has addressed virtual presentations, should be flagged to the user for clarification and further instruction.
+To enable reliable detection of such edge cases, we adopt a few-shot prompting strategy with access to existing policies (denoted as ”few-shot w/ policy”).
+We compare this approach against two baselines: (1) zero-shot prompting without any demonstration examples, and (2) few-shot prompting without access to policy references (denoted as ”few-shot w/o policy”).
+Concrete prompt templates for each setup are provided in Appendix D.2.
+To evaluate performance, we constructed a test suite of 100 examples, manually validated by the lead authors to ensure coverage and quality.
+The construction methodology for this test suite is detailed in Appendix D.2.
+All experiments were conducted using GPT-4o, with temperature set to 0.7 and a maximum token limit of 100.
+We report the average accuracy across three independent runs for each condition.
+4.2.2. Results and Findings
 As shown in Table 2, our few-shot prompting with policy method significantly outperforms both baselines, with statistically significant improvements across evaluation runs.
 These results demonstrate the effectiveness of incorporating policy context into the detection process, enabling the model to first analyze existing policies before determining whether a given case falls outside their scope.
-Based on these findings, DoubleAgents adopts the policy-aware few-shot approach as its default mechanism for edge case detection.
+Together, DoubleAgents adopts the policy-aware few-shot approach as its default mechanism for edge case detection.
 This ensures that novel or ambiguous cases are reliably flagged to users, enabling timely clarification and maintaining a high-quality user experience.
 5. User Study
-6.1. Goal and Setup
-To investigate the user-perceived effectiveness of DoubleAgents’ features and the dynamics of trust in human-AI collaboration, we recruited 10 participants to use DoubleAgents across two separate sessions.
-Our goal was to assess users’ perceived system usefulness, their reliance on the systems, and their trust and comfort in using AI to automate tasks had shifted.
-During each session, participants worked in a simulated environment to schedule a seminar with four dates and four speakers, while the system gradually adapted to their preferences.
-Participants completed a questionnaire assessing their trust in 15 different system features at three points: before the first session, after the first session, and after the second session.
-This allowed us to examine whether and when their trust and perceptions of the system changed over time.
+5.1. Goal and Overview
+To investigate how users engaged with DoubleAgents’ features and how interactive simulation shaped their use of agentic behavior, we recruited 10 participants to use DoubleAgents across two separate sessions.
+Our goal was to assess users’ perceived system usefulness, their understanding and reliance on the system, and how their understanding and calibration of agent behavior evolved over time.
+During each session, participants worked in a simulated environment to schedule a seminar with four dates and four speakers, while the system gradually adapted to participants’ values and preferences via their policies, email templates, and stop hooks.
+Participants completed two questionnaires.
+First, an AI comfortableness survey measured trust and alignment needs for proactive actions and comfort with automating planning and execution, administered before Session 1, after Session 1, and after Session 2.
+Second, a System perception survey captured UX and perceptions of 15 system features after Session 1 and after Session 2, enabling us to track how repeated use shifted alignment strategies, delegation boundaries, and reliance.
+Detailed study procedures and data analysis methods are provided in Appendix A.
 5.2. Participants
@@
 Finally, 1 participant had no prior experience, having never hosted or coordinated any tasks.
-6.3. Procedures and Data Analysis
+5.3. Result
+5.3.1. Consistently Positive User Experience
+Across both sessions, all 10 participants reported a consistently positive experience with DoubleAgents.
+Overall system ratings remained high and showed a slight improvement (from 6.30/7 to 6.62/7).
+Participants also perceived the system as effective in generating optimal plans, actions, and emails (5.90/7 to 5.88/7), and judged it as robust (5.60/7 to 5.88/7).
+Both mental and physical demand remained low across sessions, decreasing from 2.90/7 to 2.13/7 and from 1.50/7 to 1.13/7, respectively.
+Participants reported high perceived performance (6.30/7 to 6.75/7), low frustration (1.80/7 to 1.25/7), and a near-significant reduction in effort (3.20/7 to 1.88/7, $p=0.054^{{}^{\wedge}}$).
+Together, these results indicate a consistently positive and efficient user experience across sessions, with effort showing the strongest improvement.
+5.3.2. Increasing User Reliance Driven by DoubleAgents’ Design
+Despite largely stable system-level scores, participants’ perceived reliance on DoubleAgents increased significantly over sessions (from 4.90/7 to 5.75/7, $p=0.02^{*}$).
+Participants characterized this shift not as “unconditional reliance”, but as increased comfort with delegation as the system’s behavior became more user-aligned, legible, and shapeable—enabled by clearer controls, predictable execution patterns, and explicit points of intervention.
+A correlation analysis between reliance and feature-level perceptions (Spearman; FDR-corrected) showed that reliance was associated with features supporting control and explainability, including state visualization ($r=0.78$, $p<0.05^{*}$) and confirmation-based user control ($r=0.51$, $p<0.05^{*}$).
+For example, P6 found the Communication Visualization section of the interface useful in showing the full email exchange histories between the organizer and each speaker.
+She mentioned, “Everything is consistent inside the system—when some speakers were not followed up, I could map it back to the visualization and confirm accuracy… [It] helped me build trust… For tomorrow, I’ll spend less time checking the email replies [inside the chat interface] one by one, and instead [just check] the Communication Visualization on the side.” Increases in reliance over time were strongly linked to DoubleAgents’ human-AI alignment features.
+In particular, reliance increased with perceived system robustness demonstrated through interactive simulation ($r=0.54$, $p<0.05^{*}$), and was reinforced by repeated successful issue flagging via stop hooks ($r=0.65$, $p<0.01^{**}$) and perceived value alignment through coordination policies and email writing ($r=0.54$, $p<0.05^{*}$).
+Reliance was negatively correlated with mental demand (r=0.47r=-0.47, $p<0.05^{*}$) and showed no significant correlation with other system experience or perception items.
+Collectively, these findings suggest that reliance increased not merely due to perceived system capability or general UX, but because of DoubleAgents’ support for alignment in design.
+The following section details how participants experienced and leveraged this alignment through interactive simulation.
+Figure 7.
+Average user perception scores toward comfort with automating the (a) planning and (b) execution stage in seminar organizing, across three stages (average scores across 10 users: before Session 1, after Session 1, and after Session 2).
+denotes statistical significance at $p<0.1$, * at $p<0.05$, ** at $p<0.01$, *** at $p<0.005$.
+5.3.3. Increasing User Comfort with AI Automating Human Coordination Tasks: Contributing Alignment Supports.
+Over interactive simulations, participants became more comfortable delegating both planning and execution tasks to agentic AI.
+For planning (Figures 7a, participants became more comfortable letting the AI visualize current status (4.80/7 to 6.50/7, $p<0.05^{*}$), propose time assignments (5.10/7 to 6.38/7, $p<0.1^{\vee}$), decide daily next steps (4.70/7 to 6.50/7, $p<0.005^{***}$), and monitor/flag issues (5.50/7 to 6.50/7, $p<0.01^{**}$).
+For execution, comfort increased for drafting-only emails across all four types (from around 6.0 before to nearly 7.0/7 after, $p<0.01^{\vee}$), and also increased for draft-and-send tasks on users’ behalf—e.g., initial outreach rose from around 2.7/7 to 4.6/7 ($p<0.05^{*}$), with similar gains for follow-ups and final confirmation ($p<0.1^{\vee}$).
+Participants attributed this increased comfort to DoubleAgents’ alignment supports: interactive simulation; user-editable coordination policies that guide next-step decisions; email templates and style controls that realize those actions; and stop hooks for issue escalation.
+Together, these supports advanced bidirectional alignment: the AI agent’s actions became better aligned with users’ needs and values, and users developed greater understanding and trust, becoming more comfortable with the AI taking initiative in automating this human coordination task.
+Interactive simulation as a testbed for policy, email, and stop-hook calibration.
+Simulations supported alignment by letting participants safely probe edge cases, identify where the agent’s behavior diverged from expectations, and iteratively refine (1) what the agent should do (coordination policies), (2) how it should carry out those actions in email (templates and style), and (3) when it should pause for confirmation (stop hooks).
+Participants emphasized that interactive simulations increased confidence in their perception of the system’s robustness (5.80/7 to 6.38/7).
+They found simulated test cases realistic and representative of real-life scenarios (5.90/7 to 6.75/7), and especially useful for surfacing potential problems (5.80/7 to 6.50/7).
+P3 described how simulation exposed realistic exceptions and prompted advance policy-setting: “The simulations [over the 2 days] felt realistic and true to the person [they] were simulating, like their personality and emails… The simulation elicited possible problems that would happen in real life… Like [having] people asking for funding or [needing] to reschedule—it forced me to think about how I would handle that in real life and to create relevant policies, and that was good.
+I can trust the system to handle that in the future for me if scenarios like that occur, since it already happens in the simulation.
+Email templates and coordination policies fostered increased comfort with AI automation.
+Participants’ comfort with AI automating planning and execution increased over time as the system’s actions became more consistent, better grounded in the users’ own values and rationale, and increasingly “in their voice.” Carrying over value-aligned coordination policies made participants more comfortable delegating planning—e.g., having the AI decide daily next steps, track and visualize progress, make assignment decisions, and monitor/escalate issues when cases did not fit existing policies (Figure 7a).
+Email-template carryover particularly supported execution delegation.
+Comfort increased for drafting-only emails and for draft-and-send tasks (Figure 7b).
+One exception was drafting and sending clarification emails without any user confirmation, which showed little change (3.00/7 to 3.90/7 to 3.90/7).
+P10 noted that clarifications often require nuanced, user-specific judgment: “For clarification emails, I might want to double-check… [they’re] usually not in the plan [or policies]… I’m not comfortable [offloading that] to AI…” As a result, participants preferred to remain in the loop for clarification messages.
+Overall, participants described an early sense of being understood and represented by the system, as if it “got them,” when iterating on email templates and coordination policies.
+For example, P4 noted that edits to greetings, closings, and phrasing carried over into later drafts: “Compared to the last [usage], DoubleAgents get my style… so I edited fewer times this time.” Similarly, after refining policies, P3 observed that the agent’s choices matched their expectations: “I found that the agent’s recent actions followed exactly the policies I had in mind—super aligned.” This growing alignment helped the system reflect their intended actions and communication style, increasing their comfort with offloading tasks to AI agents.
+Stop hooks helped establish safe delegation boundaries, fostering users’ comfort with delegation.
+Participants highlighted stop-hook flagging as particularly important for maintaining safe delegation boundaries.
+For monitoring and surfacing potential issues, P1 described issue flagging as “one of the most important features of the system” and shared, “[With the issue flagging,] I skim [through the other features] now… It made me trust the system because I know it won’t just execute something random… It flags when it doesn’t know… That helps elicit my values and how I want to handle things—and [then DoubleAgents] respects that…” P8 described a case where flagging prevented a misinterpretation of progress—when they skimmed the respondent agent’s reply and thought everything was resolved, the flag showed that it was not.
+Thus, with these “stop hooks” continuously monitoring for exceptions and prompting intervention when needed, participants felt much more comfortable offloading tasks to the AI.
+Figure 8.
+Average user perceptions of alignment needs, expectations, and trust in agentic AI systems (average scores across 10 users: before Session 1, after Session 1, and after Session 2).
+* denotes statistical significance at $p<0.05$ and ** at $p<0.01$.
+Increasing Expectations and Needs for Alignment in Agentic AI
+Across sessions, participants reported stronger expectations for how proactive AI in general should behave and align with them (Figure 8).
+In the AI Comfortableness Survey, participants reported a significant increase in wanting the AI’s actions to align with their preferences, values, and priorities, from 6.10/7 (before using the system, based on their previous AI experiences) to 7.00/7 after two uses of DoubleAgents ($p<0.05^{*}$).
+Participants’ desire to stay in control remained high (5.30 to 5.75).
+They also reported an increased need for detailed system rationales and explanations before the system acts, from 5.00/7 to 6.00/7 ($p<0.05^{*}$).
+Participants framed clear and abundant explanations as a core alignment mechanism for building an accurate mental model of the AI’s decision process, enabling verification, and determining when and how to intervene.
+P6 mentioned, “[With the system explanation] I felt like I could tell exactly what it’s doing at every stage so there’s a lot of transparency… if I didn’t get that, I would feel lost [and] not know what’s going on…” We interpret this as evidence that DoubleAgents’ interaction design helped participants recognize the importance of alignment and articulate stronger alignment expectations for future proactive AI usage.
+Furthermore, participants also reported statistically significant increases in trust to let similar AI systems act on their behalf, from 3.20/7 to 5.00/7 ($p<0.01^{**}$), even as their assessments of AI reliability remained relatively the same.
+This suggests that alignment features can increase comfort with delegation beyond improvements in perceived raw capability.
+P1 noted that even though DoubleAgents represented only one specific experience, “I felt more optimistic… [and] in seeing similar systems like DoubleAgents in the future, I would generally trust such AI systems more.”.
+6. Deployment Studies
+6.1. Goal
+As the in-lab study demonstrated, the simulation was helpful, and participants felt they were getting more comfortable use agentic AI help planning and executing human coordination tasks.
+However, would real-world seminar organizers find DoubleAgents useful and relevant?
+If so, how would they use it-would they plan, draft, and send invitation, follow-up, or clarification emails in different ways?
+Would alignment be established in a similar manner?
+What additional challenges might arise in real-world use?
+To explore these questions and evaluate the real-world deployability of DoubleAgents, we invited organizers who were managing actual events across three real-world deployments.
+6.2. Deployment Participants and Procedures
+6.2.1. Use Cases and Participants
+In our deployment study, DoubleAgents was used in three real-world event-planning workflows involving two organizers (P11, P12).
+P11 used DoubleAgents to coordinate a Fall 2025 HCI seminar series (Aug 2025; Deployment 1, D1), inviting 10 speakers from a pool of 23 to fill 10 slots; we observed and logged system use for 7 days during this process.
+P11 later used DoubleAgents again for a Spring 2026 HCI seminar series (Jan 2026; Deployment 3, D3), inviting 11 speakers from a pool of 20 to fill 11 slots; we tracked use for 11 days.
+P12 used DoubleAgents to organize a conference workshop panel (Sep 2025; Deployment 2, D2), inviting 5–6 panelists from a pool of 18 to fill a single slot; we tracked use for 7 days.
+We recruited P11 and P12, two experienced seminar organizers (one male, one female; average age 31), via personal connections.
+Both were PhD students who had previously organized departmental research seminars and had prior experience using generative AI for writing emails.
+6.2.2. Introduction, Setup, and Interactive Simulation Rounds
+Organizers were first informed about the study and provided consent.
+They were then introduced to the study procedures and the DoubleAgents system through a demonstration, which included planning–execution loops, illustrated with a simulation example from the user study.
+Organizers also completed the initial survey covering AI perception and demographics.
+To set up the system, we disabled the respondent module from the lab version of DoubleAgents and connected the SendEmail() and checkEmailUpdates() tools to Gmail’s SMTP service, enabling organizers to send emails and receive replies through their personal Gmail accounts.
+For the simulation rounds, organizers prepared a simulation spreadsheet using their real organizing data (including real potential speakers’ names), first with placeholder email addresses and later with the real ones, and provided their Gmail SMTP credentials.
+Clear instructions and consent procedures ensured again that participants understood data usage, privacy rights, and the option to withdraw.
+Before connecting their real email accounts, participants completed a 2-hour interactive simulation session using pre-generated speaker personas based on the real candidates, allowing them to safely familiarize and themselves align with the system’s workflows.
+Finally, they tested the real email-sending module twice by sending messages to placeholder addresses (the researchers’ and organizers’ personal emails) to confirm successful delivery; afterward, we switched the recipient addresses to the invitees’ real Gmail accounts.
+6.2.3. Live Email Execution and Response Handling
+After the simulations, organizers used DoubleAgents to initiate the coordination process and track it for over a few days.
+They continued receiving system updates over the following week (D1: 7 days, D2: 4 days, D3: 11 days).
+We communicated with participants consistently, logging interactions, usage patterns, and feedback on the system’s support for scheduling, tracking, and learning.
+Organizers’ system interactions, including email sends, scheduling updates, and system adaptations, were tracked.
+Qualitative feedback was collected via check-in meetings, while quantitative system logs were collected for system learning over the week.
+6.3. Deployment Results
+6.3.1. DoubleAgents’ Perceived Usefulness and Relevance in Real-Life Human Coordination Problems.
+After the interactive simulations, both organizers (P11 and P12) reported that DoubleAgents was highly relevant to their existing seminar-organizing workflows.
+They highlighted that the system fit seamlessly into their email-heavy tasks, and that its control settings and edge-case flagging helped them handle complex coordination scenarios that conventional scheduling tools (e.g., Calendly) cannot easily manage.
+P11 emphasized that crafting good email templates is a delicate and important task because it requires personalizing messages to “make the speakers feel they are special,” even when requesting a favor.
+Both P11 and P12 found DoubleAgents capable of generating strong starter emails for them to edit early on, as well as producing acceptable final emails to send.
+P12 especially appreciated having explicit coordination policies in the system to guide next-step decisions in ways that matched his decision-making habits and values.
+He also found it useful that rationales were always visible and editable, in contrast to AI tools that hide their reasoning.
+After using DoubleAgents, P12 remarked, “I trust AI [systems] for specific use cases more,” explaining that he found DoubleAgents dedicated to and capable of generating high-quality emails.
+P11 found the color-coded visualization of speaker constraints and response states (e.g., not yet outreached, tentative yes, yes, no) especially useful; after D1, she adopted a similar color-coded scheme for other tracking tasks she previously managed in spreadsheets.
+6.3.2. Interactive Simulation as an Alignment Method for Complex, Realistic Tasks.
+Both P11 and P12 emphasized that interactive simulation was important for understanding DoubleAgents and for setting up initial coordination policies and email templates.
+They appreciated that simulations used real invitee names and incorporated persona cues and edge cases (e.g., funding requests, cancellations, Zoom requests), which helped them anticipate challenges and craft templates that would transfer to live deployment.
+For example, P12 used early simulation rounds to iterate on what information to include and where to place key details (e.g., the workshop webpage link) in his email templates, similar to how he had previously tested invitation templates by emailing himself.
+P11’s D1 simulation surfaced multiple policy- and tone-sensitive situations (e.g., rescheduling after a slot had already been assigned), which prompted iterative template and policy revisions.
+P11 reported that the system’s drafts were largely aligned with her language, but she removed or adjusted certain lines to better match her preferences and to clarify what information should be included when edge cases occur.
+Overall, P11’s D1 interactive simulation took four days across five meetings: the first two meetings focused on ensuring the system ran smoothly for a high-complexity scenario (20+ candidates, 10 slots, and substantial context per speaker).
+During setup, P11 revised invitation and follow-up templates more than four times and updated speaker descriptions so email generation could incorporate richer context.
+After we refined the system and onboarding protocols, P12’s D2 simulation took two days across two meetings.
+P12’s panel invitation task was also simpler (e.g., shorter descriptions and a single-slot panel), which further reduced the needed time.
+Both organizers reported that practice simulation rounds were valuable, and that sending test emails via SMTP to his own personal gmail helped him gradually understand system capabilities and limitations.
+Overall, these deployments suggest that simulation and onboarding effort scales with task complexity and the amount of required context.
+During D1, P11 specifically highlighted the importance and difficulty of providing sufficient context for highly personalized emails, especially for real-world cases.
+She noted that “[persona description management] is really tricky.” For instance, one invitee had previously coordinated with a different organizer in a prior semester (Friday availability), but the D1 seminar series had shifted to Thursdays.
+To avoid missing or conflicting context across past threads, schedule changes, and templates, P11 updated the persona description used for simulation with an additional line noting that the invitee had been told the wrong day in the past (Friday 4–5pm) and that the seminar was now on Thursday 4–5pm.
+In another case, P11 had already discussed the seminar with a speaker on Discord but initially omitted this context during early simulations; she then added it to the persona description so the follow-up email could adopt the appropriate tone and content.
+After rehearsing with real-speaker data in simulation, P11 finalized a preferred template for live sending.
+P11 also noted that even in D3 (her second deployment), simulation remained necessary because “real life is super complex.” She observed new patterns in edge cases, such as delayed responses or vague “check back next month” replies from senior faculty, which led her to revise policies (e.g., longer follow-up intervals) and adopt a more formal email tone.
+Across both deployments, simulation served not just as a testing ground, but as a means of progressive alignment: helping users adapt policies, templates, and expectations to shifting social and contextual dynamics.
+These findings suggest that onboarding and setup effort scale with task complexity and that simulation continues to play an ongoing role in supporting user-aligned adaptation, even after initial deployment.
+6.3.3. Example: Live Email Execution, Response Handling, and Edge Cases
+During D3, P11 sent emails to 11 selected speakers and relied more heavily on templates and system-generated drafts on Day 1 (a Friday), reflecting increased familiarity with the workflow.
+DoubleAgents then suggested waiting for responses over the following two days due to the weekend, which P11 accepted.
+Late on Day 3 (Sun), Invitee J and Invitee A replied with questions, which DoubleAgents surfaced in P11’s Day 4 (Mon) DoubleAgents session.
+On the morning of Day 4, Invitee B replied with availability.
+Based on these responses, DoubleAgents suggested replying first to Invitees A and J, but not to Invitee B, to preserve future scheduling flexibility and avoid potential conflicts.
+P11 accepted this suggestion and drafted replies accordingly.
+Specifically, Invitee A asked a funding-related question and included a personal greeting (e.g., “Happy New Year… if it doesn’t materialize…”).
+P11 revised the AI-generated reply by incorporating similar phrasing from their own writing style, but chose not to save this edit as a reusable template, noting that it reflected a personal, context-specific language choice.
+On Day 5, Invitees I and J replied, and DoubleAgents suggested allocating them to earlier-semester slots, which P11 accepted and responded to.
+Similar reply-and-assign cycles occurred on Day 6.
+On Day 7, DoubleAgents suggested following up with all the invitees who had not yet responded, in accordance with P11’s newly updated D3 policy to follow up every four working days (rather than every two, as in D1).
+P11 proceeded using a personalized follow-up template that incorporated questions added on Day 4.
+Also on Day 7, P11 handled additional questions and confirmed slots with Invitees C and H, who had replied promptly with apologies and availability.
+Later that evening, Invitee I emailed to report a schedule change and requested reassignment to an unclaimed date.
+On Day 8, DoubleAgents suggested reassigning Invitee I and requesting additional availability from Invitees H and B to resolve emerging conflicts.
+P11 accepted and carried out these actions.
+Over the weekend (Days 9 and 10), Invitee H responded with new dates but noted that P11 should follow up again the following month, as their schedule was still tentative.
+DoubleAgents accordingly scheduled a follow-up for Day 35, while on Day 11 it prompted P11 to address remaining unresponsive and unresolved cases.
+Collectively, these interactions illustrate the complexity and multi-threaded nature of this real-world task.
+As P11 reflected, “Even though I do this [seminar scheduling] all the time, [DoubleAgents] was really helpful and saved me a lot of time… Policy-abiding [actions and template-following] emails are really useful.
+The new policies and templates I developed through this semester’s simulation—since everyone is a professor—worked really well.
+The simulation part cannot be skipped.” This case highlights DoubleAgents’ strong use case, demonstrating how simulation, policies, and email templates work together to support alignment between the system’s actions and users’ intentions.
+7. Discussion
+7.1. Human-AI Alignment via Interactive Simulations
+As AI systems evolve from passive tools to agentic collaborators that plan and act on users’ behalf, human–AI alignment becomes not just a technical challenge but an interactive one.
+Our findings suggest that interactive simulation offers a compelling design mechanism to support this alignment by foregrounding user agency, surfacing system behavior early, and fostering a co-adaptive relationship between humans and AI.
+Across our lab study and deployments, we observed that interactive simulation enabled users to engage with alignment as a gradual process, rather than a one-time configuration step.
+Participants entered the system unsure of their own preferences and reluctant to delegate; yet through repeated simulated cycles, each revealing system actions, uncertainties, and edge cases, they clarified values, authored and refined policies, and ultimately built enough user comfort to let the system act on their behalf.
+This shift mirrors real-world dynamics of human delegation: people do not fully trust a new assistant immediately but gain confidence through concrete interactions over time.
+Simulation thus serves multiple alignment functions simultaneously.
+First, it externalizes system behavior, making invisible decision logic (e.g., why the agent followed up or chose a certain speaker) legible to users.
+Second, it creates safe interaction spaces, akin to usability testing for AI plans.
+Like sandboxed environments in traditional software testing, simulations help users anticipate downstream consequences without real-world stakes, crucial in settings like email coordination where tone, timing, and etiquette are socially sensitive.
+Third, simulation supports user-side alignment, helping users identify preferences they hadn’t articulated upfront (e.g., when to prioritize seniority, how long to wait before following up, or what tone feels appropriate across contexts).
+Rather than requiring exhaustive foresight, DoubleAgents invites users to iterate on policies as new situations emerge.
+This design reframes alignment as a human-AI collaboration process: the agent proposes actions, but the human remains in the loop, interpreting, approving, and updating the future behavior of the system.
+In contrast to approaches that aim to extract user goals from demonstrations or feedback, we emphasize the value of direct, interactive co-construction of alignment artifacts, such as stop hooks, policy rules, and email templates, which evolve with use cases.
+This opens new design directions: how can we better scaffold early-stage alignment for users who don’t yet know their own preferences?
+What kinds of interfaces help users develop trust while retaining oversight?
+How might simulation be extended to other high-stakes or interpersonal agentic domains, such as hiring, negotiation, or caregiving?
+Our work suggests that alignment is not a prerequisite to agentic AI, but a continuous, situated interaction practice.
+By embedding simulation into the human–AI workflow, we give users the tools not just to observe and react to agent behavior, but to actively shape and steer it.
+In doing so, DoubleAgents contributes to a broader vision of agentic AI that is transparent, participatory, and grounded in human values from the start.
+7.2. Designing User Onboarding Experiences for Agentic AI
+Onboarding is a critical site for agentic AI systems.
+Unlike traditional tools, agentic systems require users to make early decisions about goals, decision boundaries, tone and formality, and what to delegate.
+They also create distinct discoverability challenges: users often do not know what the agents can do and what they can do.
+Key modes and safeguards (e.g., auto-send vs.
+draft-only) may be implicit, and failures can appear only in rare or edge cases—so there is no single, linear “progress bar” that reveals system status or readiness.
+As a result, effective onboarding is necessary for users to form accurate mental models, judge reliability, and feel comfortable delegating and deploying the agent in higher-stakes settings (e.g., a seminar organizer connecting to Gmail to send real invitations to senior faculty—who may also be potential postdoc advisors—where mistakes can harm professional relationships and reputation).
+In DoubleAgents, we found that interactive simulation provides a powerful onboarding scaffold, helping users observe, evaluate, and refine system behavior before real-world use.
+Rather than presenting configuration as a static setup phase, DoubleAgents embeds onboarding through repeated rounds of interactive simulation.
+During these rounds, edge cases (e.g., a speaker canceling or requesting Zoom) expose users to socially complex scenarios, help them become familiar with the system’s behavior, and elicit their personal preferences.
+Participants in our study often hesitated to delegate initially, not due to lack of interest, but due to uncertainty about agent behavior.
+Using repeated rounds of interactive simulation as an onboarding task lowered this barrier by helping participants move past the initial learning curve and novelty effects, and by revealing how the agent behaved across diverse cases.
+Like training a new assistant, onboarding becomes a co-configuration process in which expectations, autonomy, perceived usefulness, and trust are developed together.
+We argue that onboarding for agentic AI should not be treated as a brief tutorial animation or tooltip overlay.
+Instead, it can be designed as a structured, interactive process where users and systems collaboratively define goals, boundaries, and escalation rules.
+DoubleAgents shows how early simulations can accelerate alignment, ease delegation, and set the foundation for long-term, trustworthy collaboration.
+8. Limitations and Future Work
+Our work has several limitations that point to promising directions for future research.
+First, DoubleAgents was initially evaluated through simulated seminar-organizing scenarios.
+While these simulations were reviewed by domain experts and designed to reflect realistic coordination challenges, they cannot fully capture the diversity and unpredictability of real-world interactions.
+Future work could strengthen alignment by grounding both the coordination agent and respondent simulations in users’ real historical data, such as past email threads, organizing decisions, and communication patterns.
+This would enable more personalized modeling of users’ values, tone, and escalation preferences, as well as more behaviorally faithful respondent agents.
+Second, DoubleAgents currently relies on different language models for different subtasks (e.g., o3 for reasoning and 4o-mini for generation).
+Although our technical evaluation suggests this division performs adequately, future models with stronger social reasoning and long-horizon planning capabilities may further improve coordination quality.
+In addition, tighter integration between user-specific data and model reasoning, such as learning when and how to apply policies from prior interactions, remains an open challenge.
+Third, while we conducted real-world deployments with two seminar organizers, discussions with additional organizers revealed coordination scenarios that our current system does not fully support.
+For example, bilingual invitations, multi-organizer decision-making, and negotiation among organizers introduce shared authority, language switching, and conflicting preferences.
+Extending DoubleAgents to support collaborative alignment, where policies, stop hooks, and responsibilities are jointly defined and negotiated, represents an important direction for future work.
+Finally, all study participants were experienced users of AI tools, which may have positively influenced their willingness to delegate and their trust in the system.
+Future studies should include participants with a wider range of AI familiarity to better understand onboarding challenges, mental model formation, and how interactive simulation can support alignment for users with less technical background.
+9. Conclusion
+We introduced DoubleAgents, a system for aligning agentic AI through interactive simulation, policy-grounded planning, and transparent oversight.
+In coordination tasks like speaker scheduling, alignment is not just about correctness; it’s about respecting user preferences, adapting to edge cases, and maintaining trust.
+By embedding simulation into the agent workflow, DoubleAgents helps users rehearse edge cases, refine their preferences, and gradually scale up delegation.
+Our lab and real-world deployments show that users increasingly trusted and customized the system, not because it was perfect, but because they could iteratively shape it.
+This work offers a practical path toward alignment in real-world agentic AI: one grounded not in static prompts, but in interaction, iteration, and co-adaptation.
+10. Generative AI Usage Disclosure
+We used generative AI tools (e.g., OpenAI’s ChatGPT) as support during several parts of this project.
+Specifically, we used them to help debug the system, draft non-critical implementation and UI code, write analysis scripts for logs and study data, and edit text for grammar and clarity during paper writing.
+Appendix A Seminar Planning: A Human Coordination Problem
+Human coordination—such as assigning work shifts or scheduling job interviews—is both cognitively demanding and socially delicate.
+Unlike tasks with stable inputs and predictable outcomes, coordination requires managing uncertainty, reconciling multiple perspectives, and sustaining relationships over time.
+Even small lapses—missed follow-ups, unclear tone, or misaligned expectations—derail coordination and strain relationships.
+These challenges make coordination both essential and persistently difficult in many areas of knowledge work.
+We study these challenges through one concrete and consequential setting: organizing an academic seminar series.
+A typical series spans an entire semester and involves inviting a new external speaker nearly every week, amounting to 8–12 speakers per term.
+Early on, organizers must identify potential speakers, gather their availability constraints, and balance their own preferences-such as ensuring topical diversity across weeks and securing highly sought-after senior researchers who are harder to schedule.
+However, the most unpredictable and challenging part is dealing with responses.
+Some people do not reply at all and require multiple follow-ups to elicit a response.
+Some give only vague and tentative replies that need repeated confirmation.
+Some can be flakey and unreliable, offering slots but later canceling.
+And many unexpected problem cases arise: people might be on parental leave, ask for visa letters or funding support, recommend a different speaker, request a different date or talk modality, or raise logistical demands.
+In such human coordination tasks, you can rarely predict what is coming.
+Therefore, conventional scheduling tools such as Calendly [^1] https://calendly.com/ or When2Meet [^2] https://www.when2meet.com/ are ill-suited to this setting.
+They assume stable constraints and one-shot coordination, yet seminar scheduling requires ongoing negotiation, persuasion, and relationship maintenance.
+Unlike ordinary meeting scheduling, seminar speakers are doing you a favor, which means you cannot simply send them a link to pick a time—you must persuade and negotiate.
+Social dynamics and etiquette also play a large role: organizers must time their emails carefully to avoid overwhelming, annoying, or losing momentum with potential speakers, who are often busy scholars.
+Knowing when to nudge, follow up, or wait is subtle.
+Email remains the standard medium for professional and academic outreach, allowing organizers to adapt tone—formal, casual, or enthusiastic—to different speakers in ways scheduling links cannot.
+It also supports sharing rich context, such as talk topics, background, past speakers, or travel plans—far beyond simply choosing a calendar slot.
+Thus, experienced organizers adopt personal strategies to cope with such a demanding task.
+For example, some follow a rule of “wait two days, then follow up” to nudge busy people who may not read emails right away, while others wait a full week before reaching out again.
+Some prefer contacting close collaborators or junior researchers first, while others prioritize securing prominent professors or distant speakers early to reduce travel constraints.
+When conflicts arise between speakers, organizers might “give faculty priority and ask students to be more flexible,” carefully deciding whom to approach for alternative availability to minimize disruption.
+For non-responders or last-minute cancellations, they must judge when and how often to follow up before moving on to new options—and sometimes even decide when it is appropriate to “beg” busy invitees for additional availability.
+Because plans are constantly evolving, organizers become locked in a cognitively and time-intensive loop of tracking shifting states, replanning, and executing revised plans as new information arrives—yet they often cannot proceed with the next steps or finalize allocations until replies are received.
+Since seminar scheduling requires nuanced, context-dependent decisions that reflect each organizer’s values and preferences, it is difficult to fully automate.
+An effective seminar scheduling tool should not attempt to handle the process independently; instead, it must learn from and adapt to the organizer, respecting social norms, relational dynamics, and the inherent unpredictability of responses.
+Appendix B User Study Procedures and Data Analysis
 Test Scenario Preparation.
-We select eight distinct personas identified by three organizing experts, representing realistic and diverse roles, research areas, and backgrounds (e.g., a fifth-year PhD student studying VR currently on the job market; a busy junior professor in NLP who spends much of her time on her startup).
-For each persona, we specify reply behaviors (responsiveness, email writing style; e.g., slow responder, writes only short emails) and optional issues (potential concerns they might raise with speakers; e.g., requesting additional funding even though their in-person visit is already funded).
+We selected eight distinct personas identified by three organizing experts, representing realistic and diverse roles, research areas, and backgrounds (e.g., a fifth-year PhD student studying VR currently on the job market; a busy junior professor in NLP who spends much of her time on her startup).
+For each persona, we specified reply behaviors (responsiveness, email writing style; e.g., slow responder, writes only short emails) and optional issues (potential concerns they might raise with speakers; e.g., requesting additional funding even though their in-person visit is already funded).
 We split them into two groups and created seminar organizing instances.
@@
 Day 1, Preparation and Setup (30min)
-Participants first provided informed consent and completed a short AI perception survey assessing their perceptions and experiences with AI automation tools, human coordination tasks, and demographics.
+Participants first provided informed consent and completed an AI comfortableness survey assessing their perceptions and expectations with AI automation tools, human coordination tasks, and demographics.
 They also edited a few seeded templated seminars organizing emails according to their tone to feed the system their writing style.
 They were then introduced to the DoubleAgents interface and task.
-One researcher gave a brief walkthrough using a visual diagram and demonstration, covering the system’s core components, including the planning-execution orchestrator agent loo, respondent agent simulation, policy and context modules, visualization, and the flag mechanism.
+One researcher gave a brief walkthrough using a visual diagram and demonstration, covering the system’s core components, including the planning-execution coordination agent loop, respondent agent simulation, policy and context modules, visualization, and the flag mechanism.
 Day 1, First System Usage and Evaluation (60min)
@@
 During the task, participants articulated and shared their mental models and usage strategies for DoubleAgents.
-Afterward, they completed another time of AI perception survey and a system experience survey assessing experience, workload, and attitudes toward specific components such as control, usefulness, and trust.
-Survey questions were adapted from NASA TLX (Hart, 2006) and prior AI perception and long-term evaluation studies (Ross et al., 2023; Long et al., 2024).
+Afterward, they completed another time of AI comfortableness survey and a system experience survey assessing experience, workload, and attitudes toward specific components such as control, usefulness, and trust.
+Survey questions were adapted from NASA TLX (Hart, 2006) and prior AI perception, reliance, and long-term evaluation studies (Ross et al., 2023; Basoah et al., 2025; Long et al., 2024).
 Finally, a short interview asked participants to reflect on their overall experience, including any insights or policies they would like the system to incorporate before Day 2.
@@
 Interview data were analyzed using thematic analysis to understand how users perceive the usefulness and trust drive about each feature, and quantitative data were analyzed using paired-samples t-tests to examine whether participants’ experiences and perceptions of the DoubleAgents and AI tools across usage and days.
-6.4. Result: Users’ Perceptions in AI Systems Shifted After Usages
-Figure 6.
-Average user perception scores toward proactive AI and comfort with automating the planning stage in seminar organizing, across three stages (average scores across 10 users: before Session 1, after Session 1, and after Session 2).
-denotes statistical significance at $p<0.1$; * at $p<0.05$; ** at $p<0.01$; and *** at $p<0.005$.
-Increasing Trust and Desire for Proactive Control.
-Figure 6(a) shows statistically significant changes in participants’ trust in AI systems to act on their behalf, from 3.20/7 (before using the system, based on their previous AI experiences), to 5.00/7 after two times of using DoubleAgents.
-P1 noted that even though DoubleAgents represented just one specific experience with a particular AI system, “I felt more optimistic… [and] in seeing similar systems like DoubleAgents in the future, I would generally trust such AI systems more.
-Additionally, there was a significant increase in wanting detailed explanations before the system acts, from 5.00/7 (before) to 6.00/7 (after).
-This suggests that seeing explanations from DoubleAgents made similar explanations from AI systems more desirable.
-P6 mentioned, “[With the system explanation] I felt like I could tell exactly what it’s doing at every stage so there’s a lot of transparency, [which] I think is pretty good… if I didn’t get that, I would feel lost [and] not know what’s going on…
-Figure 7.
-Average user perception scores toward comfort with automating the execution stage in seminar organizing, across three stages (average scores across 10 users: before Session 1, after Session 1, and after Session 2).
-denotes statistical significance at $p<0.1$ and * at $p<0.05$.
-Increasing Comfort with Automating the Planning Stage.
-Figure 6(b) shows participants’ attitudes towards proactive AI in planning became significantly more positive.
-All of these metrics showed statistically significant differences: they felt more comfortable letting AI visualize the current status (4.80/7 to 6.50/7), figure out potential time assignments in the planning (5.10/7 to 6.38/7), decide daily next steps (4.70/7 to 6.50/7), and act as an assistant to monitor or flag issues (5.50/7 to 6.50/7).
-For example, P6 found the Communication Visualization section of the interface useful in showing the full email exchange histories between the organizer and each speaker.
-She mentioned, “Everything is consistent inside the system—when some speakers were not followed up, I could map it back to the visualization and confirm accuracy… [It] helped me build trust… For tomorrow, I’ll spend less time checking the email replies [inside the chat interface] one by one, and instead [just check] the Communication Visualization on the side.
-For letting AI help monitor and point out potential issues, P1 described flagging as “one of the most important features of the system” and shared, “[With the issue flagging,] I skim [through the other features] now… It made me trust the system because I know it won’t just execute something random… It flags when it doesn’t know… That helps elicit my values and how I want to handle things—and [then DoubleAgents] respects that…
-Increasing Comfort with Automating the Execution Stage.
-Figure 7 shows that participants’ attitudes toward proactive AI in execution also became significantly more positive.
-For the tasks involving only drafting emails, as noted in the left column, all attitudes significantly improved, with comfort levels starting high and rising across the four types of emails (from around 6.0 before to nearly 7.0/7 after): initial outreach, follow-ups, clarification, and final confirmation.
-For the right column, draft-and-send tasks on users’ behalf, comfort also increased over time—rising for initial outreach from around 3.0/7 before to about 4.5/7 after, as well as for follow-ups and final confirmation.
-Participants demonstrated this after seeing DoubleAgents’ consistent performance over sessions: “I think [the system] acted quite reliably across two days, similar in terms of how it drafts emails [to the speakers], how it organizes language…” (P5).
-Participants also appreciated that the emails became more aligned to their tone and style over time—their email edits carried over to subsequent messages.
-For example, P4 mentioned that the system was learning her email writing style and that it remembered the phrasing she favored in email starters: “Compared to last time, I edited fewer times this time.,” which contributed to their trust in offloading more execution tasks to the AI, as well as a significantly increased desire for a system that aligns with and adapts to their styles over time (6.10/7 to 7.00/7, in Figure 6a).
-One notable exception is drafting-and-sending clarification emails without approval, which showed little change (3.00/7 to 3.90/7 to 3.90/7).
-P10 explained that clarification emails often require nuanced, user-specific context or subjective judgment: “For clarification, I might want to double-check… because clarification sort of means that AI is also encountering a situation that was usually not in the plan [or policies]… I’m not comfortable [offloading that] to AI… AI is already very good at pretending like it knows what it is talking about…” As a result, they preferred to remain in the loop, reviewing or guiding these clarification messages before sending.
-Additional Trust Driver: Simulation
-All 10 participants emphasized that the run-throughs of simulated test cases helped establish confidence in the AI system’s robustness, contributing to trust.
-They found the simulations useful for surfacing a variety of problems that were realistic and representative of real-life cases.
-For example, P3 mentioned “The simulations [over the 2 days] felt realistic and true to the person [they] were simulating, like their personality and emails… The simulation elicited possible problems that would happen in real life… Like [having] people asking for funding or [needing] to reschedule — it forced me to think about how I would handle that [in real life], and that was good.
-I can trust the system to handle that in the future for me if scenarios like that occur, since it already happens in the simulation.
-7. Deployment Study
-As the in-lab study demonstrated, the simulation was helpful, and participants felt comfortable trusting AI in planning and executing human coordination tasks.
-However, would real-world seminar organizers use DoubleAgents?
-If so, how would they use it—would they plan, draft, and send invitation, follow-up, or clarification emails in different ways?
-Would trust be established in a similar manner?
-What additional challenges might arise in real-world use?
-To explore these questions and evaluate the real-world deployability of DoubleAgents, we invited two organizers who were managing actual events during the preparation of this work.
-The goal was to observe how the system functions in realistic settings—including scheduling, tracking, email communication, and adaptive learning—and to understand how comfortable organizers are when the system sends real emails on their behalf.
-7.1. Deployment Users and Procedures
-Participants
-We recruited two experienced seminar organizers (1 male, 1 female; average age 31), both with a good general understanding of generative AI and experience using AI for work—one on a daily basis.
-They are PhD students who have experience organizing research seminars for their departments.
-In this deployment, each is responsible for managing speaker events with audiences of over 30 participants: D1 is coordinating the Fall 2025 HCI seminar series at a university, selecting 10 speakers from a potential pool of 23 to fill 10 slots.
-D2 is organizing a panel discussion for a conference workshop, inviting 5–6 panelists from a potential 18 to fill a single slot.
-Each organizer was introduced to the system and used it to plan and manage their events.
-Introduction, Setup, and Simulation Rounds
-Organizers were first informed about the study and provided consent.
-They were then introduced to the study procedures and the DoubleAgents system through a demonstration, which included planning–execution loops and trust mechanisms, illustrated with a simulation example from the user study.
-Organizers also completed the initial survey covering AI perception and demographics.
-To set up the system, organizers prepared a realistic start sheet by replacing simulated data with real organizing data (real potential speakers’ names, first with fake email addresses and later with the real ones) and provided their Gmail SMTP credentials.
-Clear instructions and consent procedures ensured again that participants understood data usage, privacy rights, and the option to withdraw.
-Before connecting their real email accounts, participants completed a 2-hour simulation round using pre-generated speaker personas based on the real potential speakers, allowing them to safely familiarize themselves with the system’s workflows.
-Finally, they tested the real email-sending module twice to confirm that emails were delivered correctly.
-Deployment
-After the simulation familiarization, organizers used DoubleAgents to initiate the coordination process and track it for over a few days.
-They continued receiving system updates over the following week.
-We communicated with participants consistently, logging interactions, usage patterns, and feedback on the system’s support for scheduling, tracking, and learning.
-Organizers’ system interactions—including email sends, scheduling updates, and system adaptations—were tracked.
-Qualitative feedback was collected via check-in meetings, focusing on usability, trust, concerns, and code support, while quantitative system logs were collected for system learning over the week.
-7.2. Deployment Results
-DoubleAgents’ Perceived Usefulness and Relevance in Live Deployment.
-After the demonstration and simulation run-through, both organizers (D1 and D2) reported that DoubleAgents was highly relevant to their seminar-organizing workflow.
-They highlighted that the system fit seamlessly into their email-heavy tasks and that control settings and edge-case flagging enabled handling complex coordination scenarios that conventional scheduling tools like Calendly cannot manage.
-D1 emphasized that crafting emails is a delicate and important task, as it involves personalizing messages in order to “make the speakers feel they are special,” even when requesting a favor.
-Both D1 and D2 found DoubleAgents capable of generating good starter emails for edits at the beginning, as well as producing an acceptable final email to send.
-Organizers specifically appreciated that the simulation included the names of the real speakers and incorporated aspects of their personalities in constructing the respondent simulations and edge cases, which helped them anticipate challenges in live deployment.
-System Onboarding, Testing, and Understanding in Live Deployment.
-D1’s deployment setup took four days and five meetings.
-The first two meetings were devoted to debugging and ensuring the system ran smoothly for a scenario involving more than 20 speakers, 10 slots, and substantial context information for each speaker.
-D1 modified their invitation and follow-up email templates more than four times.
-Additionally, D1 updated the seminar speakers’ descriptions to enable the email generation module to generate precise, high-quality emails that incorporated extensive context information.
-After hardening the system and refining our onboarding protocols, D2’s setup took two days and two meetings.
-D2’s application in panelists invitation were also simpler (e.g., shorter panelist descriptions, invitations for one-day panels), which contributed to the faster setup.
-Both D1 and D2 reported that the practice simulation rounds were valuable, and that sending testing emails through SMTP allowed them to gradually grasp the system’s capabilities and limitations.
-This illustrates that system setup and onboarding are consistent during deployment, while the required effort scales with task complexity and the amount of context data.
-Specifically, organizers emphasized the importance of supplying sufficient context to generate highly personalized emails.
-D1 mentioned, “[context management] is really tricky.” For example, one speaker D1 tried to invite had previously corresponded with a different organizer in a prior semester about her seminar availability on Fridays in a separate email thread.
-This semester, the seminar had shifted to Thursdays.
-To manage the context from past email threads, schedule updates, and templates, D1 opted to start a new thread with the speaker, as DoubleAgents at the time worked with new threads rather than replying within existing Gmail threads.
-In another case, D1 had already discussed seminar opportunities with a speaker on Discord but initially omitted this context during the early simulations.
-When composing the email, they noticed that the system had selected a generic template, so they updated the backend context with the appropriate tone and information to create a suitable follow-up.
-After practicing with real-speaker data in simulation, D1 finalized a preferred template for live sending.
-During the actual sending, they continued to refine messages—adding details and small fillers—and observed that these edits were retained and applied to subsequent emails.
-Email Initiation, Response Handling, and Edge Cases in Live Deployment.
-During live deployment, D1 sent emails to nine speakers.
-Over the earlier simulation rounds, she carefully perfected her initial templates and preferred writing styles via the implicit style transfer modules.
-Despite this, she still spent significant time refining the first few real emails.
-For the remaining speakers, she adopted a common real-world practice: copying previous templates into the text editor and adapting them for each recipient, which sped up the process.
-She mentioned that “the system’s email suggestions are helpful, though for similar invitation messages, I don’t always need a new draft each time.” D1 also used the system’s built-in AI chat feature seven times during initial email sending for phrasing recommendations.
-She noted that this support reduced mental demand and felt similar to receiving quick email feedback from a co-organizer, a practice she commonly uses in real-life organizing.
-D2, on the other hand, relied on the system’s automated drafting and style adaptation, demonstrating that trust and efficiency increased with experience.
-Interestingly, for some longer emails requiring scrolling to read, D2 skipped reviewing the full content and proceeded directly, trusting the system’s style transfer.
-At the end of the invitation email sending day, D2 sent emails to five potential panelists, relying directly on system-generated drafts based on his perfected templates and implicit style transfers from the previous simulation rounds.
-After the initial invitation email were sent, D1 received responses as quickly as 4 minutes, 1 hour, and 2 hours.
-As replies came in, organizers encountered more complex edge cases than those seen during simulation: all three responding speakers had overlapping availability on the same dates, creating interlocked scheduling conflicts.
-DoubleAgents logged the replies in the backend and suggested that D1 wait for the remaining six speakers while coordinating with the student speaker and the faculty member closest to her school to adjust availability.
-D1 instead opted to wait for three days before responding (one day more than the policy instructed), as she was occupied with other commitments and felt comfortable replying later.
-In contrast, D2 received three negative responses after a full day and did not encounter similar complications, and the system suggested waiting for the remaining two people to reply.
-Trust Establishment in Live Deployment
-Over time, D1 emphasized that trust depended on the system’s closeness to ground truth—specifically, whether emails were actually sent and whether responses were logged correctly and efficiently.
-Accurate alignment between the system’s state visualization and the real state of email threads was critical to their trust.
-Also, D1 noted that, over time, understanding the system’s capabilities and limitations helped her feel more comfortable using the tool.
-D2 reported that they appreciated having policies and rationale always visible and editable, in contrast to AI tools that hide their reasoning.
-After sending real emails, D2 remarked, “I trust AI for specific use cases more,” explaining that they found the system to be dedicated to and capable of generating high-quality email.
-At the same time, D2 compared DoubleAgents with other tools they were familiar with, abstracting their trust establishment process and the underlying barriers.
-He noted that the email interface differed from Gmail, introducing a small trust barrier.
-They compared Gmail’s richer text editor, easier insertion of tables and visuals, and familiar thread visualization—where they know what to expect when clicking—to our interface.
-They also expressed concern about formatting consistency for recipients (e.g., indentation, font sizes) in emails sent via DoubleAgents’ Gmail backend, noting that such issues can occur even within Gmail phone apps, which made them unsure how messages would appear to recipients.
-This concern was addressed through several rounds of email formatting and link embedding tests, during which they sent mails to their personal Gmails and checked them in parallel.
-Regarding integration, D2 felt that strong state visualization and tracking supported their trust through natural mapping.
-They believed their current practice of using spreadsheets to track speaker invitation progress could be integrated into the system’s visualization.
-For example, he usually rank the lists to monitor progress, and instead of serving solely as a display, the spreadsheet also functions as a goal-setting and action-taking tool.
-Thus, he expressed interest in making DoubleAgents’ visualization more interactive and editable, allowing them to directly initiate system actions.
-For actual adoption, he emphasized the need for scalability and further automation, requesting dynamic fields features similar to other email client tools like MailChimp.
-They also compared real email SMTP sending to Zoom microphone testing, emphasizing the value of having a trial, closed, and safe environment with strong validation—such as sending emails to themselves first.
-Trust Establishment Mechanism in DoubleAgents How Humans Build Trust With Humans Nielsen Heuristics (Nielsen, 1992) Human-AI Design Challenges (Yang et al., 2020)
-(1) Consistent and robust
-Policy-based conditional reasoning grounded in human social cognition; reliable handling of varied test suites and edge cases
-Trust grows when others are reliable, predictable, and follow shared norms or values.
-Consistency over time signals dependability.
-(Lansing et al., 2023; Rempel et al., 1985; Mayer et al., 1995; Beratan, 2007) - Consistency and standards
-- Error prevention
-- Help users recognize, diagnose, and recover from errors
-- Unpredictability of AI behavior: users need repeated, consistent outputs to form expectations
-- Data dependency: robustness across edge cases builds confidence
-(2) Controllable and aligned with user values
-User can confirm and intervene at every step; system adapts to users’ editing history and evolving policies
-Trust strengthens when people respect autonomy, allow mutual control and negotiation, and align with shared goals or values (Lansing et al., 2023; Kramer, 1999). - User control and freedom
-- Flexibility and efficiency of use
-- Conceptual misalignment: user oversight helps align system behavior with goals
-- Evolving mental models: adaptation supports shifting user practices
-(3) Understandable and explainable
-Real-time state tracking and visualization; clear reasoning and a chat window to ask questions
-Humans trust others who are transparent, communicate intentions, and explain their actions clearly (Lansing et al., 2023; Kramer, 1999). - Visibility of system status
-- Help and documentation
-- Match between system and the real world
-- Explainability gap: reasoning and state tracking reduce opacity and support accurate mental models
-Table 3.
-How DoubleAgents’ trust-building dimensions map to human–human trust building literature (Rempel et al., 1985; Mayer et al., 1995; Kramer, 1999; Lansing et al., 2023; Beratan, 2007), Nielsen heuristics (Nielsen, 1992), and human-AI interaction design challenges (Yang et al., 2020).
-8. Discussion: Trust Establishment with AI Agents Over Time, Consistency, Control, and Understanding
-DoubleAgents illustrates how users establish trust in a novel AI system over time.
-Users initially expressed low trust because they were unfamiliar with the system, but through exploration, testing, learning, and habituation, they gradually began to rely on it.
-This aligns with prior longitudinal studies on AI tools (Long et al., 2024).
-Understanding how this trust emerged sheds light on the mechanisms by which human–AI relationships can evolve.
-As shown in Table 3, DoubleAgents’ trust-establishment mechanisms center on three interdependent qualities: consistency, control, and understanding.
-These dimensions align with patterns of human–human trust development, are supported by classic usability heuristics (Nielsen, 1992), and address well-known challenges in human–AI interaction design (Yang et al., 2020).
-First, one main design goal of DoubleAgents is to be consistent and robust.
-The system employs policy-driven, conditional reasoning to stabilize its outputs and ground its behavior in shared human values and social norms.
-This design enables reliable performance even across diverse testing suites and edge cases, allowing users to form stable expectations—echoing how consistency signals dependability in human relationships and mitigating issues of unpredictability and data dependency in AI systems.
-Second, DoubleAgents incorporates mechanisms to remain controllable and aligned with user values.
-The system enforces stepwise confirmations and allows user intervention throughout planning and execution, while adaptively learning from users’ past editing histories and evolving policy additions.
-These mechanisms promote a sense of autonomy and mutual negotiation, reflecting how humans build trust through control-sharing, and addressing conceptual misalignments and evolving mental models in human–AI collaboration.
-Third, DoubleAgents is designed to be understandable and explainable.
-It offers real-time state tracking, clear visualizations, explicit reasoning for its actions, and a chat interface for user queries.
-These mechanisms increase system transparency and communicability, similar to how open communication fosters trust between humans, and reduce the explainability gap in AI systems.
-Collectively, these design mechanisms enabled users to gradually transition from skepticism to reliance.
-By operationalizing consistency, control, and understanding, DoubleAgents supported the iterative habituation process through which users internalized the system’s behavior and incorporated it into their workflows.
-This suggests that trust in proactive AI systems is not a fixed property but an emergent outcome—shaped by interaction designs that allow users to explore, test, and progressively align the system with their own goals and values.
-Several participants compared building trust in DoubleAgents to their prior experiences with coding automation tools such as Cursor and CLine, reflecting the CS background of most participants.
-For example, P10 described first trying the tool on a safe project in a controlled environment (not pushing to a live website) to see how well it worked, illustrating the role of consistent system performance in building trust.
-They also highlighted that Cursor offered strong control and state-reverting capabilities, which allowed them to confirm each step and explore outcomes safely.
-Over time, they learned how to craft better prompts, demonstrating alignment with their goals and values.
-Finally, participants emphasized the importance of understandable reasoning: both Cursor and DoubleAgents provide stepwise explanations for each action, which helped them make sense of the system’s behavior and develop confidence in its outputs.
-As P4 reflected: “Most of the trust I had with Cursor was because it would ask at each step and then explain its reasoning for that step—and it was accurate.
-The same thing here [in DoubleAgents]: it’s accurate, and each step provides reasoning.
-It’s from the frequent confirmation and the reasoning always being correct that I built trust.”
-These accounts illustrate that consistent performance, controllable interactions, and clear explanations collectively scaffolded the sophisticated trust formation process between AI and users, aligning users’ experiences with the mechanisms designed into DoubleAgents and with broader theories of habituation and technology acceptance.
-9. Limitations and Future Works
-First, DoubleAgents was initially evaluated through simulated seminar-organizing scenarios.
-While these simulations were reviewed by domain experts and deemed realistic, they may still lack the richness of real-world settings.
-Future work could make the ReAct-based coordination agent more user-aligned by incorporating users’ past email or organizing data and modeling their personal communication styles and values.
-Similarly, the respondent agent simulation could be made more realistic by drawing on a more diverse set of real human behavioral data.
-Second, our system used different models for different tasks—o3 for reasoning tasks and 4o-mini for other tasks.
-Although our technical evaluation showed they performed well, future models with stronger social and human cognition capabilities, as well as incorporating users’ data, may achieve better planning and execution.
-Third, although we deployed the system with two real organizers in real life, we also consulted others whose scenarios introduced additional complexities where the system may not perform perfectly.
-For example, one seminar organizer mentioned theirs involved bilingual invitations, and others described co-organizing dynamics, both of which introduce challenges not addressed in our current design.
-Finally, all our participants were experienced users of AI tools.
-Their familiarity may have positively biased their trust and comfort levels.
-Future work should include participants with a broader range of AI experience to better understand adoption barriers among less experienced users.
-10. Conclusion
-We present DoubleAgents, an agentic planning system that incorporates clear human intervention points, rich state visualizations, explicit policies, and uncertainty/edge‑case flagging to automate human coordination.
-In a two‑session lab study (n=10) and a week‑long deployment with two organizers, participants reported significantly greater comfort delegating planning and execution to DoubleAgents.
-Simulated respondents provided a safe proving ground to observe policy‑aligned planning and consistent system performance.
-From these studies, we distill three trust‑by‑design mechanisms—consistency, control, and understanding—operationalized via policy‑driven reasoning, stepwise approvals, implicit system adaptive learning, and a real-time system state.
-These patterns generalize beyond seminar scheduling to proactive agents design that must act while remaining auditable and aligned.
-Appendix A DoubleAgents Prompts
-A.1. Coordination Agent
-A.1.1. Progress Summary
+Appendix C DoubleAgents Prompts
+C.1. Coordination Agent
+C.1.1. Progress Summary
 You are an expert seminar organizer assistant.
@@
 ---------- [Task Rules] [Possible Dates] [Time Context] [Speaker Information] Here are the history of Approved Plans and Actions: [Approved Plans and Actions] -------------------------------- Here is the history of past user edits and implicit learning (recent, action edits only): [Implicit Learning] -------------------------------- Past Email Log: [Email Log]
-A.1.2. Policy Selection
+C.1.2. Policy Selection
 You are a helpful assistant.
@@
 [Implicit Learning/User Edit History (recent, action edits only)]
-A.1.3. Action Generation
+C.1.3. Action Generation
 You are an expert seminar organizer assistant.
@@
 -------------------------- Time Context: [Time Context] Task Progress Summary: [Progress Summary] -------------------------- Applied Coordination Policies: [Applied Policy] -------------------------- Implicit Learning/User Edit History (recent, action edits only): [Implicit Learning - Action Edits] --------------------------
-A.1.4. Email Drafting
+C.1.4. Email Drafting
 I am [Organizer Name], a seminar/workshop organizer.
@@
 [Implicit Learning - Email Edits]
-A.1.5. Implicit User Preference Learning
+C.1.5. Implicit User Preference Learning
 Plan and Action Edits
@@
 Inputs: history_context: {history_context}" "ai_text: {ai_text}" "user_text: {user_text}"
-A.2. Edge Case Detection
+C.2. Edge Case Detection
 You are an expert seminar organizer assistant.
@@
 Here are the policies: [Policies]
-A.3. Simulated Respondents
+C.3. Simulated Respondents
 You are simulating a seminar speaker.
@@
 Format your response as: ACTION: <reply|ask_question|wait> WAIT_UNTIL_DAY: <integer or ’none’> EMAIL: <JSON object or ’none’> RATIONALE: <your reasoning> EMAIL should be like: "from": "XX", "to": "XX", "subject": "XXX", "body": "XXXXX", "timestamp": "XX" dont output json ‘‘‘ etc
-A.4. Free-form QA Chat
+C.4. Free-form QA Chat
 You are an expert assistant.
@@
 [EMAIL LOG] [USER QUESTION]
-A.5. Daily Summary
+C.5. Daily Summary
 Today is simulation day [Day ID].
@@
 Dont include today’s date in the summary -- just summarize the progress.
-Appendix B Technical Evaluation Details
-B.1. Policy Selection
+Appendix D Technical Evaluation Details
+D.1. Policy Selection
 Here are the prompts used in the technical evaluations for policy selection:
-B.1.1. Experiment prompts
+D.1.1. Experiment prompts
 speaker-only
@@
 Main reasoning result: {progress_summary} Coordination Policies: {policy_list} Return ONLY the JSON list of indices:
-B.2. Edge Case Detection
-B.2.1. Test suite construction
+D.2. Edge Case Detection
+D.2.1. Test suite construction
 We employ GPT‑5 to generate a set of 10 edge-case (flaggable) test cases and 10 non-flagged (clear) test cases, with all examples validated for quality by the lead authors.
@@
 In total, 100 test cases are constructed and used as our test suite for edge case detection.
-B.2.2. Prompts to generate linguistic variations and test data
+D.2.2. Prompts to generate linguistic variations and test data
 You are a helpful assistant.
@@
 Please generate exactly five variations of the response for each case
-B.2.3. Experiment prompts
+D.2.3. Experiment prompts
 zero-shot

### v3

@@
-DoubleAgents: Interactive Simulations for Alignment in Agentic AI
+DoubleAgents: Human-Agent Alignment in a Socially Embedded Workflow
 Abstract.
-Agentic workflows promise efficiency, but adoption hinges on whether people can align systems that act on their behalf with their goals, values, and situational expectations.
-We present DoubleAgents, an agentic planning tool that embeds transparency and control through user intervention, value-reflecting policies, rich state visualizations, and uncertainty flagging for human coordination tasks.
-A built-in respondent simulation generates realistic scenarios, allowing users to rehearse, refine policies, and calibrate their use of agentic behavior before live deployment.
-We evaluate DoubleAgents in a two-day lab study (n=10), three deployment studies, and a technical evaluation.
-Results show that participants initially hesitated to delegate but used simulation to probe system behavior and adjust policies, gradually increasing delegation as agent actions became better aligned with their intentions and context.
-Deployment results demonstrate DoubleAgents’ real-world relevance and usefulness, showing that simulation helps users effectively manage real-world tasks with higher complexity and uncertainty.
-We contribute interactive simulation as a practical pathway for users to iteratively align and calibrate agentic systems.
-generative AI, AI agents, trust, reliance, transparency, state tracking, seminar organizing, control, planning, human coordination, reasoning, constitutional AI
-copyright: acmcopyrightjournalyear: 2025conference: 2026; 2026; arxivbooktitle: Proceedings of the 2026price: 15.00isbn: 978-1-4503-XXXX-X/18/06ccs: Human-centered computing Interactive systems and tools
+Aligning agentic AI with user intent is critical for delegating complex, socially embedded tasks, yet user preferences are often implicit, evolving, and difficult to specify upfront.
+We present DoubleAgents, a system for human–agent alignment in coordination tasks, grounded in distributed cognition.
+DoubleAgents integrates three components: (1) a coordination agent that maintains state and proposes plans and actions, (2) a dashboard visualization that makes the agent’s reasoning legible for user evaluation, and (3) a policy module that transforms user edits into reusable alignment artifacts, including coordination policies, email templates, and stop hooks, which improve system behavior over time.
+We evaluate DoubleAgents through a two-day lab study (n=10), three real-world deployments, and a technical evaluation.
+Participants’ comfort in offloading tasks and reliance on DoubleAgents both increased over time, correlating with the three distributed cognition components.
+Participants still required control at points of uncertainty — edge-case flagging and context-dependent actions.
+We contribute a distributed cognition approach to human-agent alignment in socially embedded tasks.
+generative AI, AI agents, trust, reliance, transparency, state tracking, seminar organizing, control, human coordination, constitutional AI
+copyright: acmcopyrightjournalyear: 2025conference: 2026; 2026; UISTbooktitle: Proceedings of the UIST 2026price: 15.00isbn: 978-1-4503-XXXX-X/18/06ccs: Human-centered computing Interactive systems and tools
 Figure 1.
-DoubleAgents is an agentic AI coordination tool that uses interactive simulation to let users experience and align the full ReAct cycle before live deployment.
-Starting from an initial setup (left)—including a simulation goal, simulated test-case personas, and an initial set of policies, and email templates—users can undergo multiple simulated ReAct cycles (middle) to rehearse planning, review and approve actions, execute plans, detect new edge cases, and iteratively refine policies via stop hooks, and email templates over repeated test cases and time.
-This process results in (right) the final speaker-slot allocation, as well as a user-aligned set of policies, emails, and stop hooks before the system sends real emails.
+DoubleAgents aligns agentic AI with user preferences for socially embedded coordination tasks such as organizing seminars.
+From an initial setup (left), including a goal, simulated personas, and seed policies and templates, users iteratively align the system (middle) by reviewing proposed plans, approving actions, handling edge cases, and refining policies, templates, and stop hooks across repeated runs.
+This produces (right) both a finalized speaker-slot allocation and a set of user-aligned artifacts — policies, email templates, and stop hooks — that persist and transfer to future coordination tasks.
 1. Introduction
-Figure 2.
-System diagram of DoubleAgents illustrating a day-by-day ReAct workflow that couples policy-guided planning with human oversight and LLM simulation.
-Inputs (top left) include the organizer’s goal, seminar slots, speaker personas, and a toolbox of callable functions.
-The coordination agent (center) iteratively: summarizes state; selects applicable policies; proposes a plan and action that the user can Regenerate or Approve; executes by drafting and sending emails, or waiting for responses.
-If a reply falls outside policy coverage, issueFlag escalates an edge case to the user for clarification before proceeding.
-The context management (right) maintains time information, past plans/actions, email logs, user preferences, and persona data, continuously updating to ground subsequent steps.
-Simulated respondent agents (bottom right) generate realistic, persona-consistent behaviors and replies that drive the loop across days.
-For a detailed walkthrough, please refer to Section 4.
-Diagram showing the architecture of the DoubleAgents system, with components for goal setting, agent coordination, context management (including edge case detection), and various submodules such as toolbox, orchestrator, and simulated respondents.
-Human–AI alignment is one of the defining challenges of the 21st century, especially for agentic AI: systems that go beyond answering questions to taking actions on users’ behalf (Leike et al., 2018; Soares and Fallenstein, 2014; Hendrycks et al., 2021).
-Following the Bidirectional Human–AI Alignment framework defined by Shen et al.
-(2025), we view alignment in agentic settings as an ongoing co-adaptation process between users and systems.
-On the system side, agents must be continuously shaped through interaction history and curated context so their actions reflect an individual user’s goals, preferences, and social norms (Zhang et al., 2025; Gao et al., 2024).
-On the human side, users must continually develop appropriate mental models, usage patterns, and AI literacy so they can effectively direct, evaluate, and calibrate delegation (Vanderlyn et al., 2025; Andrews et al., 2023).
-As agentic AI becomes more common, we need practical ways to assess and improve whether these systems remain aligned as they act in the world.
-Today, many popular agentic use cases, such as agentic coding or “vibe coding”, are relatively low-stakes and operate in closed sandbox environments (Horikawa et al., 2025).
-In these domains, alignment is often easier because systems can leverage large-scale, well-structured public data (e.g., GitHub), and success criteria are comparatively legible (e.g., whether code compiles, passes tests, or produces the desired output) (Chen et al., 2021; Jimenez et al., 2024).
-However, in more complex tasks that must adhere to human and domain knowledge as well as social cognition and norms, such as writing emails to real people to request favors, where etiquette is highly context-dependent, agents cannot reliably align on their own (Afroogh et al., 2024; Gabriel, 2020).
-These tasks often require nuanced, case-by-case judgment that is difficult to infer from generic training data alone and may demand substantial human oversight if deployed naively.
-One representative scenario is human coordination, such as scheduling workers with the right skills and experience for shifts they are available for or inviting researchers to a speaker series.
-In this work, we focus on the speaker-series scheduling setting.
-Organizers must secure a speaker for each slot (e.g., 4-5pm on Fridays), coordinate availability across multiple candidates, and handle uneven responsiveness (e.g., some invitees reply late or not at all).
-Responses can also introduce socially sensitive and context-dependent requests: “Can you reimburse travel?”, “Can I do Zoom?”, “I can’t make it on Fridays, what about Thursdays 4-5pm?” “Sorry for the last-minute cancellation—I now have a conflict,” or “My schedule is uncertain; can you follow up next month?” Determining the right next step is rarely optimal in a mathematical sense; it depends on the organizer’s values and preferences, which we refer to as “policies” (e.g., “Offer Zoom only if in-person is not feasible and no one else can do that slot,” “Prioritize senior faculty over PhD students when conflicts happen,” or “Follow up after two business days unless urgent”).
-In practice, because these “actions” are carried out almost entirely through email communication, the organizer must decide which email to send next (and to whom), how to phrase each email with the right content and right tone, and when to send it.
-While an LLM-based agent may generate plausible actions and write acceptable emails, it often lacks the organizer’s specific norms and cannot reliably anticipate how preferences should change across cases.
-Alignment therefore must happen on both sides: the system needs ways to acquire and apply three user-specific automation refinements: (1) coordination policies that guide planning and action choices, (2) email templates and writing style that match the user’s voice and social intent for each action, and (3) stop hooks that define when the agents need to pause and request for user confirmation; and the user needs support to form a clear mental model and specify these three preferences, especially early on, when they may not yet know what rules, tone, or escalation boundaries they want.
-The central challenge lies in how can we accelerate this bidirectional alignment so the agent can act more autonomously over time, without requiring constant babysitting.
-Our core approach is interactive simulation to support iterative automation refinement, enabling alignment between users and agents.
-As modern AI can simulate diverse human behaviors (Park et al., 2023, 2022), we use simulation as a testbed to expose users to realistic coordination scenarios, including problematic or edge-case behaviors, and to help them define and refine their communication strategies.
-During simulations, the agent proposes plans and actions (e.g., inviting multiple candidates, deciding when to wait versus follow up, or holding a slot for a high-priority speaker), and the user evaluates whether these choices are socially and contextually appropriate.
-This mirrors ideas from software engineering such as unit and integration testing: a system earns confidence not by succeeding on a single happy path, but by handling a range of representative cases.
-In our context, however, completing a task is not enough: agentic AI systems must be user-specific and value-aligned, reflecting personal priorities, norms, and preferences.
-We present DoubleAgents, a system for aligning agentic AI through interactive simulation.
-DoubleAgents integrates generative AI to simulate two complementary roles, as Figure 2 shows.
-First, it uses a ReAct loop (Yao et al., 2023) that pairs coordination agents for planning and execution.
-The agents can call tools to retrieve relevant context, query the current task state, and execute actions such as drafting and sending emails to potential speakers.
-We extend the loop with explicit interaction checkpoints so users can review, intervene, and approve decisions throughout planning and execution.
-Second, DoubleAgents includes an LLM-based simulator of respondents that generates diverse, realistic replies for different personas (e.g., a professor who responds slowly or a respondent who introduces funding or logistics constraints).
-This simulation enables repeated run-throughs of difficult cases, helping users understand how the agent behaves, stress-test policies against edge cases, and iteratively refine preferences and autonomy boundaries.
-Across interactive simulation runs, the primary outputs are not only a candidate schedule (speaker–slot allocation), but also a reusable set of user-specific alignment artifacts: refined policies, updated email templates, and calibrated stop hooks that can be applied when the system is deployed to send real emails.
-We evaluate DoubleAgents through a technical evaluation, a two-day lab study (n=10) in which participants used DoubleAgents with simulated respondents, and three real-world deployments.
-In the lab study, participants initially hesitated to delegate.
-However, over time, through multiple simulation rounds and the two study sessions, participants increasingly delegated, as visible state, policy-grounded reasoning, and clear intervention points (including stop-hook flagging) helped them calibrate autonomy and build confidence in the agent’s planning and execution.
-By the end, participants also produced reusable policies, stop hooks, and email templates aligned with their voice and style.
-Three week-long deployment studies demonstrate DoubleAgents’ usefulness and relevance for real-world, complex human coordination tasks.
-DoubleAgents was connected to organizers’ real email workflows and sending emails to real invitees, as well as incorporated incoming real speakers’ replies into its ongoing coordination state.
-Organizers reported that interactive simulation beforehand helped them prepare for deployment by crafting strong initial policies and email templates for real-world use.
-Our work demonstrates how agentic systems can support reliable, human-centered coordination by embedding simulation, transparency, and interactive oversight into the agentic AI workflow.
-Specifically, we contribute:
-  • A design perspective on alignment in agentic AI, realized through a policy-driven, intervention-friendly ReAct framework that supports transparency, human oversight, and alignment with human goals and values.
-  • An interactive simulation mechanism that models realistic human responses, enabling users to safely explore agent behavior, understand system dynamics, rehearse interventions, and iteratively refine policies, email references, and stop hooks before real-world application.
-  • Empirical results from a user study (n=10) and three real-world deployments, examining how interactive simulation and policy refinement support alignment in agentic AI for human coordination.
+Alignment between humans and AI is the foundation of useful agentic systems.
+To act on a user’s behalf, an agent must understand intent that is often implicit, evolving, and context-dependent — internalizing tacit norms of when (not) to act, when (not) to ask, and how to match a user’s style without being told each time.
+Without alignment, small misunderstandings scale into systemic problems across workflows and social relationships, eroding the trust that makes delegation possible.
+Even well-aligned AI still needs human oversight, both to correct errors and to guide new edge cases in complex, real-world settings.
+The end goal of alignment is not to fully replace human judgment but to let humans offload as much as possible to AI while maintaining control.
+How much is offloaded should be up to the human: the more aligned the system, the more confidently a user can delegate.
+But alignment cannot be specified upfront.
+Users often do not know what they want until they see the system act, and their preferences evolve as the task unfolds.
+The central question is how humans and AI can work together through repeated interaction and refinement to reach alignment and then comfortable offloading.
+Distributed cognition (Hollan et al., 2000; Hutchins and Klausen, 1996) offers a principled framework for designing this kind of human-AI partnership.
+Research on complex sociotechnical systems has shown that for many hard problems, like flying an airplane, no single individual is fully responsible for the cognition.
+Instead, cognition is distributed across humans and artifacts: the pilot, co-pilot, autopilot, cockpit instruments, ground control, and maintenance crews all contribute.
+Three components make this work.
+First, tracking state: the system must maintain a shared sense of where things stand, as cockpit instruments monitor altitude, speed, and position in real time.
+Second, representing information: critical state must be displayed in the right form so actors can make judgments at a glance, as cockpit displays show system overviews rather than raw data.
+Third, adapting policies: all actors follow procedures and checklists that encode accumulated knowledge about safe operation, as aviation checklists evolve after incidents to cover new situations.
+We apply this framework to design DoubleAgents, an interactive system for human-agent alignment in socially embedded coordination tasks.
+The name reflects the system’s two types of AI agents: a coordination agent that manages the task on behalf of the user, and simulated respondent agents — either real recipients in live deployment or simulated AI agents that stand in for real people during rehearsal runs.
+This dual mode lets users iterate on alignment through either live deployment with realistic feedback or simulation that compresses weeks into minutes before real social stakes are involved.
+Our domain is organizing a university speaker seminar series, a task where a human must monitor evolving situations over weeks, follow up with the nonresponsive, and handle unexpected scenarios, all while navigating real social stakes (you cannot unsend an email, nor be too pushy or too passive).
+DoubleAgents distributes cognition through three mechanisms: a coordination agent that tracks all state and proposes next actions, serving as the organizer’s memory and reasoning partner; a dashboard with temporal, social, and procedural views that make the agent’s reasoning legible so the user can verify and evaluate proposals; and a policy module that turns users’ direct edits into reusable alignment artifacts (coordination policies, email templates, and stop hooks) that improve over repeated runs through either live deployment or simulation.
+Overall, we make three contributions:
+  • A distributed cognition framework for human–AI alignment in socially embedded tasks, identifying three core components: tracking coordination state, externalizing information through visualization, and adapting behavior via policies, and articulating how these jointly support effective alignment.
+  • DoubleAgents, a system that operationalizes this framework through three components: (1) a coordination agent that maintains shared state and proposes actions, (2) a visualization layer that makes reasoning legible for user evaluation, and (3) a policy module that enables users to iteratively refine reusable alignment artifacts.
+The system supports both simulation for rapid iteration and live deployment for real-world coordination.
+  • Empirical insights from controlled studies and real-world deployments demonstrating the distinct role of each component, policies for shaping planning, visualizations for supporting evaluation, and stop hooks for governing execution, and showing that iterative alignment process produces reusable alignment artifacts that persist across sessions.
 2. Related Works
@@
 As AI agents take on more complex tasks, ensuring they align with human values has become a central research challenge.
-A key strategy is reinforcement learning from human feedback (RLHF), which fine-tunes models to produce outputs consistent with user expectations (Ouyang et al., 2022).
-However, alignment doesn’t end at training.
-Studies show that users often push back on AI responses that conflict with their values, steering behavior through real-time feedback, a process termed user-driven value alignment (Fan et al., 2025).
-Such interactive alignment is critical in high-stakes contexts like medical decision support, where LLM-powered co-pilot systems have helped clinicians catch more errors than either humans or AI alone (Ong et al., 2025).
-These findings show the importance of transparency, oversight, and human agency in AI deployment.
-Another approach, Constitutional AI, replaces constant human feedback with rule-based self-supervision.
-Models are trained to revise outputs using explicit principles like “Be respectful” (Bai et al., 2022).
-Recent work has expanded this into participatory frameworks such as Collective Constitutional AI, where large groups co-author alignment rules to reflect diverse values (Huang et al., 2024).
-Tools like ConstitutionMaker further help users turn natural language feedback into usable policies (Petridis et al., 2024).
-DoubleAgents contributes to this space by embedding alignment into interactive simulation.
-Users iteratively refine coordination policies, email tone, and stop hooks through rehearsed ReAct loops, treating alignment not as a one-time step, but as a continuous, participatory process.
-2.2. Agentic AI in Task Planning and Execution
-AI agents, systems that perform tasks on users’ behalf, operate through tool use (e.g., APIs) (Schick et al., 2023) or direct interface manipulation (Anthropic, 2025; OpenAI, 2025; Manus AI, 2025; Labs, 2025).
-While these agents show growing capabilities, delegating control to them still raises concerns around trust, transparency, and alignment (Afroogh et al., 2024).
-Planning and execution remain core challenges.
-Benchmarks like Tau-Bench, NaturalPlan, TravelPlanner, and AppWorld (Yao et al., 2024; Zheng et al., 2024; Xie et al., 2024; Trivedi et al., 2024) highlight the difficulty of robust multi-step planning, especially in open-ended tasks.
-To address this, methods like task decomposition (Zhang et al., 2024) and ReAct, which interleaves reasoning and acting, have been shown to improve agent adaptability (Yao et al., 2023), especially when used in collaborative or multi-agent domains (Lal et al., 2023).
-However, fully autonomous agents like AutoGPT often struggle with robustness and transparency in real-world use (Zheng et al., 2024; Yao et al., 2024).
-Cognitive research suggests humans rarely plan exhaustively upfront (Krieger et al., 2009; Teevan et al., 2016), pointing to a need for agents that support incremental and context-sensitive planning.
-Recent studies show that exposing agent reasoning and allowing user intervention improves both performance and trust (He et al., 2025).
-In sum, planning in agentic AI is not just a technical hurdle; it is a design problem.
-Transparent reasoning, iterative refinement, and human-in-the-loop oversight are essential for building trustworthy agents that can meaningfully support real-world tasks.
+Following the Bidirectional Human–AI Alignment framework (Shen et al., 2025), we view alignment in agentic settings as an ongoing co-adaptation process.
+On the system side, agents must be continuously shaped through interaction history and curated context so their actions reflect a user’s goals, preferences, and social norms (Zhang et al., 2025; Gao et al., 2024).
+On the human side, users must develop appropriate mental models and usage patterns to effectively direct, evaluate, and calibrate delegation (Vanderlyn et al., 2025; Andrews et al., 2023).
+Operationalizing this bidirectional view requires mechanisms that let users specify intent, oversee execution, and revise expectations over time.
+Terry et al.
+(2023) decomposes this challenge into specification, process, and evaluation alignment, capturing the full arc from goal communication through output verification.
+Subsequent work has addressed each stage: Shankar et al.
+(2024) shows with EvalGen that users’ evaluation criteria themselves drift as they observe model outputs, while Epperson et al.
+(2025) supports process-level steering with AGDebugger, enabling what-if probing and checkpoint-based resets to diagnose misalignment in multi-agent workflows.
+Together, these efforts outline a design space where guardrails, prompt optimization, and interactive debugging serve as complementary post-hoc alignment mechanisms.
+Constitutional AI offers a different strategy, replacing constant human feedback with rule-based self-supervision using explicit principles (Bai et al., 2022).
+This has been expanded into participatory frameworks such as Collective Constitutional AI, where large groups co-author alignment rules to reflect diverse values (Huang et al., 2024), and ConstitutionMaker, which helps users turn natural language feedback into usable policies (Petridis et al., 2024).
+DoubleAgents contributes to this space by embedding alignment into an interactive process.
+Following Suchman’s view of plans as situated resources rather than fixed programs (Suchman, 1987), policies and stop hooks in our system function as evolving artifacts that users negotiate in context.
+Through interactions, users iteratively refine coordination policies, email tone, and stop conditions, treating alignment as a continuous, participatory process shaped by the situations in which it unfolds.
+2.2. Agentic AI: From Task Planning to Socially Embedded Workflows
+Delegating complex tasks to AI agents raises concerns around trust, transparency, and alignment (Afroogh et al., 2024), motivating designs that keep humans in the loop.
+Exposing agent reasoning with user intervention improves both performance and trust (He et al., 2025), yet cognitive research suggests humans rarely plan exhaustively upfront (Krieger et al., 2009; Teevan et al., 2016), pointing to a need for agents that support incremental, context-sensitive planning with human oversight.
+Whether agents operate through tool use (Schick et al., 2023) or direct interface manipulation (Anthropic, 2025; OpenAI, 2025; Manus AI, 2025; Labs, 2025), the difficulty of robust multi-step planning has been consistently highlighted in recent benchmark works (Yao et al., 2024; Zheng et al., 2024; Xie et al., 2024; Trivedi et al., 2024), underscoring that oversight is not merely desirable but necessary.
+These challenges intensify when agents operate in socially embedded contexts, where correctness depends on interpersonal norms and situated judgment rather than functional specifications.
+Benchmarks for social intelligence reveal a wide capability gap: SOTOPIA (Zhou et al., 2024) found that LLM agents significantly underperform humans on social commonsense, while CASA (Qiu et al., 2025) showed high norm violation rates in agentic settings, suggesting task-directed focus actively crowds out social sensitivity.
+DoubleAgents targets the gap this literature exposes: the absence of end-to-end systems for multi-turn, socially embedded coordination.
+Rather than maximizing autonomy, our system helps users iteratively author and refine agent behavior through calibrating delegation boundaries, coordination policies, and communication tone, for tasks where social context and timing are integral to correctness.
 2.3. Human Simulations with LLMs
-Recent work has shown that LLMs in multi-agent simulations can exhibit surprisingly human-like behaviors, forming routines, organizing events, and coordinating spontaneously without explicit instructions (Park et al., 2023; Horton, 2023; Park et al., 2022).
-These behaviors often reflect human reasoning and social strategies, aligning with patterns observed in economic games (Sreedhar et al., 2025; Sreedhar and Chilton, 2025; Aher et al., 2023), judicial decisions (Hamilton, 2023), and over 80% of psychology experiments (Hewitt et al., 2024).
-This suggests LLM agents can serve as useful proxies for modeling human behavior at scale.
-Beyond simulation, LLMs are demonstrating real-time social competence.
-CICERO, a dialogue-enabled Diplomacy agent, ranked in the top 10% of human players by integrating language with strategic planning (Team et al., 2022).
-Similarly, the MATRIX framework used simulated AI dialogues to surface moral tradeoffs in Constitutional AI evaluations (Pang et al., 2024), showing how LLMs can assist in value alignment.
-LLM simulations also support red teaming and adversarial testing, enabling safe, scalable evaluation of edge cases and vulnerabilities (Purpura et al., 2025).
-While imperfect, these agents provide a reproducible and low-risk method for probing system behavior in socially sensitive contexts, offering a powerful complement to human-centered alignment and evaluation.
+Recent work has shown that LLM-based multi-agent simulations can exhibit human-like behaviors in social and coordination settings, forming routines, organizing events, and interacting in ways that resemble human reasoning patterns (Park et al., 2023; Horton, 2023; Park et al., 2022).
+These simulations have been used to approximate human behavior in domains such as economic decision-making (Sreedhar et al., 2025; Sreedhar and Chilton, 2025; Aher et al., 2023), judicial reasoning (Hamilton, 2023), and psychology experiments (Hewitt et al., 2024).
+They have also supported applications such as strategic gameplay (Team et al., 2022), adversarial testing (Purpura et al., 2025), and value alignment through simulated interactions (Pang et al., 2024).
+In DoubleAgents, LLM-simulated respondents are used to provide a fast and flexible way to surface a range of coordination scenarios, including potential edge cases, so that users can iteratively refine their delegation strategies.
+These interactions primarily serve to accelerate the alignment process by making system behavior observable early, while similar workflows can also be carried out during real-world deployment with actual participants.
 3. DoubleAgents System
 Figure 2.
-A screenshot of DoubleAgents coordinating the assignment of four speakers to four seminar slots.
-(A) Policy Panel – Displays coordination policies that reflect user values and preferences.
-Users can add, edit, or delete policies, which guide the reasoning and planning of the coordination agent.
-(B) Interactive Chat Interface – Shows the full history of plans and actions, allowing users to monitor planning and execution, confirm or modify agent suggestions, and intervene when necessary.
-(C) Plan and Action Pop-up – A focused view within the chat history where users can review, edit, regenerate, or approve individual plans and actions.
-(D) Assignment Tracker – Visualizes tentative speaker-to-slot assignments and indicates each speaker’s availability.
-Dark purple denotes confirmed availability; light purple signals ambiguous availability requiring user confirmation.
-(E) Communication History Viewer – Displays the messaging history between the coordination agent and each speaker, organized by date and interaction.
-(F) Calendar View – Provides temporal context by showing the current simulation and upcoming seminar dates, color-coded for clarity.
-We present DoubleAgents, a system for aligning agentic AI through interactive simulation.
-Built on the ReAct framework (Yao et al., 2023), DoubleAgents supports iterative planning and execution via a team of AI agents that collaborate with users on complex coordination tasks.
-In this paper, we focus on a concrete use case: assigning seminar speakers to available time slots (see Appendix A for additional background).
-DoubleAgents takes as input a user-defined goal (e.g., “Organize a seminar with four speakers and four slots”), persona descriptions for simulated respondent agents, the user’s existing email templates, and a set of generic scheduling policies.
-Through iterative simulations, the system produces a finalized assignment of speakers to slots, along with refined policies, updated email templates, and executable stop hooks that users can later deploy in real-world agentic workflows.
-The system comprises six core components organized into three functional categories.
-First, DoubleAgents includes the Core ReAct Agents, which implement the reasoning, planning, and simulation logic that drives the coordination process:
-  • Coordination Agent - Responsible for reasoning over the current problem state using a set of predefined policies. It proposes plans and actions, and invokes appropriate tools to execute them.
-  • Simulated Respondents - Represent background agents that emulate human respondents’ behavior by responding to user-generated emails based on predefined personas and descriptions.
-Second, DoubleAgents incorporates the System Orchestration Modules, which manage the internal context and ensure robustness during decision-making:
-  • Context Management Module - Maintains all relevant contextual information related to the task, including simulated respondents’ details (e.g., personas and descriptions), email logs from prior time steps, and the history of user-approved plans, actions, and emails.
-  • Edge Case Detector - Identifies situations that fall outside the coverage of existing policies and escalates them to the user for clarification and additional instructions.
-These escalations are recorded as stop hooks, enabling iterative refinement and extension of the policy set.
-Lastly, the User Interaction Interfaces components enable transparency, user control, and real-time feedback during the coordination process:
-  • Interactive Chat Interface - Facilitates communication between the user and the system.
-Users can approve or revise generated plans, actions, and emails, and respond to flagged edge cases by providing natural language feedback.
-  • Visualization Dashboard - Provides real-time feedback on the state of the task, including temporal progress, assignment status, and an activity tracker to help users monitor the overall coordination process.
-In this section, we provide a walkthrough example to illustrate how DoubleAgents operates in practice.
-We also describe the implementation details of DoubleAgents.
-Consider the example of Lee, a PhD student organizing the HCI seminar series at his university.
-Lee plans to invite four speakers–Prof.
-A (a famous professor), Prof.
-B (a rising star professor), Student X (a senior PhD student), and Student Y (a PhD student who is a friend of Lee) – to deliver talks during four available time slots: Sep 5, Oct 4, Nov 7, and Dec 5.
-Lee begins at the landing page, where he specifies his goals to the system.
-He types: “I am trying to organize a seminar with four speakers and four slots, help me reach out”.
-He then provides additional details, including the seminar location, and the dates of the four time slots.
-Next, Lee enters persona information for each speaker he wants to invite, which the system uses to generate simulated speaker profiles.
-The landing page is shown in Figure 4.
-To further customize communication, Lee uploads five of his past emails, allowing the system to adapt its generated messages to his style; for instance, casual for friends and formal for faculty.
-Finally, Lee initiates the simulation by clicking the “Start Planning” button, prompting the system to set up the scenario with the specified speakers and time slots.
-At this point, he is also shown a list of preset policies (see Figure 3(A)), which serve as guidelines for the coordination agent to begin planning.
-Lee reviews the policies, all of which appear reasonable.
-He may modify existing policies or introduce new ones to better guide the agent’s planning process.
-Figure 4. The landing page to specify the user goals, seminar details, and the persona info for the speakers.
-3.1. Context Management Module
-At the beginning of the scheduling process, all information provided by Lee on the landing page is compiled into a structured problem context that the system consults at each step to solve the assignment task.
-DoubleAgents represents this problem context using JSON structures, where each key denotes a specific type of contextual information (e.g., speaker profiles, time constraints, email history, etc.), and each value contains the corresponding content.
-As shown in Figure 2, an example snapshot of the problem context includes keys such as ”time info”, which reflects the seminar slots specified by Lee.
-The context is dynamically updated throughout the interaction, incorporating user feedback, agent actions, and simulated responses to maintain an up-to-date representation of the task state.
-3.2. Interactive Chat Interface
-Upon clicking “Start Planning,” Lee is directed to the main Interactive Chat Interface (Figure 3(B)), where he can view the full history of system interactions and monitor the planning and execution progress carried out by the core ReAct agents.
-All plans, actions, and emails generated by the system are accessible through this interface.
-Each system update appears as an interactive pop-up window (Figure 3(C) shows a pop-up plan item), enabling Lee to inspect and confirm the system output in a controllable and traceable way.
-The interface also supports natural language queries, giving Lee full control over what to view.
-For example, he can type “What did Prof.
-B reply in the last email?” to retrieve the relevant message history directly within the chat.
-This chat functionality is powered by GPT-4o, with prompt details provided in the Appendix C.4.
-(a) The pop-up window for a generated plan.
-(b) The pop-up window for generated actions following the plan.
-(c) The pop-up window for a generated email.
-Figure 5. Examples for plan and action generation.
-3.3. Coordination Agent
-3.3.1. Progress Understanding
-When Lee initiates the planning process, the coordination agent assesses the overall state of the problem by generating a concise summary of the progress so far.
-To do that, it first reviews the current state of the scheduling process by analyzing the problem context, including speakers’ info, email logs from previous days, and the history of plans and actions previously approved by Lee.
-Based on this information, the coordination agent generates a summary of the current status, indicating which days have been scheduled, what actions have been taken, and the most recent contact date for each speaker to help monitor responses and follow-ups.
-The prompt used in generating the summary is illustrated in the Appendix C.1.1.
-3.3.2. Policy Selection
-To design a set of general policies applicable to a broad range of planning problems, the lead authors brainstorm and define rules tailored to address common scheduling challenges.
-These challenges include: (1) individuals being reluctant to share their full availability, (2) ambiguous or unclear responses, (3) the need for multiple follow-ups to elicit responses, (4) resolving scheduling conflicts between two speakers, and (5) initiating the process by broadly reaching out to potential participants.
-Based on these identified challenges, we developed six initial ”seed policies” as a core set of guidelines for the system to reference and select from when generating plans.
-Each policy is designed to address a specific challenge, with two policies dedicated to handling non-responsive participants: one focused on strategies to cope with the lack of response (follow-up frequency) and another providing a framework for when to give up on follow-ups.
-In Lee’s example, since it is the first day of scheduling, the progress summary informs the system that no speaker availabilities have been collected yet.
-As a result, the coordination agent invokes the policy selection tool, which identifies the appropriate policies based on the current problem state.
-Drawing from the list of ”seed policies”, the system selects: “If we are just beginning and have not yet contacted any speakers for availability, then ask each speaker for their availability across all slots.” The selected policy is then highlighted in the Policy Panel (as shown in Figure 3(A)), providing Lee with the rationale to generate the plan.
-The policy acts as the planning guideline, informing the agent’s next steps.
-Details of the prompts used for the policy selection module are provided in the Appendix C.1.2.
-3.3.3. Plan Generation
-Guided by this policy, the coordination agent generates the following plan (shown in Figure 5(a)) for Lee: “Send email to Prof.
-A, Prof.
-B, Student X, and Student Y to request availability for slots on Sep 5, Oct 4, Nov 7, and Dec 5.”.
-Lee considers the plan and the rationale logical, since scheduling cannot proceed without first collecting availability information, and clicks ’Approve & Continue’ to move forward.
-The plan and its rationale are automatically recorded in the interactive chat interface (shown in Figure 2(B)), allowing Lee to review past activities and understand the rationale behind the plan.
-If Lee is dissatisfied with the proposed plan, he can either request the system to regenerate an alternative or directly edit the current plan himself, as shown in Figure 5(a).
-3.3.4. Action Generation
-To operationalize the plan, the coordination agent then generates a set of concrete actions to be executed (see Figure 5(b)).
-Specifically, it proposes sending individual emails to each speaker, with the stated goal: “Request for availability for Seminar Slots (Sep 5, Oct 3, Nov 7, Dec 5).” These proposed actions are presented to Lee for review.
-He examines each action and its corresponding goal, deems them appropriate, and clicks Approve to proceed.
-Importantly, Lee retains full control over the suggested actions: he may edit the goals, remove actions that are misaligned, or instruct the system to regenerate the set based on revised input.
-In general, there are three types of actions: (1) drafting emails to selected respondents, (2) sending emails to specified recipients using their provided addresses, and (3) awaiting responses from certain respondents.
-The detailed prompts used for plan and action generation are provided in the Appendix C.1.3.
-3.3.5. Action Execution
-Following approval, the coordination agent proceeds to instantiate each action by generating email drafts for the designated speakers, including both subject lines and message bodies (Figure 5(c)).
-This is done by calling the email writing tool, as specified in the tool box shown in Figure 2.
-Specifically, it calls the writeEmail() API, which drafts emails given the current problem state, seminar details, user-provided templates, and historical email edits.
-These elements allow the system to learn and adapt to the Lee’s preferred communication style, capturing tone, level of formality, and domain-specific phrasing, when generating new emails.
-Note that email drafting is performed using GPT‑4o, which is invoked by the coordination agent during the execution of the writeEmail() tool.
-Lee reviews each draft and identifies one instance where the tone is not appropriate—the email to Prof.
-A, a distinguished professor, feels overly informal.
-To address this, Lee types: “Please make the email sound more formal, as Prof.
-A is a distinguished professor I’d like to invite.” The system then regenerates a more suitable version.
-Lee further refines the message by editing the first paragraph to better reflect his personal voice: “My name is Lee, a PhD student in Computer Science.
-Your work in [XXX] has greatly inspired us, and I would be honored to invite you to speak at our HCI seminar series.” After reviewing and finalizing all drafts, Lee is satisfied and clicks ‘Approve & Send’, prompting the system to dispatch the emails to all invited speakers.
-This is accomplished by invoking the sendEmail() API by the coordination agent, targeting the appropriate speaker email addresses specified in the problem context.
-In addition, all email-related activities are logged in the dashboard (Figure 2(E)), enabling Lee to easily monitor the status and history of interactions with each speaker.
-Besides writing and sending emails, the coordination agent may invoke actions involving delayed responses, such as waiting for a speaker’s reply.
-This is achieved by calling the waitResponse() tool.
-In such cases, no message is generated or sent; instead, the agent pauses execution and either proceeds to the next pending action or advances to the next simulated day.
-Overall, we use GPT-o3 as the backbone of the coordination agent, given its strong reasoning capabilities.
-3.3.6. Implicit User Preference Learning
-DoubleAgents supports implicit learning of user preferences by observing user-initiated edits to system-generated plans, actions, or emails.
-For instance, if the coordination agent proposes to reach out to Student X to ask for more availability, Lee may choose to remove the action “Send initial email to Student X”—perhaps because Student X’s prior responses have been vague or unhelpful.
-The system treats such edits as preference signals.
-To act on these signals, the coordination agent summarizes the edit into a concise statement, logs it in a user edit history, and incorporates it into the problem context.
-This updated context is then referenced in subsequent planning steps, enabling the agent to evolve its behavior in alignment with the user’s preferences over time.
-To streamline the summarization process, we leverage GPT-4o.
-Details of the prompt used for generating summaries of user edits across various content types can be found in Appendix C.1.5.
-The user-initiated edits, the corresponding system-generated drafts, and the generated summaries of the edits are all preserved in the user edit history.
-This comprehensive history serves as an evolving knowledge base, directly influencing the system’s generation of future plans, actions, and emails, and ensuring that it remains responsive to the nuances of user behavior.
-3.4. Simulated Respondent Agent
-To construct a realistic simulation environment for Lee, we draw inspiration from (Park et al., 2023) and employ LLMs to simulate each respondent agent, generating actions that align with their assigned personas.
-Concretely, upon receiving Lee’s emails, each simulated respondent agent first reasons over its persona and then generates an action that adheres strictly to those characteristics.
-For instance, the agent representing Prof.
-A opts not to respond on the same day (shown in Figure 2), as his persona specifies: “You are very busy and reply very late.
-Normally, it takes you three follow-up emails to respond.” In contrast, Prof.
-B’s persona indicates she is prompt in communication, leading her agent to reply immediately.
-Given that Prof.B’s persona also specifies preferences, such as “prefers presenting over Zoom”, the generated action will explicitly request a Zoom presentation option.
-The prompt for speaker agent, both for generating actions and drafting email responses, is detailed in the Appendix C.3.
-To execute the selected actions, the simulated respondent agents compose and send email responses, which are then received by Lee on his end.
-Figure 6(a) shows an example response received by Lee as shown in the interface.
-These responses advance the overall scheduling process and enable Lee to continue coordinating subsequent steps.
-Similar to the coordination agent, we adopt GPT-o3 as the backbone of each respondent agent to perform reasoning and generate plans and actions.
-GPT-4o is used when invoking the sendEmail() tool to generate email responses.
-(a) Email sent by the simulated Prof. A.
-(b) Edge case detection given response from Prof. B asking about Zoom options.
-Figure 6. The example email sent by the simulated respondents and the system message when detecting edge case.
-3.5. Edge Case Detector
-Within a day, Lee receives an email from Prof.
-B stating: “Thanks for reaching out, Lee!
-I was wondering if there might be an option to present via Zoom instead of in person?
-That would provide more flexibility for me.” As this type of request had not been previously addressed by any existing policy, representing an edge case, the edge case detector issues a flag to prompt clarification from Lee.
-Through the Interactive Chat Interface, Lee sees the following system message: “Prof.
-B requests a Zoom option due to circumstances preventing in-person attendance.
-Should I confirm Zoom availability for their seminar or insist on in-person participation?” Preferring in-person interactions for seminars, Lee responds with the instruction: “Tell Prof.
-B we prefer in-person due to the potentially large audience.” The coordination agent then processes this clarification and generates a new action to inform Prof.
-B accordingly.
-Lee reviews the suggested action, finds it appropriate, and clicks Approve.
-An email draft is subsequently generated; after confirming the message aligns with his intent, Lee clicks ’Approve & Send’ to dispatch the response to Prof.
-B.
-Lee’s clarification is recorded as a stop hook and used to refine the policy set, explicitly specifying that Zoom is not preferred when speakers request remote participation.
-In this way, user interventions during edge cases are converted into reusable system policies.
-In general, the edge case detector acts as a tool for the coordination agent to check if the message sent from the respondent agent can be addressed with the existing set of policies in order to make the next plans.
-When the detector identifies an edge case, the coordination agent escalates the issue to the user by presenting a clarification question in the interactive chat interface, prompting the user to provide guidance on how to proceed.
-The edge case detector is implemented with GPT-4o (see prompts in the Appendix C.2).
-3.6. Dashboard Visualization
-To support user understanding and foster alignment, DoubleAgents provides clear visual feedback on the task state and system progress throughout the scheduling process.
-We implement a Flask-based dashboard interface designed to make the planning workflow interactive, legible, and actionable.
-The dashboard is composed of several layered visual components, shown in Figure 3(D), (E), and (F), which we describe in detail below:
-  • Assignment Grid and Progress Tracker (Figure 3D): This panel visualizes tentative and confirmed speaker-to-slot assignments.
-It highlights which seminar slots remain unfilled and which speakers have been successfully scheduled.
-This layer helps Lee assess how close the current plan is to completion at a glance.
-  • Calendar and Timelines (Figure 3F): Anchored in simulated days, this temporal layer helps Lee reason about timing, deadlines, and follow-up actions.
-The calendar highlights both the current simulated day and the seminar dates, allowing Lee to prioritize actions based on urgency and scheduling constraints.
-  • Communication Visualization (Figure 3E): This panel displays the message history between the coordination agent and each speaker, offering a transparent view of the outreach and response status.
-It enables Lee to track ongoing communications, detect gaps, and assess follow-up needs.
-These visual layers operate in parallel, allowing the system to efficiently track and update the state of each element in real time.
-Throughout the process, Lee interacts with the coordination agent and the respondent agents to collaboratively establish speaker-to-slot assignments for the seminar.
-Together, these visualization features provide Lee with a clear overview of assignment progress and a sense of what needs to be done on the current and subsequent days.
-At the end of each simulated day, the system generates a daily summary that highlights key events and decisions.
-Specifically, we use GPT-4o to generate the daily summaries, with prompt details provided in the Appendix C.5.
-As shown in Figure 3(B), one example summary is presented within the Interactive Chat Interface and reads: “Invitations for the HCI seminar were sent to Profs.
-A, B, and Students X, Y.
+A snapshot of DoubleAgents coordinating four speakers across seminar slots.
+Grounded in distributed cognition, the system organizes interaction across three components: (a) Policies - encoding evolving user preferences and procedures; (b) The Coordination Agent - tracking state, plans next steps, suggests actions, and surfaces the policies guiding its decisions; and (c) The Dashboard Visualization - externalizing key information, including assignments, timelines, and communication history, enabling users to quickly interpret the situation.
+In use, the coordination agent monitors progress, proposes a plan, and explains its rationale; the user reviews the state through the dashboard, verifies that the plan aligns with expectations, and approves it to proceed.
+Through interactions, policies are refined, allowing the system to better align with the user over time.
+DoubleAgents is an interactive system for aligning agentic AI behavior with user preferences in coordination tasks.
+We ground the system in a concrete use case: organizing a university speaker seminar series (see Appendix A for additional background), a setting where users must manage evolving constraints, follow up with nonresponsive participants, and handle unexpected situations under real social stakes.
+Consistent with our distributed cognition framework, DoubleAgents distributes coordination across three components: a coordination agent that tracks task state and proposes next plans and actions, a dashboard that presents temporal, social, and procedural views to make the agent’s reasoning legible for user evaluation, and a policy module that turns user edits into reusable alignment artifacts (e.g., coordination policies, email templates, and stop hooks).
+Together, these components support iterative refinement of agent behavior over time, enabling users to progressively align the system through interaction in both simulated and real-world deployment settings.
+In this section, we first present a user walkthrough (Section 3.1), followed by the system implementation (Section 3.2).
+3.1. User Walkthrough
+To illustrate how DoubleAgents operates in practice, we walk through a concrete example.
+Consider Lee, a PhD student organizing the HCI seminar series at his university.
+Lee plans to invite four speakers, Prof.
+Adrian Smith (a famous professor), Prof.
+Lucy Tran (a rising star professor), Michael Chen (a senior PhD student), and Sarah Malik (a PhD student who is a friend of Lee) to deliver talks during four available time slots: Sep 5, Oct 4, Nov 7, and Dec 5.
+We demonstrate three example cases of how Lee uses DoubleAgents to solve the problem: (a) starting email bunch, (b) handling edge cases, and (c) following up with non-responsive speakers.
+3.1.1. Setting up
+Lee begins at the landing page (Figure 6), where he specifies his goal: “I am trying to organize a seminar with four speakers and four slots—help me reach out.” He then provides additional details, including the seminar location and the dates of the four time slots.
+Next, Lee enters information for each speaker, such as their names and email addresses.
+When operating in a simulated setting, Lee can optionally include brief persona descriptions to approximate the behavior of the speaker.
+For example, Prof.
+Adrian Smith is described as “a famous professor who is very busy and often does not reply to emails,” while Sarah Malik is “a friend who is generally responsive.”.
+In real-world deployment, these personas are not required, as interactions occur with actual participants.
+Finally, Lee uploads several past emails he has sent to seminar speakers.
+This allows the system to adapt to his communication style, generating messages that are more casual for friends and more formal for faculty.
+With this information, the system could send emails to speakers on Lee’s behalf.
+But Lee is nervous about this.
+He wants to make sure the system sends well-written, personalized emails; follows up appropriately if speakers do not respond; nudges people without being pushy; fills all four slots without double-booking; and, critically, does not offend a famous faculty member with an overly casual message.
+Lee does not yet know how the system will behave, so rather than letting it act in the real world, he starts a simulation that will not send any actual emails.
+Instead, the simulation lets him see exactly how the system reasons, what plans it proposes, and how it responds to different speaker behaviors, all guided by a set of policies he can inspect and refine.
+Lee is also shown a list of preset policies in the Policy Panel (Figure 2(Left)).
+These are natural-language rules that will guide the AI’s reasoning: for example, when to follow up, how to handle conflicts, and when to give up on a non-responsive speaker.
+Lee reviews them, finds them reasonable, and is now ready to start the simulation.
+3.1.2. DoubleAgents starts and policy selection.
+When the simulation starts, Lee sees the Control Panel (Figure 2(Middle)) alongside a dashboard showing the assignment progress, calendar, and communication history (Figure 2(Right)).
+The calendar shows that in the simulation, today is August 5, and the first seminar date to fill is September 5, so he should start reaching out now.
+Obviously, nobody has been contacted yet.
+Lee does not have a clear strategy for who to email first or what to say, so he lets the AI suggest a plan.
+The coordination agent assesses the current state: no speakers emailed, no availability collected, four slots to fill, and looks for a relevant policy in the existing policy bank, which contains “If someone hasn’t responded to the email after three reach-outs, then don’t contact them again.”, “Don’t be shy to ask for more availability if they don’t give you many.
+They will probably reply more if you beg.” and a few others.
+Then the agent finds one that’s most relevant: “If we are just beginning and have not yet contacted any speakers for availability, then ask each speaker for their availability across all slots.” The selected policy is clearly highlighted to Lee in the Policy Panel (Figure 2(Left)), so Lee can see exactly why the AI is about to do what it does.
+Based on this policy, the AI subsequently proposes a plan (Figure 7(a)): “Send email to Prof.
+Adrian Smith, Prof.
+Lucy Tran, Michael Chen, and Sarah Malik to request availability for slots on Sep 5, Oct 4, Nov 7, and Dec 5.”
+This makes sense to Lee: scheduling cannot proceed without first collecting availability.
+So he clicks ‘Approve & Continue.’ The system then generates four concrete actions, one email per speaker (Figure 7(b)), which Lee reviews and approves.
+Now the AI drafts the actual emails (Figure 7(c)).
+Lee reads the first draft carefully.
+He notices the date format is inconsistent and wants to add a note mentioning that his advisor recommended the speaker.
+He makes these edits and clicks ‘Regenerate.’ The revised draft incorporates all his changes.
+For the next email to Prof.
+Lucy Tran, the system has already learned Lee’s preferred date format and applies it automatically, requiring no further corrections.
+Lee approves and sends all four emails.
+Even though these are simulated, Lee is reassured: the AI is learning his style, and he can see exactly what would be sent in the real world.
+3.1.3. Handling responses from speakers.
+In reality, Lee would have to wait days or weeks for responses.
+In the simulation, time is compressed: a full day in the real world takes only one to two minutes.
+So Lee quickly sees what happens next.
+By the end of Day 1, three of the four speakers have responded.
+Michael Chen and Sarah Malik reply with their availability.
 Prof.
-B and Student Y expressed interest, with scheduling pending.
-Both Prof.
-B and Student X inquired about Zoom options instead of in-person attendance.” After reviewing the summary, Lee confirms the day’s progress and clicks Continue to Next Day to proceed.
-This iterative, day-by-day process continues until all seminar slots are successfully assigned.
+Lucy Tran also replies promptly, but with an unexpected request: “Thanks for reaching out, Lee!
+I was wondering if there might be an option to present via Zoom instead of in person?” Prof.
+Adrian Smith, true to his persona, does not respond at all.
+The assignment progress tracker immediately updates to reflect who has replied and what slots they are available for.
+The communication log records every exchange.
+Lee can see at a glance where things stand and where the gaps are.
+3.1.4. Edge cases handling.
+Prof.
+Lucy Tran’s Zoom request creates a problem: none of the existing policies address remote presentations.
+The AI does not try to handle this on its own.
+Instead, the edge case detector flags it and asks Lee for guidance: “Prof.
+Lucy Tran requests a Zoom option.
+Should I confirm Zoom availability or insist on in-person participation?” This is exactly the kind of situation Lee was worried about: the AI encountering something unexpected and making a bad call.
+But instead of acting autonomously, the system pauses and asks.
+Lee decides that Zoom is acceptable and types: “Zoom is allowed.” The AI generates a response to Prof.
+Lucy Tran, which Lee reviews and sends.
+Importantly, Lee’s decision does not resolve this case alone.
+The system records it as a stop hook and updates the policy set to include a new rule about allowing Zoom when speakers request it.
+If a similar situation arises later, for example, Michael Chen also asks about Zoom, the system will already know how to handle it without bothering Lee again.
+At the end of Day 1, the system generates a daily summary of everything that happened: which emails were sent, who replied, and what decisions were made.
+Lee reviews it, finds it accurate, and clicks ‘Continue to Next Day.’
+3.1.5. Clarifying, waiting, and following up.
+The calendar advances.
+Prof.
+Adrian Smith has not yet responded, while Prof.
+Lucy Tran’s Zoom details remain unresolved.
+The AI generates a plan: clarify Zoom logistics with Prof.
+Tran and continue waiting for Prof.
+Smith (Figure 2(Middle)).
+Lee checks the rationale, confirms it references the appropriate policies, and approves.
+The system drafts a follow-up to Prof.
+Tran, who replies confirming November 7th via Zoom.
+The dashboard updates accordingly, leaving October 4th as the only open slot.
+On Day 3, Prof.
+Smith has still not responded.
+The coordination agent, guided by the follow-up policy specifying action after two workdays of silence, drafts a gentle nudge.
+Lee notices it matches his usual follow-up style: short and to the point, with no unnecessary pressure.
+This is something the system picked up from uploaded emails and his earlier edits.
+He approves without changes.
+Later on, Prof.
+Smith replies with availability on November 7th, creating a new conflict: three speakers now favor the same date while October 4th remains unfilled.
+Lee can see the conflict clearly in the assignment progress tracker.
+From this point onward, the system continues to generate plans grounded in policies, learn from Lee’s past choices, and execute actions adaptively, resolving the scheduling conflict and filling the remaining slot until the seminar assignment is complete.
+3.1.6. Usage conclusion.
+By the time the DoubleAgents usage concludes, Lee has not only produced a tentative speaker-to-slot assignment but has also built up a set of refined policies, customized email templates, and executable stop hooks that align with his actual preferences.
+He now knows how the AI handles follow-ups, edge cases, and conflicts, because he has seen it happen in a safe and simulated environment.
+When moving to real-world deployment, such as sending actual emails to speakers, these artifacts directly guide the agent’s actions, reducing uncertainty and enabling more reliable coordination.
+Over time, this process accumulates into a reusable set of alignment artifacts that can be applied and further developed in future coordination tasks.
+3.2. System Implementations
+We implemented DoubleAgents as an interactive simulation system built on the ReAct framework (Yao et al., 2023).
+The system architecture comprises four key components: (1) a Simulation Engine that drives an iterative planning-execution loop through a coordination agent and LLM-based simulated respondent agents, (2) a Policy Module that encodes user values as natural language rules, supports policy selection and refinement through stop hooks, and captures implicit user preferences from edit behavior, (3) an Edge Case Detector that identifies situations beyond existing policy coverage and escalates them for user clarification, and (4) a Dashboard Visualization built on Flask that provides real-time visual feedback on assignment progress, communication history, and scheduling timelines.
+Users interact with the system via a control panel that supports plan review, action editing, and natural-language queries.
+Full implementation details, including prompts and system architecture, are provided in Appendix B.
 4. Technical Evaluation
-DoubleAgents enables users to engage in iterative task planning and action execution through two core functionalities: (1) policy selection, which guides the generation of plans aligned with predefined coordination strategies; and (2) robust edge case detection, which identifies problematic or ambiguous situations and escalates them to the user for clarification and resolution.
-In this section, we describe the technical evaluation of the two components.
+DoubleAgents enables iterative task planning and action execution through two core mechanisms: (1) policy selection, which steers plan generation toward user-aligned coordination strategies; and (2) edge case detection, which identifies ambiguous or potentially problematic situations and escalates them for user intervention.
+Together, these components support a structured interaction loop in which users can guide, refine, and correct agent behavior over time.
+We present a technical evaluation of these two components.
 4.1. Policy Selection Experiments
 4.1.1. Experiment Settings
-Policy selection refers to the process of identifying relevant coordination policies that guide the generation of high-quality, context-appropriate plans.
-Specifically, given a particular problem state (e.g., unfilled time slots and unresponsive speakers), the policy selection module is expected to retrieve all applicable policies that can serve as guidelines for the next planning step.
+Policy selection refers to identifying coordination policies that guide the generation of high-quality, user-aligned plans.
+Given a particular problem state (e.g., unfilled time slots and unresponsive speakers), the policy selection module retrieves relevant policies that can inform the next planning step.
 For example, if no speakers have been contacted yet, a suitable policy would be: “If we are just beginning and have not yet contacted any speakers for availability, then ask each speaker for their availability across all slots.” In our implementation, we adopt a seed set of general-purpose coordination policies curated by domain experts.
-To inform policy selection, we use the output of the getProgressSummary() API, which provides a high-level snapshot of the current task state.
+To study how different representations of the problem state affect policy selection, we compare four input conditions that vary in the level of abstraction and context provided to the model:
+  • Speaker-only context, which includes information about each speaker’s persona.
+For example: “Speaker 1 (Prof.
+Adrian Smith): busy, often replies only after multiple follow-ups; Speaker 2 (Dr.
+Lucy Tran): usually responds within 1–2 days; Speaker 3 (Sarah Malik): highly responsive and flexible; Speaker 4 (Michael Chen): prefers early scheduling but has not replied recently.”.
+  • TimeSlot-only context, which includes the current status of all time slots and their constraints.
+For example: “Four slots: Sep 5 (unfilled), Oct 3 (filled), Nov 7 (filled), Dec 5 (unfilled).
+The Nov 7 slot has been confirmed with a speaker, while Oct 3 is tentatively held.
+Remaining open slots require confirmation.
+  • Combined TimeSlot and Speaker context, which includes both sources of information without additional abstraction.
+For example: “Speaker 1 (busy, slow to reply) and Speaker 2 (moderately responsive) were last contacted two days ago; Speaker 3 (responsive) has not yet been contacted; Speaker 4 prefers early slots but has not replied.
+Slot status: Oct 3 and Nov 7 are filled (Nov 7 confirmed), while Sep 5 and Dec 5 remain open.
+  • Summary-based context, which includes the information of the speakers, time slots, and the summary of the current task state to provide a structured, high-level snapshot of the system.
 An example summary reads: “One slot has been secured: speaker 4 confirmed for Nov 7 (2025-11-07) and received a confirmation email on 2025-08-07.
@@
 Last outreach dates: speaker 1 and speaker 2 were last emailed on 2025-08-08; speaker 3 was last emailed on 2025-08-07.
-We evaluated the effectiveness of our approach by comparing it against three baseline variants, each using a reduced context: (1) speaker-only context, which includes only information about speaker responsiveness; (2) slot-only context, which includes only slot-level availability (e.g., filled vs.
-unfilled); and (3) combined speaker and slot context, which merges the two but omits higher-level summaries.
-Prompt templates for all conditions are provided in Appendix D.
-To evaluate the performance of each approach, we constructed a test suite of 30 examples, all sourced from real user study sessions and curated to reflect a diverse set of scheduling scenarios.
-For each example, we treat the set of policies approved by users during the study as the ground truth for evaluation.
-All evaluations were conducted using the o4-mini model with a default temperature of 1.
-We report the average F1 score across three independent runs for each experimental setting.
+Prompt templates for all conditions are provided in Appendix E.
+We construct a test suite of 30 examples sourced from real user study sessions, covering a diverse set of scheduling scenarios.
+For each example, we treat the set of policies approved by users during the study as ground truth.
+All evaluations are conducted using the o4-mini model with a default temperature of 1, and we report the average F1 score across three independent runs for each condition.
 4.1.2. Results and findings
-As summarized in Table 1, our summary-based context approach achieves the highest F1 score for predicting relevant policies given a specific problem state.
-The improvement is statistically significant when compared to all three baseline conditions, with an average F1 score of 0.70, demonstrating the effectiveness of using structured progress summaries to guide policy selection.
-Notably, simply aggregating speaker- and slot-related context without summarization does not improve performance; the combined-context baseline yields only 0.27 average F1, indicating that raw context alone is insufficient for accurate policy retrieval.
-Based on these results, we adopt the summary-based approach as the default mechanism for selecting policies, which are then passed to the orchestrator agent to guide plan generation in a policy-aligned manner.
-Prompting Techniques F1 Score
-Mean SD
+As shown in Table 1, the summary-based context yields the highest F1 score for predicting relevant policies given a specific problem state, outperforming all other input conditions.
+The improvement is statistically significant compared to the three baselines, indicating that providing a structured, high-level representation of task progress leads to more accurate policy selection.
+In contrast, simply exposing the model to raw contextual information, either from speaker personas, slot availability, or their combination, results in substantially lower performance.
+Notably, the combined-context condition does not improve over simpler variants, suggesting that aggregating low-level context without abstraction is insufficient for effective policy retrieval.
+Based on these findings, we adopt the summary-based representation as the default input for policy selection in DoubleAgents.
+The selected policies are then passed to the coordination agent to guide plan generation in a policy-aligned manner.
+4.2. Edge Case Detection Experiments
+4.2.1. Experiment settings
+Edge case detection aims to determine whether a respondent’s message introduces a scenario that falls outside the scope of existing coordination policies.
+For instance, if a speaker requests to present via Zoom while no policy addresses virtual presentations, the system should flag this case for user clarification and further instruction.
+To study how different prompting strategies affect detection performance, we compare three conditions that vary in the level of guidance and policy awareness provided to the model: (1) zero-shot prompting, where the model is given only the task instruction and the incoming message without any examples; (2) few-shot prompting without policy context (few-shot w/o policy), where the model is provided with representative examples of edge and non-edge cases but does not have access to the current set of coordination policies; and (3) few-shot prompting with policy context (few-shot w/ policy), where the model is given both demonstration examples and the existing policies, allowing it to assess whether a new situation is already covered or requires escalation.
+Prompt templates for all conditions are provided in Appendix E.2.
+We construct a test suite of 100 examples, manually validated by the lead authors to ensure diversity and coverage across common coordination scenarios and edge cases.
+The construction methodology is detailed in Appendix E.2.
+All experiments are conducted using GPT-4o, with temperature set to 0.7 and a maximum token limit of 100.
+We report the average accuracy across three independent runs for each condition.
+4.2.2. Results and findings
+As shown in Table 1, the few-shot prompting with policy context achieves the highest accuracy among the three conditions, outperforming both zero-shot and few-shot w/o policy baselines with statistically significant improvements.
+These results suggest that access to explicit policy context enables the model to more reliably determine whether a given scenario is already covered or requires escalation.
+In comparison, zero-shot prompting struggles to consistently distinguish between routine variations and true edge cases, while few-shot prompting without policy context benefits from examples but lacks the ability to ground its decisions in the current policy set.
+This often leads to either over-flagging benign cases or missing subtle policy gaps.
+Based on these findings, DoubleAgents adopts the policy-aware few-shot approach as its default mechanism for edge case detection.
+By grounding decisions in both prior examples and explicit policy coverage, the system more reliably surfaces novel or ambiguous situations, enabling timely user intervention while maintaining a smooth coordination workflow.
+Prompting Techni. Mean SD
 Speaker-only .27 .054
-Slot-only .11 .038
-Slot + Speaker .27 .017
+TimeSlot-only .11 .038
+TimeSlot + Speaker .27 .017
 Summary-based .70*** .035
-Table 1.
-The technical evaluation results for Policy Section comparing four prompting techniques.
-The best scores are shown in bold.
-Statistically significant improvements (independent two-sample t-test, $p<0.001$) over all baselines are marked with ***.
-Prompting Techniques Accuracy
-Mean SD
+(a) Policy Selection (F1 Score).
+Prompting Techni. Mean SD
 Zero-shot .79 .021
 Few-shot w/o Policy .88† .000
-Few-shot w Policy .93** .015
-Table 2.
-The technical evaluation results for Edge Case Detection comparing different prompting techniques.
-The best scores are shown in bold.
-Statistically significant improvements (independent two-sample t-test, p ¡ 0.01) over both baselines are marked with **, improvements over one baseline are marked with †.
-4.2. Edge Case Detection Experiments
-4.2.1. Experiment Settings
-Edge case detection involves identifying whether a respondent’s message presents a scenario that falls outside the coverage of existing coordination policies.
-For example, a speaker requesting to present via Zoom, when no policy has addressed virtual presentations, should be flagged to the user for clarification and further instruction.
-To enable reliable detection of such edge cases, we adopt a few-shot prompting strategy with access to existing policies (denoted as ”few-shot w/ policy”).
-We compare this approach against two baselines: (1) zero-shot prompting without any demonstration examples, and (2) few-shot prompting without access to policy references (denoted as ”few-shot w/o policy”).
-Concrete prompt templates for each setup are provided in Appendix D.2.
-To evaluate performance, we constructed a test suite of 100 examples, manually validated by the lead authors to ensure coverage and quality.
-The construction methodology for this test suite is detailed in Appendix D.2.
-All experiments were conducted using GPT-4o, with temperature set to 0.7 and a maximum token limit of 100.
-We report the average accuracy across three independent runs for each condition.
-4.2.2. Results and Findings
-As shown in Table 2, our few-shot prompting with policy method significantly outperforms both baselines, with statistically significant improvements across evaluation runs.
-These results demonstrate the effectiveness of incorporating policy context into the detection process, enabling the model to first analyze existing policies before determining whether a given case falls outside their scope.
-Together, DoubleAgents adopts the policy-aware few-shot approach as its default mechanism for edge case detection.
-This ensures that novel or ambiguous cases are reliably flagged to users, enabling timely clarification and maintaining a high-quality user experience.
+Few-shot w/ Policy .93** .015
+(b) Edge Case Detection (Accuracy)
+Table 1. Technical evaluation results. Best scores in bold. Significance via independent two-sample t-test: † $p<0.05$ over one baseline, ∗∗ $p<0.01$, ∗∗∗ $p<0.001$ over all baselines.
 5. User Study
-5.1. Goal and Overview
-To investigate how users engaged with DoubleAgents’ features and how interactive simulation shaped their use of agentic behavior, we recruited 10 participants to use DoubleAgents across two separate sessions.
-Our goal was to assess users’ perceived system usefulness, their understanding and reliance on the system, and how their understanding and calibration of agent behavior evolved over time.
-During each session, participants worked in a simulated environment to schedule a seminar with four dates and four speakers, while the system gradually adapted to participants’ values and preferences via their policies, email templates, and stop hooks.
+5.1. Study Design
+5.1.1. Overview
+To investigate how users and AI can reach alignment through repeated interaction and refinement — and how this alignment enables comfortable offloading — we recruited 10 participants across two sessions over two consecutive days.
+Each participant completed approximately 2.5 hours total: a 30-minute preparation and setup phase, followed by a 60-minute usage and evaluation session each day.
+Our goal was to assess whether comfort with delegation evolves over time, how the system’s three distributed cognition components support users in learning what is safe to delegate, and where users still need to maintain control.
+During each session, participants worked in a simulated environment to schedule a seminar with four dates and four speakers, while the system gradually adapted to participants’ values and preferences through their policies, email templates, and stop hooks.
 Participants completed two questionnaires.
-First, an AI comfortableness survey measured trust and alignment needs for proactive actions and comfort with automating planning and execution, administered before Session 1, after Session 1, and after Session 2.
-Second, a System perception survey captured UX and perceptions of 15 system features after Session 1 and after Session 2, enabling us to track how repeated use shifted alignment strategies, delegation boundaries, and reliance.
-Detailed study procedures and data analysis methods are provided in Appendix A.
-5.2. Participants
-We recruited 10 participants (7 female, 3 male; aged 21–30, M = 25.7) from our university community.
-All participants (10 of 10) reported using generative AI automation tools daily.
-Participants primarily used tools such as AI writing assistants (10 of 10), AI coding assistants (7 of 10), and AI design/creative tools (5 of 10).
-All participants reported solid understanding of generative AI: 7 described having in-depth knowledge of models and tools, while 3 reported a good general understanding of key concepts.
-Participants’ experience with human coordination tasks varied widely.
-2 participants had moderate to extensive experience, having led team schedules or group planning multiple times.
-3 participants had some experience, having organized small tasks or events independently.
-4 participants had minimal experience, primarily helping or shadowing others in coordination tasks.
-Finally, 1 participant had no prior experience, having never hosted or coordinated any tasks.
-5.3. Result
-5.3.1. Consistently Positive User Experience
-Across both sessions, all 10 participants reported a consistently positive experience with DoubleAgents.
-Overall system ratings remained high and showed a slight improvement (from 6.30/7 to 6.62/7).
-Participants also perceived the system as effective in generating optimal plans, actions, and emails (5.90/7 to 5.88/7), and judged it as robust (5.60/7 to 5.88/7).
-Both mental and physical demand remained low across sessions, decreasing from 2.90/7 to 2.13/7 and from 1.50/7 to 1.13/7, respectively.
-Participants reported high perceived performance (6.30/7 to 6.75/7), low frustration (1.80/7 to 1.25/7), and a near-significant reduction in effort (3.20/7 to 1.88/7, $p=0.054^{{}^{\wedge}}$).
-Together, these results indicate a consistently positive and efficient user experience across sessions, with effort showing the strongest improvement.
-5.3.2. Increasing User Reliance Driven by DoubleAgents’ Design
-Despite largely stable system-level scores, participants’ perceived reliance on DoubleAgents increased significantly over sessions (from 4.90/7 to 5.75/7, $p=0.02^{*}$).
-Participants characterized this shift not as “unconditional reliance”, but as increased comfort with delegation as the system’s behavior became more user-aligned, legible, and shapeable—enabled by clearer controls, predictable execution patterns, and explicit points of intervention.
-A correlation analysis between reliance and feature-level perceptions (Spearman; FDR-corrected) showed that reliance was associated with features supporting control and explainability, including state visualization ($r=0.78$, $p<0.05^{*}$) and confirmation-based user control ($r=0.51$, $p<0.05^{*}$).
-For example, P6 found the Communication Visualization section of the interface useful in showing the full email exchange histories between the organizer and each speaker.
-She mentioned, “Everything is consistent inside the system—when some speakers were not followed up, I could map it back to the visualization and confirm accuracy… [It] helped me build trust… For tomorrow, I’ll spend less time checking the email replies [inside the chat interface] one by one, and instead [just check] the Communication Visualization on the side.” Increases in reliance over time were strongly linked to DoubleAgents’ human-AI alignment features.
-In particular, reliance increased with perceived system robustness demonstrated through interactive simulation ($r=0.54$, $p<0.05^{*}$), and was reinforced by repeated successful issue flagging via stop hooks ($r=0.65$, $p<0.01^{**}$) and perceived value alignment through coordination policies and email writing ($r=0.54$, $p<0.05^{*}$).
-Reliance was negatively correlated with mental demand (r=0.47r=-0.47, $p<0.05^{*}$) and showed no significant correlation with other system experience or perception items.
-Collectively, these findings suggest that reliance increased not merely due to perceived system capability or general UX, but because of DoubleAgents’ support for alignment in design.
-The following section details how participants experienced and leveraged this alignment through interactive simulation.
+An AI perception survey measured alignment expectations, trust, and comfort with automating planning and execution; it was administered before Session 1, after Session 1, and after Session 2.
+A DoubleAgents experience survey captured UX, task load, and perceptions of system features after each session.
+Together, these instruments allowed us to track how repeated use shifted alignment, delegation boundaries, and the evolving balance between offloading and control.
+Detailed study procedures and data analysis methods are provided in Appendix C.
 Figure 3.
-Average user perception scores toward comfort with automating the (a) planning and (b) execution stage in seminar organizing, across three stages (average scores across 10 users: before Session 1, after Session 1, and after Session 2).
-denotes statistical significance at $p<0.1$, * at $p<0.05$, ** at $p<0.01$, *** at $p<0.005$.
-5.3.3. Increasing User Comfort with AI Automating Human Coordination Tasks: Contributing Alignment Supports.
-Over interactive simulations, participants became more comfortable delegating both planning and execution tasks to agentic AI.
-For planning (Figures 7a, participants became more comfortable letting the AI visualize current status (4.80/7 to 6.50/7, $p<0.05^{*}$), propose time assignments (5.10/7 to 6.38/7, $p<0.1^{\vee}$), decide daily next steps (4.70/7 to 6.50/7, $p<0.005^{***}$), and monitor/flag issues (5.50/7 to 6.50/7, $p<0.01^{**}$).
-For execution, comfort increased for drafting-only emails across all four types (from around 6.0 before to nearly 7.0/7 after, $p<0.01^{\vee}$), and also increased for draft-and-send tasks on users’ behalf—e.g., initial outreach rose from around 2.7/7 to 4.6/7 ($p<0.05^{*}$), with similar gains for follow-ups and final confirmation ($p<0.1^{\vee}$).
-Participants attributed this increased comfort to DoubleAgents’ alignment supports: interactive simulation; user-editable coordination policies that guide next-step decisions; email templates and style controls that realize those actions; and stop hooks for issue escalation.
-Together, these supports advanced bidirectional alignment: the AI agent’s actions became better aligned with users’ needs and values, and users developed greater understanding and trust, becoming more comfortable with the AI taking initiative in automating this human coordination task.
-Interactive simulation as a testbed for policy, email, and stop-hook calibration.
-Simulations supported alignment by letting participants safely probe edge cases, identify where the agent’s behavior diverged from expectations, and iteratively refine (1) what the agent should do (coordination policies), (2) how it should carry out those actions in email (templates and style), and (3) when it should pause for confirmation (stop hooks).
-Participants emphasized that interactive simulations increased confidence in their perception of the system’s robustness (5.80/7 to 6.38/7).
-They found simulated test cases realistic and representative of real-life scenarios (5.90/7 to 6.75/7), and especially useful for surfacing potential problems (5.80/7 to 6.50/7).
-P3 described how simulation exposed realistic exceptions and prompted advance policy-setting: “The simulations [over the 2 days] felt realistic and true to the person [they] were simulating, like their personality and emails… The simulation elicited possible problems that would happen in real life… Like [having] people asking for funding or [needing] to reschedule—it forced me to think about how I would handle that in real life and to create relevant policies, and that was good.
-I can trust the system to handle that in the future for me if scenarios like that occur, since it already happens in the simulation.
-Email templates and coordination policies fostered increased comfort with AI automation.
-Participants’ comfort with AI automating planning and execution increased over time as the system’s actions became more consistent, better grounded in the users’ own values and rationale, and increasingly “in their voice.” Carrying over value-aligned coordination policies made participants more comfortable delegating planning—e.g., having the AI decide daily next steps, track and visualize progress, make assignment decisions, and monitor/escalate issues when cases did not fit existing policies (Figure 7a).
-Email-template carryover particularly supported execution delegation.
-Comfort increased for drafting-only emails and for draft-and-send tasks (Figure 7b).
-One exception was drafting and sending clarification emails without any user confirmation, which showed little change (3.00/7 to 3.90/7 to 3.90/7).
-P10 noted that clarifications often require nuanced, user-specific judgment: “For clarification emails, I might want to double-check… [they’re] usually not in the plan [or policies]… I’m not comfortable [offloading that] to AI…” As a result, participants preferred to remain in the loop for clarification messages.
-Overall, participants described an early sense of being understood and represented by the system, as if it “got them,” when iterating on email templates and coordination policies.
-For example, P4 noted that edits to greetings, closings, and phrasing carried over into later drafts: “Compared to the last [usage], DoubleAgents get my style… so I edited fewer times this time.” Similarly, after refining policies, P3 observed that the agent’s choices matched their expectations: “I found that the agent’s recent actions followed exactly the policies I had in mind—super aligned.” This growing alignment helped the system reflect their intended actions and communication style, increasing their comfort with offloading tasks to AI agents.
-Stop hooks helped establish safe delegation boundaries, fostering users’ comfort with delegation.
-Participants highlighted stop-hook flagging as particularly important for maintaining safe delegation boundaries.
-For monitoring and surfacing potential issues, P1 described issue flagging as “one of the most important features of the system” and shared, “[With the issue flagging,] I skim [through the other features] now… It made me trust the system because I know it won’t just execute something random… It flags when it doesn’t know… That helps elicit my values and how I want to handle things—and [then DoubleAgents] respects that…” P8 described a case where flagging prevented a misinterpretation of progress—when they skimmed the respondent agent’s reply and thought everything was resolved, the flag showed that it was not.
-Thus, with these “stop hooks” continuously monitoring for exceptions and prompting intervention when needed, participants felt much more comfortable offloading tasks to the AI.
-Figure 8.
-Average user perceptions of alignment needs, expectations, and trust in agentic AI systems (average scores across 10 users: before Session 1, after Session 1, and after Session 2).
-* denotes statistical significance at $p<0.05$ and ** at $p<0.01$.
-Increasing Expectations and Needs for Alignment in Agentic AI
-Across sessions, participants reported stronger expectations for how proactive AI in general should behave and align with them (Figure 8).
-In the AI Comfortableness Survey, participants reported a significant increase in wanting the AI’s actions to align with their preferences, values, and priorities, from 6.10/7 (before using the system, based on their previous AI experiences) to 7.00/7 after two uses of DoubleAgents ($p<0.05^{*}$).
-Participants’ desire to stay in control remained high (5.30 to 5.75).
-They also reported an increased need for detailed system rationales and explanations before the system acts, from 5.00/7 to 6.00/7 ($p<0.05^{*}$).
-Participants framed clear and abundant explanations as a core alignment mechanism for building an accurate mental model of the AI’s decision process, enabling verification, and determining when and how to intervene.
-P6 mentioned, “[With the system explanation] I felt like I could tell exactly what it’s doing at every stage so there’s a lot of transparency… if I didn’t get that, I would feel lost [and] not know what’s going on…” We interpret this as evidence that DoubleAgents’ interaction design helped participants recognize the importance of alignment and articulate stronger alignment expectations for future proactive AI usage.
-Furthermore, participants also reported statistically significant increases in trust to let similar AI systems act on their behalf, from 3.20/7 to 5.00/7 ($p<0.01^{**}$), even as their assessments of AI reliability remained relatively the same.
-This suggests that alignment features can increase comfort with delegation beyond improvements in perceived raw capability.
-P1 noted that even though DoubleAgents represented only one specific experience, “I felt more optimistic… [and] in seeing similar systems like DoubleAgents in the future, I would generally trust such AI systems more.”.
+Average user-reported scores across 10 participants for comfort with automating (a) planning and (b) execution tasks, across sessions over time (1: strongly disagree, 7: strongly agree).
+$p<0.1$,  $p<0.05$, ∗∗ $p<0.01$, ∗∗∗ $p<0.005$.
+5.1.2. Participants
+We recruited 10 participants (7 female, 3 male; aged 21–30, $M=25.7$) from our university community via personal contacts and departmental mailing lists.
+All reported daily use of generative AI tools, including AI writing assistants (10/10), coding assistants (7/10), and design tools (5/10).
+Seven described in-depth knowledge of generative AI models, while three reported a good general understanding of key concepts.
+Participants’ experience with human coordination tasks varied widely: 2 had moderate to extensive experience, having led team schedules or group planning multiple times; 3 had some experience, having organized small tasks or events independently; 4 had minimal experience, primarily helping or shadowing others in coordination tasks; and 1 had no prior experience.
+5.2. Results
+We organize our findings around the central question of how humans and AI reach alignment through repeated interaction and refinement, enabling comfortable offloading.
+We examine three aspects: whether comfort with delegation evolves over time (§5.2.1), how the three distributed cognition components support participants in learning what is safe to delegate (§5.2.2), and where participants still need to maintain control (§5.2.3).
+5.2.1. Did participants become more comfortable delegating control over time?
+Participants reported increased comfort delegating both planning and execution tasks to AI
+For planning (Figure 3a), comfort ratings rose across all four sub-tasks (letting the AI track and visualize task state, solve scheduling constraints such as time-slot assignments, reason about context to determine next steps, and monitor for edge cases requiring user attention), with averages climbing from roughly 5/7 before deployment to around 6.5/7 afterward.
+Constraint solving for slot allocation is a comparatively task-specific capability and showed only marginally significant improvement.
+In comparison, the planning sub-tasks, tracking, monitoring, and flagging shows strong comfort gains over time, consistent with the distributed-cognition affordances of the coordination agent that help organizers offload ongoing awareness work.
+For execution (Figure 3b), the pattern split by autonomy level: participants were already fairly comfortable letting the system draft emails on their behalf, and comfort rose further (from around 5.7/7 to nearly 6.8/7 across all four email types, each at least marginally significant).
+By contrast, comfort with the system both drafting and sending emails—where the social stakes of an error are higher—started considerably lower and remained so, though it still improved meaningfully (from roughly 3/7 to 4.3/7 across email types).
+Among the four email types, initial outreach showed strong gains, likely because its content is relatively formulaic and low-variance; follow-up and confirmation emails improved modestly, while clarification emails showed no significant change over the alignment process—an exception we return to in Section 5.3.2.
+Participants’ perceived reliance on DoubleAgents, as measured by the DoubleAgents Experience survey, increased significantly (4.90/7 to 5.75/7, $p=0.02^{*}$).
+Participants characterized this not as unconditional reliance, but as increased comfort with delegation as the system’s behavior became more predictable and legible — feeling both aligned with how they wanted to work and comfortable letting the system act on their behalf.
+P2 articulated this arc explicitly: “I would intervene more at first but if it can demonstrate that it’s consistently able to do it correctly, I’m going to let it act on its own on my behalf more autonomously.”
+5.2.2. How did participants use and rate the three distributed cognition components?
+Policies and templates helped participants encode their values, reducing effort over time.
+Carrying over user-defined coordination policies strengthened alignment in planning: participants found the AI’s decisions on next steps, progress tracking, and escalation increasingly reflective of their intent.
+Email-template carryover similarly improved alignment in execution, with participants reporting that drafts better matched their expectations over successive sessions.
+Quantitative measures supported this pattern: perceived value alignment via policies and templates increased from 5.50/7 to 6.12/7, drafting comfort rose from approximately 6.0 to nearly 7.0/7 across all email types, and effort dropped from 3.20/7 to 1.88/7 ($p=0.054^{\wedge}$).
+P4 noted that edits to greetings, closings, and phrasing carried over into later drafts: “Compared to the last [usage], DoubleAgents get my style… so I edited fewer times this time.” P3 observed: “I found that the agent’s recent actions followed exactly the policies I had in mind — super aligned.
+Visualizations made the agent’s reasoning legible, enabling confident evaluation.
+The dashboard’s temporal, social, and procedural views let participants verify what the agent was doing without parsing raw data.
+P6 found the communication visualization useful for showing full email exchange histories: “Everything is consistent inside the system — when some speakers were not followed up, I could map it back to the visualization and confirm accuracy… [It] helped me build trust… For tomorrow, I’ll spend less time checking the email replies [inside the chat interface] one by one, and instead [just check] the Communication History on the side.
+The coordination agent served as memory and reasoning partner, removing tracking burden.
+By maintaining full state across days and weeks of coordination—who was contacted, who responded, what is pending—the agent freed participants from the cognitive overhead of tracking progress themselves.
+This let them focus on judgment and decision-making rather than bookkeeping.
+P7 noted that DoubleAgents was “pretty good at figuring out plans, keeping track of who’s responded, who’s not, who’s available, who’s not, like what to do moving forward.” P1 similarly shared that “now …I didn’t have to use my brain too much [to make sure things are consistent], just a little bit—so it handled the task really well.
+5.2.3. Where did participants still need control rather than full delegation?
+While comfort with delegation grew, participants identified boundaries where human judgment remained essential.
+Stop hooks for flagging edge cases.
+Participants highlighted stop-hook flagging as critical for maintaining safe delegation boundaries.
+P1 described issue flagging as “one of the most important features of the system”: “[With the issue flagging,] I skim [through the other features] now… It made me trust the system because I know it won’t just execute something random… It flags when it doesn’t know… That helps elicit my values and how I want to handle things — and [then DoubleAgents] respects that…” P8 described a case where flagging prevented a misinterpretation of progress: when they skimmed a reply and thought everything was resolved, the flag showed that it was not.
+Stop hooks gave participants confidence to offload routine execution precisely because they knew the system would pause at points of uncertainty.
+Stop hooks gave participants confidence to offload routine execution precisely because they knew the system would pause at points of uncertainty.
+Comfort in offloading, in other words, was built not by the agent performing flawlessly, but by making its boundaries of confident action transparent and actionable.
+Clarification emails as a natural boundary for delegation.
+The notable exception to increasing delegation comfort was clarification emails sent without user confirmation, which showed little change across sessions ($p>0.1$).
+P10 explained: “For clarification emails, I might want to double-check… [they’re] usually not in the plan [or policies]… I’m not comfortable [offloading that] to AI…” Unlike routine emails that follow established templates, clarification emails require the agent to identify ambiguity, judge what information is missing, and formulate the right question — all of which demand user-specific judgment that resists encoding into policies.
+This suggests a natural boundary for delegation: as agents take on more execution, the critical human role shifts toward decision-making at points of uncertainty, precisely the moments where knowing “when and how to ask the user” becomes the key alignment challenge.
+5.2.4. Simulation as an accelerator for alignment
+The three alignment mechanisms described above—policies, visualizations, and the coordination agent—improve through repeated interaction.
+But in live deployment, edge cases surface slowly: a speaker may take days to reply, and unusual requests arise only once across an entire series.
+Simulation accelerates this by compressing weeks of coordination into minutes while maintaining realistic diversity of cases.
+Participants found simulations realistic (5.90/7 to 6.75/7), useful for surfacing potential problems (5.80/7 to 6.50/7), and confidence in DoubleAgents robustness increased through simulation (5.80/7 to 6.38/7).
+P3 described: “The simulations felt realistic and true to the person [they] were simulating… it elicited possible problems that would happen in real life… Like people asking for funding or needing to reschedule—it forced me to think about how I would handle that and to create relevant policies.
+I can trust the system to handle that in the future since it already happens in the simulation.
+Simulation served all three mechanisms as a testbed: it provided a safe environment to probe whether agent behavior matched user intent (policies), populated the dashboard with realistic state for practicing oversight (visualizations), and surfaced ambiguous situations where the agent should pause (stop hooks)—letting participants define delegation boundaries in advance rather than discovering them through live mistakes.
+This parallels unit testing in software engineering, but where correctness is user-specific rather than universally defined.
+Simulation is not required—alignment can develop through live use—but it compresses the iteration cycle and lets participants build confidence before real social stakes are involved.
 6. Deployment Studies
-6.1. Goal
-As the in-lab study demonstrated, the simulation was helpful, and participants felt they were getting more comfortable use agentic AI help planning and executing human coordination tasks.
-However, would real-world seminar organizers find DoubleAgents useful and relevant?
-If so, how would they use it-would they plan, draft, and send invitation, follow-up, or clarification emails in different ways?
-Would alignment be established in a similar manner?
-What additional challenges might arise in real-world use?
-To explore these questions and evaluate the real-world deployability of DoubleAgents, we invited organizers who were managing actual events across three real-world deployments.
-6.2. Deployment Participants and Procedures
+6.1. Study Design
+6.1.1. Goal
+The lab study demonstrated that interactive simulation helped participants become more comfortable delegating socially embedded tasks to agentic AI.
+The deployment study examines whether this alignment enables organizers to confidently offload real coordination tasks, and how alignment evolves when the agent operates in live, socially complex environments.
 6.1.2. Use cases and participants
-In our deployment study, DoubleAgents was used in three real-world event-planning workflows involving two organizers (P11, P12).
-P11 used DoubleAgents to coordinate a Fall 2025 HCI seminar series (Aug 2025; Deployment 1, D1), inviting 10 speakers from a pool of 23 to fill 10 slots; we observed and logged system use for 7 days during this process.
-P11 later used DoubleAgents again for a Spring 2026 HCI seminar series (Jan 2026; Deployment 3, D3), inviting 11 speakers from a pool of 20 to fill 11 slots; we tracked use for 11 days.
-P12 used DoubleAgents to organize a conference workshop panel (Sep 2025; Deployment 2, D2), inviting 5–6 panelists from a pool of 18 to fill a single slot; we tracked use for 7 days.
-We recruited P11 and P12, two experienced seminar organizers (one male, one female; average age 31), via personal connections.
-Both were PhD students who had previously organized departmental research seminars and had prior experience using generative AI for writing emails.
-6.2.2. Introduction, Setup, and Interactive Simulation Rounds
-Organizers were first informed about the study and provided consent.
-They were then introduced to the study procedures and the DoubleAgents system through a demonstration, which included planning–execution loops, illustrated with a simulation example from the user study.
-Organizers also completed the initial survey covering AI perception and demographics.
-To set up the system, we disabled the respondent module from the lab version of DoubleAgents and connected the SendEmail() and checkEmailUpdates() tools to Gmail’s SMTP service, enabling organizers to send emails and receive replies through their personal Gmail accounts.
-For the simulation rounds, organizers prepared a simulation spreadsheet using their real organizing data (including real potential speakers’ names), first with placeholder email addresses and later with the real ones, and provided their Gmail SMTP credentials.
-Clear instructions and consent procedures ensured again that participants understood data usage, privacy rights, and the option to withdraw.
-Before connecting their real email accounts, participants completed a 2-hour interactive simulation session using pre-generated speaker personas based on the real candidates, allowing them to safely familiarize and themselves align with the system’s workflows.
-Finally, they tested the real email-sending module twice by sending messages to placeholder addresses (the researchers’ and organizers’ personal emails) to confirm successful delivery; afterward, we switched the recipient addresses to the invitees’ real Gmail accounts.
-6.2.3. Live Email Execution and Response Handling
-After the simulations, organizers used DoubleAgents to initiate the coordination process and track it for over a few days.
-They continued receiving system updates over the following week (D1: 7 days, D2: 4 days, D3: 11 days).
-We communicated with participants consistently, logging interactions, usage patterns, and feedback on the system’s support for scheduling, tracking, and learning.
-Organizers’ system interactions, including email sends, scheduling updates, and system adaptations, were tracked.
-Qualitative feedback was collected via check-in meetings, while quantitative system logs were collected for system learning over the week.
+DoubleAgents was used in three real-world event-planning deployments (D1–D3) involving two organizers (Q1, Q2), both PhD students (one male, one female; average age 31) with prior experience organizing departmental seminars and using generative AI for email writing.
+Q1 coordinated two HCI seminar series (D1, D3) and Q2 organized a conference workshop panel (D2).
+Table 2 summarizes each deployment.
+Deployment ID D1 D2 D3
+Organizer ID Q1 Q2 Q1
+Task Fall seminar Workshop panel Spring seminar
+Date Aug 2025 Sep 2025 Jan–Feb 2026
+Candidate pool 23 18 20
+Speakers invited 10 6 11
+Deployment length 7 days 4 days 12 days
+Table 2. Summary of three real-world deployments.
+6.1.3. Deployment procedure
+Organizers were first informed about the study and provided consent, then introduced to the system through a demonstration of planning–execution loops using a simulation example from the lab study.
+Before going live, each organizer completed at least two interactive simulation sessions (2+ hours) using personas based on their real candidates, iterating on policies, templates, and contextual inputs until the agent’s proposed actions matched the organizer’s intent.
+For deployment, we disabled the respondent module and connected the system to Gmail’s SMTP service, enabling organizers to send and receive real emails through their personal accounts.
+Once organizers judged their alignment sufficient, they used DoubleAgents to initiate and manage their coordination tasks over the live durations listed in Table 2.
+During execution, the agent operated on live emails while organizers retained approval authority over each action.
+We tracked system interactions including email sends, scheduling updates, and policy adaptations, and collected qualitative feedback via regular check-in meetings.
 6.2. Deployment Results
-6.3.1. DoubleAgents’ Perceived Usefulness and Relevance in Real-Life Human Coordination Problems.
-After the interactive simulations, both organizers (P11 and P12) reported that DoubleAgents was highly relevant to their existing seminar-organizing workflows.
-They highlighted that the system fit seamlessly into their email-heavy tasks, and that its control settings and edge-case flagging helped them handle complex coordination scenarios that conventional scheduling tools (e.g., Calendly) cannot easily manage.
-P11 emphasized that crafting good email templates is a delicate and important task because it requires personalizing messages to “make the speakers feel they are special,” even when requesting a favor.
-Both P11 and P12 found DoubleAgents capable of generating strong starter emails for them to edit early on, as well as producing acceptable final emails to send.
-P12 especially appreciated having explicit coordination policies in the system to guide next-step decisions in ways that matched his decision-making habits and values.
-He also found it useful that rationales were always visible and editable, in contrast to AI tools that hide their reasoning.
-After using DoubleAgents, P12 remarked, “I trust AI [systems] for specific use cases more,” explaining that he found DoubleAgents dedicated to and capable of generating high-quality emails.
-P11 found the color-coded visualization of speaker constraints and response states (e.g., not yet outreached, tentative yes, yes, no) especially useful; after D1, she adopted a similar color-coded scheme for other tracking tasks she previously managed in spreadsheets.
-6.3.2. Interactive Simulation as an Alignment Method for Complex, Realistic Tasks.
-Both P11 and P12 emphasized that interactive simulation was important for understanding DoubleAgents and for setting up initial coordination policies and email templates.
-They appreciated that simulations used real invitee names and incorporated persona cues and edge cases (e.g., funding requests, cancellations, Zoom requests), which helped them anticipate challenges and craft templates that would transfer to live deployment.
-For example, P12 used early simulation rounds to iterate on what information to include and where to place key details (e.g., the workshop webpage link) in his email templates, similar to how he had previously tested invitation templates by emailing himself.
-P11’s D1 simulation surfaced multiple policy- and tone-sensitive situations (e.g., rescheduling after a slot had already been assigned), which prompted iterative template and policy revisions.
-P11 reported that the system’s drafts were largely aligned with her language, but she removed or adjusted certain lines to better match her preferences and to clarify what information should be included when edge cases occur.
-Overall, P11’s D1 interactive simulation took four days across five meetings: the first two meetings focused on ensuring the system ran smoothly for a high-complexity scenario (20+ candidates, 10 slots, and substantial context per speaker).
-During setup, P11 revised invitation and follow-up templates more than four times and updated speaker descriptions so email generation could incorporate richer context.
-After we refined the system and onboarding protocols, P12’s D2 simulation took two days across two meetings.
-P12’s panel invitation task was also simpler (e.g., shorter descriptions and a single-slot panel), which further reduced the needed time.
-Both organizers reported that practice simulation rounds were valuable, and that sending test emails via SMTP to his own personal gmail helped him gradually understand system capabilities and limitations.
-Overall, these deployments suggest that simulation and onboarding effort scales with task complexity and the amount of required context.
-During D1, P11 specifically highlighted the importance and difficulty of providing sufficient context for highly personalized emails, especially for real-world cases.
-She noted that “[persona description management] is really tricky.” For instance, one invitee had previously coordinated with a different organizer in a prior semester (Friday availability), but the D1 seminar series had shifted to Thursdays.
-To avoid missing or conflicting context across past threads, schedule changes, and templates, P11 updated the persona description used for simulation with an additional line noting that the invitee had been told the wrong day in the past (Friday 4–5pm) and that the seminar was now on Thursday 4–5pm.
-In another case, P11 had already discussed the seminar with a speaker on Discord but initially omitted this context during early simulations; she then added it to the persona description so the follow-up email could adopt the appropriate tone and content.
-After rehearsing with real-speaker data in simulation, P11 finalized a preferred template for live sending.
-P11 also noted that even in D3 (her second deployment), simulation remained necessary because “real life is super complex.” She observed new patterns in edge cases, such as delayed responses or vague “check back next month” replies from senior faculty, which led her to revise policies (e.g., longer follow-up intervals) and adopt a more formal email tone.
-Across both deployments, simulation served not just as a testing ground, but as a means of progressive alignment: helping users adapt policies, templates, and expectations to shifting social and contextual dynamics.
-These findings suggest that onboarding and setup effort scale with task complexity and that simulation continues to play an ongoing role in supporting user-aligned adaptation, even after initial deployment.
-6.3.3. Example: Live Email Execution, Response Handling, and Edge Cases
-During D3, P11 sent emails to 11 selected speakers and relied more heavily on templates and system-generated drafts on Day 1 (a Friday), reflecting increased familiarity with the workflow.
-DoubleAgents then suggested waiting for responses over the following two days due to the weekend, which P11 accepted.
-Late on Day 3 (Sun), Invitee J and Invitee A replied with questions, which DoubleAgents surfaced in P11’s Day 4 (Mon) DoubleAgents session.
-On the morning of Day 4, Invitee B replied with availability.
-Based on these responses, DoubleAgents suggested replying first to Invitees A and J, but not to Invitee B, to preserve future scheduling flexibility and avoid potential conflicts.
-P11 accepted this suggestion and drafted replies accordingly.
-Specifically, Invitee A asked a funding-related question and included a personal greeting (e.g., “Happy New Year… if it doesn’t materialize…”).
-P11 revised the AI-generated reply by incorporating similar phrasing from their own writing style, but chose not to save this edit as a reusable template, noting that it reflected a personal, context-specific language choice.
-On Day 5, Invitees I and J replied, and DoubleAgents suggested allocating them to earlier-semester slots, which P11 accepted and responded to.
-Similar reply-and-assign cycles occurred on Day 6.
-On Day 7, DoubleAgents suggested following up with all the invitees who had not yet responded, in accordance with P11’s newly updated D3 policy to follow up every four working days (rather than every two, as in D1).
-P11 proceeded using a personalized follow-up template that incorporated questions added on Day 4.
-Also on Day 7, P11 handled additional questions and confirmed slots with Invitees C and H, who had replied promptly with apologies and availability.
-Later that evening, Invitee I emailed to report a schedule change and requested reassignment to an unclaimed date.
-On Day 8, DoubleAgents suggested reassigning Invitee I and requesting additional availability from Invitees H and B to resolve emerging conflicts.
-P11 accepted and carried out these actions.
-Over the weekend (Days 9 and 10), Invitee H responded with new dates but noted that P11 should follow up again the following month, as their schedule was still tentative.
-DoubleAgents accordingly scheduled a follow-up for Day 35, while on Day 11 it prompted P11 to address remaining unresponsive and unresolved cases.
-Collectively, these interactions illustrate the complexity and multi-threaded nature of this real-world task.
-As P11 reflected, “Even though I do this [seminar scheduling] all the time, [DoubleAgents] was really helpful and saved me a lot of time… Policy-abiding [actions and template-following] emails are really useful.
+6.2.1. Alignment as a precondition for offloading.
+Both organizers reported that DoubleAgents fit seamlessly into email-heavy coordination workflows that conventional tools (e.g., Calendly) cannot manage, and attributed their willingness to offload tasks to the system’s alignment mechanisms—explicit, editable policies and templates that made the agent’s reasoning legible.
+Q2 valued having editable policies to guide next-step decisions, contrasting this with AI tools that hide their reasoning; after deployment, Q2 remarked, “I trust AI [systems] for specific use cases more.” Q1 highlighted that messages must “make speakers feel they are special” and found DoubleAgents capable of generating strong starter drafts that improved over time.
+6.2.2. Building alignment through and beyond simulation.
+Both organizers used interactive simulation with real invitee names and persona-specific edge cases (e.g., funding requests, cancellations, Zoom preferences) to iteratively refine policies and templates before going live.
+Q2 used early rounds to iterate on template structure; Q1’s D1 surfaced tone-sensitive situations (e.g., rescheduling after a slot was already assigned), prompting revision of both templates and policies.
+Setup effort scaled with task complexity: Q1’s D3 (20 candidates, 11 slots, substantial per-speaker context) required four days across five meetings; Q2’s simpler panel invitation took two days across two meetings.
+A recurring challenge was context curation—Q1 noted that persona description management was “really tricky,” requiring updates such as clarifying schedule changes or adding context from prior Discord conversations so emails would adopt the right tone.
+Simulation-established alignment transferred effectively to structurally predictable scenarios—funding questions, scheduling conflicts, and reassignment requests were handled with only minor adjustments—but deployment surfaced unanticipated dynamics: noncommittal faculty responses (e.g., “check back in two months”) and cross-platform context required ongoing policy and tone adjustments.
+Even in her second deployment, Q1 found renewed simulation necessary because “real life is super complex.
+6.2.3. Distributed cognition across organizer and agent.
+Beyond alignment, the deployments revealed how DoubleAgents redistributed cognitive labor between organizer and agent through three complementary mechanisms:
+Visualization as shared group representation.
+Over deployment, Q1 found DoubleAgents’ color-coded visualization of speaker states (e.g., tentative yes, confirmed, declined) so useful that she adopted a similar color-coded scheme for other social tracking tasks she managed in spreadsheets beyond D1 and D3.
+Q2 drew on his existing practice of using ranked spreadsheets to monitor invitation progress, and described DoubleAgents’ visualization as supporting trust through “natural mapping” to that workflow.
+Policy and artifact refinement as externalized strategy.
+Over deployment, both organizers iteratively edited policies and templates to improve alignment.
+Q1’s D1 simulation surfaced policy-sensitive situations (e.g., rescheduling after a slot had already been assigned), prompting new policies.
+For D3—targeting more senior faculty than D1—she changed follow-up frequency from every two working days to every four, adopted more formal language, and revised templates more than four times to match the tone she identified after D1.
+Q2 valued explicit policies that matched his decision-making habits, appreciated that rationales were always visible and editable (unlike AI tools that hide reasoning), and drafted new policies to address gaps found during simulation.
+During deployment, Q1 observed that her edits carried forward to subsequent emails, so the artifacts accumulated organizational knowledge over time, requiring fewer changes.
+After one initial revision, Q2 showed a similar pattern: over usage, he began skipping full review of longer agent-drafted emails, trusting the style transfer and its consistency.
+Coordination agent as attention manager.
+In Q1’s D1 and D3 deployments, DoubleAgents managed parallel reply-and-assign cycles across 11 speakers over two weeks, at times handling three or four near-simultaneous replies with overlapping availability conflicts.
+Even in Q2’s smaller D2, he found it sufficient to check DoubleAgents for coordination status rather than scanning his inbox.
+Q1 shared that DoubleAgents saved her time at the end of D3: “Even though I do this [seminar scheduling] all the time, [DoubleAgents] was really helpful and saved me a lot of time… Policy-abiding [actions and template-following] emails are really useful.
 The new policies and templates I developed through this semester’s simulation—since everyone is a professor—worked really well.
-The simulation part cannot be skipped.” This case highlights DoubleAgents’ strong use case, demonstrating how simulation, policies, and email templates work together to support alignment between the system’s actions and users’ intentions.
 7. Discussion
-7.1. Human-AI Alignment via Interactive Simulations
-As AI systems evolve from passive tools to agentic collaborators that plan and act on users’ behalf, human–AI alignment becomes not just a technical challenge but an interactive one.
-Our findings suggest that interactive simulation offers a compelling design mechanism to support this alignment by foregrounding user agency, surfacing system behavior early, and fostering a co-adaptive relationship between humans and AI.
-Across our lab study and deployments, we observed that interactive simulation enabled users to engage with alignment as a gradual process, rather than a one-time configuration step.
-Participants entered the system unsure of their own preferences and reluctant to delegate; yet through repeated simulated cycles, each revealing system actions, uncertainties, and edge cases, they clarified values, authored and refined policies, and ultimately built enough user comfort to let the system act on their behalf.
-This shift mirrors real-world dynamics of human delegation: people do not fully trust a new assistant immediately but gain confidence through concrete interactions over time.
-Simulation thus serves multiple alignment functions simultaneously.
-First, it externalizes system behavior, making invisible decision logic (e.g., why the agent followed up or chose a certain speaker) legible to users.
-Second, it creates safe interaction spaces, akin to usability testing for AI plans.
-Like sandboxed environments in traditional software testing, simulations help users anticipate downstream consequences without real-world stakes, crucial in settings like email coordination where tone, timing, and etiquette are socially sensitive.
-Third, simulation supports user-side alignment, helping users identify preferences they hadn’t articulated upfront (e.g., when to prioritize seniority, how long to wait before following up, or what tone feels appropriate across contexts).
-Rather than requiring exhaustive foresight, DoubleAgents invites users to iterate on policies as new situations emerge.
-This design reframes alignment as a human-AI collaboration process: the agent proposes actions, but the human remains in the loop, interpreting, approving, and updating the future behavior of the system.
-In contrast to approaches that aim to extract user goals from demonstrations or feedback, we emphasize the value of direct, interactive co-construction of alignment artifacts, such as stop hooks, policy rules, and email templates, which evolve with use cases.
-This opens new design directions: how can we better scaffold early-stage alignment for users who don’t yet know their own preferences?
-What kinds of interfaces help users develop trust while retaining oversight?
-How might simulation be extended to other high-stakes or interpersonal agentic domains, such as hiring, negotiation, or caregiving?
-Our work suggests that alignment is not a prerequisite to agentic AI, but a continuous, situated interaction practice.
-By embedding simulation into the human–AI workflow, we give users the tools not just to observe and react to agent behavior, but to actively shape and steer it.
-In doing so, DoubleAgents contributes to a broader vision of agentic AI that is transparent, participatory, and grounded in human values from the start.
-7.2. Designing User Onboarding Experiences for Agentic AI
-Onboarding is a critical site for agentic AI systems.
-Unlike traditional tools, agentic systems require users to make early decisions about goals, decision boundaries, tone and formality, and what to delegate.
-They also create distinct discoverability challenges: users often do not know what the agents can do and what they can do.
-Key modes and safeguards (e.g., auto-send vs.
-draft-only) may be implicit, and failures can appear only in rare or edge cases—so there is no single, linear “progress bar” that reveals system status or readiness.
-As a result, effective onboarding is necessary for users to form accurate mental models, judge reliability, and feel comfortable delegating and deploying the agent in higher-stakes settings (e.g., a seminar organizer connecting to Gmail to send real invitations to senior faculty—who may also be potential postdoc advisors—where mistakes can harm professional relationships and reputation).
-In DoubleAgents, we found that interactive simulation provides a powerful onboarding scaffold, helping users observe, evaluate, and refine system behavior before real-world use.
-Rather than presenting configuration as a static setup phase, DoubleAgents embeds onboarding through repeated rounds of interactive simulation.
-During these rounds, edge cases (e.g., a speaker canceling or requesting Zoom) expose users to socially complex scenarios, help them become familiar with the system’s behavior, and elicit their personal preferences.
-Participants in our study often hesitated to delegate initially, not due to lack of interest, but due to uncertainty about agent behavior.
-Using repeated rounds of interactive simulation as an onboarding task lowered this barrier by helping participants move past the initial learning curve and novelty effects, and by revealing how the agent behaved across diverse cases.
-Like training a new assistant, onboarding becomes a co-configuration process in which expectations, autonomy, perceived usefulness, and trust are developed together.
-We argue that onboarding for agentic AI should not be treated as a brief tutorial animation or tooltip overlay.
-Instead, it can be designed as a structured, interactive process where users and systems collaboratively define goals, boundaries, and escalation rules.
-DoubleAgents shows how early simulations can accelerate alignment, ease delegation, and set the foundation for long-term, trustworthy collaboration.
+7.1. Beyond Organizing Seminars
+While DoubleAgents is instantiated for seminar organization, its underlying architecture is domain-general.
+Extending it to other coordination settings highlights new design opportunities across each component of the distributed cognition framework.
+Bespoke visualizations for different domains.
+Our dashboard provided temporal, social, and procedural views tailored to scheduling.
+Other domains will require different external representations: a hiring pipeline may benefit from a funnel view across candidate stages, while conference planning may require dependency graphs linking venues, catering, and speaker logistics.
+The insight from distributed cognition is that the right external representation enables evaluation at a glance (Hutchins and Klausen, 1996), which means the visualization layer should be treated as a bespoke design problem for each domain.
+One promising direction is to make this layer adaptive: users could specify what they need to monitor, and the system could generate or iteratively refine visual layouts.
+Such a malleable representation layer would allow the interface to co-evolve with users’ understanding of the task.
+From single-step policies to multi-step processes.
+Our current policy module encodes single-step conditionals (e.g., when X happens, do Y).
+However, many coordination tasks unfold as multi-step procedures.
+For example, inviting a senior speaker may involve an informal inquiry, followed by a formal invitation, and then logistical coordination, each governed by distinct norms.
+Supporting such process-level specifications would align the system more closely with how organizations encode knowledge: as structured workflows or playbooks rather than isolated rules (Hutchins and Klausen, 1996).
+This shift also introduces new interface challenges: users must be able to track not only which policy is active, but where each stakeholder sits within an evolving process.
+Integration with existing coordination ecosystems.
+Real-world coordination is deeply embedded in existing tools: communication occurs in Slack, tasks in Jira, schedules in shared calendars, and budgets in spreadsheets.
+Integrating DoubleAgents with these systems (e.g., via APIs or MCP servers) would allow the agent to both read from and act within the user’s existing infrastructure.
+For instance, the agent could summarize threads to Slack, create follow-up tasks, or detect stalled interactions from calendar and messaging signals.
+Beyond convenience, such integration enriches the agent’s state representation, enabling it to reason over richer, continuously updated signals and reducing the burden on users to manually maintain context.
+7.2. Generalizing to Multi-User Coordination
+Our study focuses on a single organizer, but many real-world coordination tasks involve multiple stakeholders with distinct roles.
+Distributed cognition naturally extends to such settings, where cognition is shared across people and artifacts rather than confined to an individual (Hutchins and Klausen, 1996).
+This raises new design considerations for how DoubleAgents might support team-based coordination.
+Role-specific views with shared context.
+In multi-user settings, each participant requires a role-specific interface while remaining aware of cross-role dependencies.
+For example, in workshop organization, one person may manage catering, another speaker invitations, and another the budget.
+Each role demands a tailored view (e.g., headcounts, speaker pipeline, financial tracking), yet these views are tightly coupled: budget constraints venue choice, and venue capacity constrains invitations.
+A multi-user DoubleAgents system would provide each user with their own dashboard, policies, and agent, while maintaining a shared coordination substrate that propagates updates across roles.
+Changes in one domain (e.g., new funding or speaker cancellations) would be surfaced as actionable signals in others, enabling coordinated adaptation without requiring manual information relay.
+Per-user and shared policies.
+Delegation preferences vary across individuals.
+One user may permit autonomous actions within their domain, while another may require explicit approval for every step.
+Supporting this heterogeneity requires both per-user policies, which govern behavior within individual scopes, and shared policies, which encode team-level constraints (e.g., approval requirements for large expenses).
+In this setting, stop hooks evolve into coordination mechanisms not only between a user and their agent, but across users: they can flag when an action in one role impacts another and trigger cross-role review before execution.
+This enables alignment to be negotiated at both individual and collective levels.
+From individual to team alignment.
+Our findings suggest that alignment improves through repeated interaction.
+In multi-user settings, this process becomes collective: teams develop shared expectations about when the system acts autonomously, when it defers, and how it communicates state.
+Over time, accumulated policies, templates, and stop hooks form a persistent layer of organizational memory: an encoded representation of how the team coordinates (Hutchins and Klausen, 1996).
+Similar to how teams develop shared mental models through repeated use of tools and procedures, iterative interaction with such systems can produce reusable coordination practices that transfer across team members and tasks.
 8. Limitations and Future Work
-Our work has several limitations that point to promising directions for future research.
-First, DoubleAgents was initially evaluated through simulated seminar-organizing scenarios.
-While these simulations were reviewed by domain experts and designed to reflect realistic coordination challenges, they cannot fully capture the diversity and unpredictability of real-world interactions.
-Future work could strengthen alignment by grounding both the coordination agent and respondent simulations in users’ real historical data, such as past email threads, organizing decisions, and communication patterns.
-This would enable more personalized modeling of users’ values, tone, and escalation preferences, as well as more behaviorally faithful respondent agents.
-Second, DoubleAgents currently relies on different language models for different subtasks (e.g., o3 for reasoning and 4o-mini for generation).
-Although our technical evaluation suggests this division performs adequately, future models with stronger social reasoning and long-horizon planning capabilities may further improve coordination quality.
-In addition, tighter integration between user-specific data and model reasoning, such as learning when and how to apply policies from prior interactions, remains an open challenge.
-Third, while we conducted real-world deployments with two seminar organizers, discussions with additional organizers revealed coordination scenarios that our current system does not fully support.
-For example, bilingual invitations, multi-organizer decision-making, and negotiation among organizers introduce shared authority, language switching, and conflicting preferences.
-Extending DoubleAgents to support collaborative alignment, where policies, stop hooks, and responsibilities are jointly defined and negotiated, represents an important direction for future work.
-Finally, all study participants were experienced users of AI tools, which may have positively influenced their willingness to delegate and their trust in the system.
-Future studies should include participants with a wider range of AI familiarity to better understand onboarding challenges, mental model formation, and how interactive simulation can support alignment for users with less technical background.
+Our lab study involved 10 participants in simulated scenarios, limiting generalizability, though the within-subjects design across two sessions yielded rich insights into how alignment strategies and delegation boundaries evolved with repeated use.
+Three week-long deployments validated these patterns in real-world settings.
+Larger, more diverse samples, including users with less AI familiarity, would help clarify onboarding challenges and how simulation supports alignment across experience levels.
+While our simulations reflected realistic coordination challenges, they cannot fully capture real-world unpredictability.
+Grounding both the coordination agent and respondent simulator in users’ historical data, including past emails, decisions, and communication patterns, could enable more personalized modeling of values and tone.
+Finally, DoubleAgents relies on different language models for different subtasks (e.g., o3 for reasoning, 4o-mini for generation).
+Future models with stronger social reasoning and long-horizon planning may improve coordination quality, and tighter integration between user-specific data and model reasoning remains an open challenge.
 9. Conclusion
-We introduced DoubleAgents, a system for aligning agentic AI through interactive simulation, policy-grounded planning, and transparent oversight.
-In coordination tasks like speaker scheduling, alignment is not just about correctness; it’s about respecting user preferences, adapting to edge cases, and maintaining trust.
-By embedding simulation into the agent workflow, DoubleAgents helps users rehearse edge cases, refine their preferences, and gradually scale up delegation.
-Our lab and real-world deployments show that users increasingly trusted and customized the system, not because it was perfect, but because they could iteratively shape it.
-This work offers a practical path toward alignment in real-world agentic AI: one grounded not in static prompts, but in interaction, iteration, and co-adaptation.
-10. Generative AI Usage Disclosure
-We used generative AI tools (e.g., OpenAI’s ChatGPT) as support during several parts of this project.
-Specifically, we used them to help debug the system, draft non-critical implementation and UI code, write analysis scripts for logs and study data, and edit text for grammar and clarity during paper writing.
+We present DoubleAgents, an interactive alignment tool for human coordination through user intervention, value-reflecting policies, rich state visualizations, and uncertainty flagging.
+Technical evaluations validate key system capabilities of DoubleAgents, including policy selection and edge-case detection.
+A two-session user study (N=10) shows users’ increased willingness to delegate both planning and execution, supported by policy-based alignment, interpretable visualizations, and explicit delegation boundaries.
+Three real-world deployments show that alignment learned through simulation transfers to live coordination while continuing to evolve in practice.
+Our findings frame alignment not as a one-time specification problem but as an iterative process, and position distributed cognition as a principled approach for enabling progressive delegation with retained user control.
 Appendix A Seminar Planning: A Human Coordination Problem
@@
 An effective seminar scheduling tool should not attempt to handle the process independently; instead, it must learn from and adapt to the organizer, respecting social norms, relational dynamics, and the inherent unpredictability of responses.
-Appendix B User Study Procedures and Data Analysis
+Appendix B System Implementation Details
+Figure 4.
+System architecture of DoubleAgents.
+The Policy Module supplies policies to the Coordination Agent within the Simulation Engine, which exchanges emails with Simulated Respondent Speakers to advance the task.
+When a response falls outside existing policy coverage, the Edge Case Detector escalates the situation to the user and feeds the resulting clarification back into the Policy Module as an updated policy.
+The Dashboard Visualization renders the evolving task state throughout this cycle.
+We present DoubleAgents, a system for aligning agentic AI through interactive simulation.
+Built on the ReAct framework (Yao et al., 2023), DoubleAgents supports iterative planning and execution via a team of AI agents that collaborate with users on complex coordination tasks.
+DoubleAgents takes as input a user-defined goal (e.g., “Organize a seminar with four speakers and four slots”), persona descriptions for simulated respondent agents, the user’s existing email templates, and a set of generic scheduling policies.
+Through iterative simulations, the system produces a finalized assignment of speakers to slots, along with refined policies, updated email templates, and executable stop hooks that users can later deploy in real-world agentic workflows.
+The system architecture comprises four key components: (1) a Simulation Engine that drives the iterative planning-execution loop through coordinated AI agents (Section B.1), (2) a Policy Module that governs agent reasoning and adapts to user preferences (Section B.2), (3) an Edge Case Detector that identifies situations beyond existing policy coverage and escalates them for user intervention (Section B.3), and (4) a Dashboard Visualization that provides real-time visual feedback on task state and progress (Section B.4).
+As illustrated in Figure 4, these four components interact in a loop.
+The Policy Module supplies policies to the coordination agent inside the Simulation Engine, which in turn exchanges emails with simulated respondent agents to advance the scheduling task.
+When a respondent’s reply falls outside existing policy coverage, the Edge Case Detector escalates the issue to the user and feeds the resulting clarification back into the Policy Module as an updated policy.
+Throughout this cycle, the Dashboard Visualization renders the evolving task state so that the user can monitor progress and intervene at any point.
+We describe each component in detail below.
+B.1. Simulation Engine
+The Simulation Engine is the core of DoubleAgents, implementing the iterative planning-execution loop that drives the coordination process.
+It comprises three main subcomponents: the Coordination Agent, which reasons over the task and orchestrates actions; the Simulated Respondent Agents, which emulate human behavior to create a realistic environment; and the Context Management Module, which maintains the evolving problem state.
+B.1.1. Coordination Agent
+The coordination agent is the primary reasoning component of the system, responsible for assessing the current problem state, selecting relevant policies, generating plans, and executing actions.
+It operates within the ReAct framework and proceeds through the following stages at each simulation step:
+  • Progress Understanding.
+The agent reviews the current problem context, including speaker information, email logs from previous days, and the history of user-approved plans and actions, and generates a concise summary of the current status.
+This summary indicates which days have been scheduled, what actions have been taken, and the most recent contact date for each speaker.
+The prompt used for generating the progress summary is provided in Appendix D.1.1.
+  • Policy Selection. Based on the progress summary, the agent invokes the policy selection tool to identify the most relevant policies for the current state (Section B.2).
+  • Plan Generation.
+Guided by the selected policies, the agent proposes a high-level plan (e.g., “Send email to all speakers to request availability for all seminar slots”).
+The plan and its rationale are presented to the user via the Interactive Interface (Figure 2(Middle)) for review and approval.
+Users may approve, edit, or request regeneration of the plan (Figure 7(a)).
+  • Action Generation.
+Upon plan approval, the agent generates a set of concrete actions to operationalize the plan (Figure 7(b)).
+There are three types of actions: (a) drafting emails to selected respondents, (b) sending emails to specified recipients, and (c) awaiting responses from certain respondents.
+Users retain full control over actions and may edit, remove, or request regeneration.
+Detailed prompts for plan and action generation are provided in Appendix D.1.3.
+  • Action Execution.
+The agent instantiates each approved action by invoking the appropriate tools from its toolbox (Figure 5).
+For email drafting, it calls the writeEmail() API, which generates emails given the current problem state, seminar details, user-provided templates, and historical email edits.
+Email drafting is performed using GPT-4o.
+For sending, it calls the sendEmail() API targeting the appropriate speaker email addresses.
+For delayed responses, it calls the waitResponse() tool, pausing execution and advancing to the next pending action or simulated day.
+Users interact with the coordination agent through the Interactive Interface (Figure 2(Middle)), which displays the full history of plans, actions, and emails as interactive pop-up windows.
+The interface also supports natural language queries (e.g., “What did Prof.
+Lucy Tran reply in the last email?”), powered by GPT-4o (see Appendix D.4).
+We use GPT-o3 as the backbone of the coordination agent, given its strong reasoning capabilities.
+B.1.2. Simulated Respondent Agents
+Drawing inspiration from (Park et al., 2023), we employ LLMs to simulate each respondent agent, generating actions that align with their assigned personas.
+Upon receiving emails from the coordination agent, each simulated respondent first reasons over its persona and then generates an appropriate action.
+For instance, a respondent whose persona specifies “You are very busy and reply very late.
+Normally, it takes you three follow-up emails to respond” will delay its reply accordingly, while a respondent described as prompt in communication will reply immediately.
+If a respondent’s persona includes specific preferences (e.g., “prefers presenting over Zoom”), the generated response will reflect those preferences.
+Each respondent composes and sends email responses, which advance the scheduling process and enable the user to coordinate subsequent steps.
+The prompt for the speaker agent is detailed in Appendix D.3.
+We adopt GPT-4o for email response generation via the sendEmail() tool.
+B.1.3. Context Management Module
+The Context Management Module maintains all relevant contextual information throughout the scheduling process.
+DoubleAgents represents the problem context using JSON structures, where each key denotes a specific type of contextual information (e.g., speaker profiles, time constraints, email history) and each value contains the corresponding content.
+As shown in Figure 5, the problem context includes keys such as “time info” (reflecting seminar slots), speaker details, and communication logs.
+The context is dynamically updated at each step, incorporating user feedback, agent actions, and simulated responses to maintain an up-to-date representation of the task state.
+B.2. Policy Module
+The Policy Module serves as the central mechanism through which user values and preferences are encoded and operationalized within the system.
+Policies are natural language rules that guide the coordination agent’s reasoning and planning at each step of the scheduling process.
+They are displayed in the Policy Panel (Figure 2(Left)), where users can inspect, add, edit, or delete them at any time.
+B.2.1. Seed Policies
+To bootstrap the system with broadly applicable guidelines, the lead authors designed a set of six initial “seed policies” that address common scheduling challenges: (1) individuals being reluctant to share their full availability, (2) ambiguous or unclear responses, (3) the need for multiple follow-ups to elicit responses, (4) resolving scheduling conflicts between two speakers, (5) initiating the process by broadly reaching out to potential participants, and (6) determining when to cease follow-up attempts with non-responsive participants.
+Each policy is designed to address a specific challenge and serves as a default guideline that the coordination agent can draw upon when generating plans.
+B.2.2. Policy Selection
+At each planning step, the coordination agent invokes a policy selection tool that identifies the most relevant policies given the current problem state.
+For instance, at the beginning of the scheduling process when no speaker availabilities have been collected, the system selects the policy: “If we are just beginning and have not yet contacted any speakers for availability, then ask each speaker for their availability across all slots.” The selected policy is highlighted in the Policy Panel (Figure 2(A)), providing users with a transparent rationale for the agent’s subsequent plan.
+Details of the prompts used for policy selection are provided in the Appendix D.1.2.
+B.2.3. Policy Refinement via Stop Hooks
+Policies are not static; they evolve throughout the simulation process.
+When the Edge Case Detector (Section B.3) identifies a situation not covered by existing policies and the user provides a clarification, the system records this intervention as a stop hook and uses it to refine the policy set.
+For example, if a speaker requests to present via Zoom and the user specifies a preference for in-person attendance, this clarification is converted into a new policy explicitly stating that remote participation is not preferred.
+In this way, user interventions during edge cases are transformed into reusable system policies that persist across subsequent simulation steps and can be deployed in real-world workflows.
+B.2.4. Implicit User Preference Learning
+Beyond explicit policy refinement, the system also supports implicit learning of user preferences by observing user-initiated edits to system-generated plans, actions, or emails.
+The coordination agent summarizes each edit into a concise statement, logs it in a user edit history, and incorporates it into the problem context for subsequent planning steps.
+This enables the agent to evolve its behavior in alignment with user preferences over time without requiring explicit policy authoring.
+We leverage GPT-4o for the summarization process, with prompt details provided in Appendix D.1.5.
+B.3. Edge Case Detector
+The Edge Case Detector serves as a safeguard that identifies situations falling outside the coverage of existing policies and escalates them to the user for clarification.
+It is implemented as a tool that the coordination agent invokes to check whether an incoming message from a respondent agent can be addressed using the current policy set.
+When the detector identifies an edge case—for example, a speaker requesting to present via Zoom when no existing policy addresses remote participation—it triggers a clarification request through the Interactive Chat Interface.
+The system presents the user with a concise description of the situation and a question prompting guidance (e.g., “Prof.
+Lucy Tran requests a Zoom option due to circumstances preventing in-person attendance.
+Should I confirm Zoom availability for their seminar or insist on in-person participation?”).
+The user then provides a natural language instruction (e.g., “Tell Prof.
+Lucy Tran we prefer in-person due to the potentially large audience”), which the coordination agent processes to generate an appropriate action.
+Critically, each user clarification is recorded as a stop hook and used to refine the policy set (Section B.2).
+This mechanism ensures that (1) ambiguous or novel situations are never resolved autonomously without user input, (2) user interventions are preserved as reusable policies for future reference, and (3) the policy set grows iteratively to cover an expanding range of scenarios.
+The edge case detector is implemented with GPT-4o, with prompts detailed in Appendix D.2.
+B.4. Dashboard Visualization
+To support user understanding and foster alignment, DoubleAgents provides real-time visual feedback on the task state and system progress throughout the scheduling process.
+We implement a Flask-based dashboard interface designed to make the planning workflow interactive, legible, and actionable.
+The dashboard is composed of several layered visual components, shown in Figure 2 (Right):
+  • Assignment Progress Tracker.
+This panel visualizes tentative and confirmed speaker-to-slot assignments.
+It highlights which seminar slots remain unfilled and which speakers have been successfully scheduled, enabling users to assess how close the current plan is to completion at a glance.
+  • Calendar View.
+Anchored in simulated days, this temporal layer helps users reason about timing, deadlines, and follow-up actions.
+The calendar highlights both the current simulated day and the seminar dates, allowing prioritization of actions based on urgency and scheduling constraints.
+  • Communication History.
+This panel displays the message history between the coordination agent and each speaker, offering a transparent view of the outreach and response status.
+It enables users to track ongoing communications, detect gaps, and assess follow-up needs.
+These visual layers operate in parallel, allowing the system to efficiently track and update the state of each element in real time.
+At the end of each simulated day, the system generates a daily summary that highlights key events and decisions, using GPT-4o (with prompt details in Appendix D.5).
+The summary is presented within the Interactive Interface, providing users with a consolidated overview before advancing to the next simulated day.
+Appendix C User Study Procedures and Data Analysis
 Test Scenario Preparation.
@@
 Interview data were analyzed using thematic analysis to understand how users perceive the usefulness and trust drive about each feature, and quantitative data were analyzed using paired-samples t-tests to examine whether participants’ experiences and perceptions of the DoubleAgents and AI tools across usage and days.
-Appendix C DoubleAgents Prompts
-C.1. Coordination Agent
-C.1.1. Progress Summary
+Appendix D DoubleAgents Prompts
+D.1. Coordination Agent
+D.1.1. Progress Summary
 You are an expert seminar organizer assistant.
@@
 ---------- [Task Rules] [Possible Dates] [Time Context] [Speaker Information] Here are the history of Approved Plans and Actions: [Approved Plans and Actions] -------------------------------- Here is the history of past user edits and implicit learning (recent, action edits only): [Implicit Learning] -------------------------------- Past Email Log: [Email Log]
-C.1.2. Policy Selection
+D.1.2. Policy Selection
 You are a helpful assistant.
@@
 [Implicit Learning/User Edit History (recent, action edits only)]
-C.1.3. Action Generation
+D.1.3. Action Generation
 You are an expert seminar organizer assistant.
@@
 -------------------------- Time Context: [Time Context] Task Progress Summary: [Progress Summary] -------------------------- Applied Coordination Policies: [Applied Policy] -------------------------- Implicit Learning/User Edit History (recent, action edits only): [Implicit Learning - Action Edits] --------------------------
-C.1.4. Email Drafting
+D.1.4. Email Drafting
 I am [Organizer Name], a seminar/workshop organizer.
@@
 [Implicit Learning - Email Edits]
-C.1.5. Implicit User Preference Learning
+D.1.5. Implicit User Preference Learning
 Plan and Action Edits
@@
 Inputs: history_context: {history_context}" "ai_text: {ai_text}" "user_text: {user_text}"
-C.2. Edge Case Detection
+D.2. Edge Case Detection
 You are an expert seminar organizer assistant.
@@
 Here are the policies: [Policies]
-C.3. Simulated Respondents
+D.3. Simulated Respondents
 You are simulating a seminar speaker.
@@
 Format your response as: ACTION: <reply|ask_question|wait> WAIT_UNTIL_DAY: <integer or ’none’> EMAIL: <JSON object or ’none’> RATIONALE: <your reasoning> EMAIL should be like: "from": "XX", "to": "XX", "subject": "XXX", "body": "XXXXX", "timestamp": "XX" dont output json ‘‘‘ etc
-C.4. Free-form QA Chat
+D.4. Free-form QA Chat
 You are an expert assistant.
@@
 [EMAIL LOG] [USER QUESTION]
-C.5. Daily Summary
+D.5. Daily Summary
 Today is simulation day [Day ID].
@@
 Dont include today’s date in the summary -- just summarize the progress.
-Appendix D Technical Evaluation Details
-D.1. Policy Selection
+Appendix E Technical Evaluation Details
+E.1. Policy Selection
 Here are the prompts used in the technical evaluations for policy selection:
-D.1.1. Experiment prompts
+E.1.1. Experiment prompts
 speaker-only
@@
 Main reasoning result: {progress_summary} Coordination Policies: {policy_list} Return ONLY the JSON list of indices:
-D.2. Edge Case Detection
-D.2.1. Test suite construction
+E.2. Edge Case Detection
+E.2.1. Test suite construction
 We employ GPT‑5 to generate a set of 10 edge-case (flaggable) test cases and 10 non-flagged (clear) test cases, with all examples validated for quality by the lead authors.
@@
 In total, 100 test cases are constructed and used as our test suite for edge case detection.
-D.2.2. Prompts to generate linguistic variations and test data
+E.2.2. Prompts to generate linguistic variations and test data
 You are a helpful assistant.
@@
 Please generate exactly five variations of the response for each case
-D.2.3. Experiment prompts
+E.2.3. Experiment prompts
 zero-shot
@@
 We should prioritize their availability by assigning slots to faculty over students and asking students to be more flexible or to provide additional dates to accommodate them.", - "If we are just beginning and have not yet contacted any speakers for availability, then ask each speaker for their availability across all slots." Response: response Your Output:
+Appendix F System Screenshots
+Figure 5.
+System diagram of DoubleAgents illustrating a day-by-day ReAct workflow that couples policy-guided planning with human oversight and LLM simulation.
+Inputs (top left) include the organizer’s goal, seminar slots, speaker personas, and a toolbox of callable functions.
+The coordination agent (center) iteratively: summarizes state; selects applicable policies; proposes a plan and action that the user can Regenerate or Approve; executes by drafting and sending emails, or waiting for responses.
+If a reply falls outside policy coverage, issueFlag escalates an edge case to the user for clarification before proceeding.
+The context management (right) maintains time information, past plans/actions, email logs, user preferences, and persona data, continuously updating to ground subsequent steps.
+Simulated respondent agents (bottom right) generate realistic, persona-consistent behaviors and replies that drive the loop across days.
+For a detailed walkthrough, please refer to Section 4.
+Diagram showing the architecture of the DoubleAgents system, with components for goal setting, agent coordination, context management (including edge case detection), and various submodules such as toolbox, orchestrator, and simulated respondents.
+Figure 6. The landing page to specify the user goals, seminar details, and the persona info for the speakers.
+(a) The pop-up window for a generated plan.
+(b) The pop-up window for generated actions following the plan.
+(c) The pop-up window for a generated email.
+Figure 7. Examples for plan and action generation.
+(a) Email sent by the simulated Prof. A.
+(b) Edge case detection given response from Prof. B asking about Zoom options.
+Figure 8. The example email sent by the simulated respondents and the system message when detecting edge case.